### v1

# Preparation of Papers for IEEE Sponsored Conferences & Symposia*

Abstract
Continuous-time Conflict Based-Search (CCBS) has long been viewed as the de-facto optimal solver for multi-agent path finding in continuous time (MAPFR). Recent findings, however, show that the original theoretical variant of CCBS can suffer from non-termination, while the widely used implementation can return sub-optimal solutions. We introduce an analytical framework that yields simple and sufficient conditions under which any CCBS-style algorithm is both sound, i.e., returns only optimal solutions, and solution complete, i.e., terminates on every solvable MAPFR instance. Investigating the publicly available implementation of CCBS reveals that it violates these conditions. Though this merely indicates that CCBS might be unsound, this indication is supported by counter-examples.
Leveraging the analytical framework, we propose a novel branching rule and prove that it satisfies the sufficient conditions, thereby restoring soundness and termination guarantees. Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of the discrete-time CBS for the first time in the continuous domain. We experimentally apply standard CCBS and CCBS under our branching rule to an example problem, with our branching rule returning a solution with lower sum-of-costs than standard CCBS. Because the branching rule largely only affects the branching step, it can be adopted as a drop-in replacement in existing code-bases, as we show in our provided implementation. Beyond CCBS, the analytical framework and termination criterion provide a systematic way to evaluate other CCBS-like MAPFR solvers and future extensions.
I INTRODUCTION
As robotics and automation take an increasingly larger role in society, there is a growing need to move agents from where they are to where they need to be without collision. This is the problem of Multi-agent Path Finding (MAPF) which finds application across diverse domains, from computer games [1] to warehouses [2, 3], factories [4] to airports [5]. Progress toward fast and exact solution methods has been steady for decades, targeting the many variants of MAPF.
In its classical formulation, MAPF is about finding a collision-free path for each agent on a graph from its start vertex to goal vertex. Agents are represented as points and can wait at vertices and traverse edges, with time being discrete, edges taking unit-time to traverse, and an objective function based on arrival times at goal vertices being minimized. This problem is NP-hard [6] and has many variations, comprehensively detailed in e.g. [7, 8, 9]. Early focus was placed on reducing computation and searching for sub-optimal solutions by, e.g., establishing a priority order among agents or restricting the search to a limited time-window [10, 11]. Shortly thereafter, exact methods such as $M^{*}$ [12], Increasing Cost Tree Search [13] and Conflict Based-Search (CBS) [14] were introduced. These methods cleverly avoid searching the entire solution space — which grows exponentially with the number of agents — by adopting a “lazy” approach of planning agents individually and addressing conflicts only as they occur. Of the three, CBS has enjoyed the largest adoption, with multiple continuations. To name a few, Meta-Agent CBS [15] and Improved CBS [16] for performance enhancements, Enhanced CBS [17] for bounded sub-optimal solutions, Rolling Horizon Collision Resolution [18] for lifelong MAPF, and Multi-Goal CBS [19] for MAPF where agents are assigned multiple goal vertices.
The simplifying assumptions in MAPF that time is discretized and edges take unit-time to traverse bring significant drawbacks to its practical applicability; in real-world settings, environments do not always follow grid-like structures, agent speeds and action times vary, and artificial wait times to ensure that agents move in lockstep inflates the objective value. The continuous-time MAPF variant, MAPFR [20], removes such discrete-time assumptions. In MAPFR, agents can traverse edges at any real-valued time by following a motion function that defines a continuous trajectory between two connected vertices. Additionally, agents have volumes such that two agents collide when their volumes overlap. The work in [21] addresses MAPFR for sub-optimal solutions, while [22] finds optimal solutions to a simplified MAPFR problem where agents traverse edges in straight lines at constant speeds. Currently, the only known method to address MAPFR in its entirety for optimal solutions is Continuous-time CBS (CCBS) [20] which claims to be sound and solution complete, that is, it guarantees to return an optimal solution if one exists. Since the introduction of CCBS, numerous continuation works have been published, such as [23] for performance enhancements, $T$-Robust CCBS [24] for solutions that are robust to delays, [25] for Any-Angle MAPF, Continuous-time Prioritize Lifelong Planner [26] for lifelong MAPF, and [27] where CCBS is modified and applied to quadcopter drones.
Despite CCBS’s wide-spread adoption, only recently has a critical assessment in [28] of the claims made in [20] established that CCBS does not fulfill its guarantees. The theoretical description of CCBS does not guarantee termination under the existence of an optimal solution (a condition for solution completeness) and the publicly available implementation does not guarantee that a returned solution is optimal (a condition for soundness). This is shown in [28] theoretically and with counter-examples. Therefore, these results imply that currently there does not exist a MAPFR solver for optimal solutions.
In this work, we develop an analytical framework to aid the analysis of CCBS and find sufficient but not necessary requirements for CCBS to be sound and solution complete. We then show that the branching rule in CCBS does not satisfy this requirement, which by itself does not mean that CCBS is not sound and solution complete, but it does complement the results in [28] and our experiments. Using the analytical framework, we then present a novel branching rule and prove that CCBS is sound and solution complete when using it. With this work, we therefore provide the first provably sound and solution-complete method to the MAPFR problem.
The outline of this article is the following: In Section II, we formally define the MAPFR problem, soundness, and solution completeness. CCBS is also introduced, along with details surrounding the findings in [28]. The framework for analyzing branching rules and CCBS’s algorithmic properties is presented in Section III. In Section IV, we apply the framework on CCBS’s implemented branching rule. Our branching rule is presented in Section V, along with soundness and solution-completeness proofs. Section VI presents experimental results when applying standard CCBS and CCBS using our branching rule to an example problem. Finally, conclusions can be found in VII.
II Background
In this section, we formally introduce the MAPFR problem and define what we mean by soundness and solution completeness. Thereafter, CCBS [20] and the recent findings on CCBS’s guarantees in [28] are presented.
II-A Problem Formulation
A MAPFR problem is defined in [20] by a tuple $\langle\mathcal{G},\mathcal{M},S,G,\mathit{coord},\mathcal{A}\rangle$, where $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ represents a graph with vertices $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, $\mathcal{M}$ denotes a 2D metric space, $S$ and $G$ respectively map agents to start and goal vertices, $\mathit{coord}:\mathcal{V}\rightarrow\mathcal{M}$ maps vertices to coordinates in $\mathcal{M}$, and $\mathcal{A}$ comprises a finite set of move actions.
A move action $m=(m_{\varphi},m_{D})$ offers a way for an agent to traverse an edge $(v,v^{\prime})\in\mathcal{E}$. The motion function $m_{\varphi}:[0,m_{D}]\rightarrow\mathcal{M}$ describes a continuous trajectory with duration $m_{D}$ starting at $v$ and ending at $v^{\prime}$; that is, $m_{\varphi}(0)=\mathit{coord}(v)$ and $m_{\varphi}(m_{D})=\mathit{coord}(v^{\prime})$. We denote the source and target vertices as $\mathit{from}(m)=v$ and $\mathit{to}(m)=v^{\prime}$, respectively. Each edge in $\mathcal{E}$ admits a finite number of move actions, providing multiple options for an agent to traverse an edge.
Figure 1 illustrates this concept through edge $(v,v^{\prime})$ and three associated move actions $m_{1}$, $m_{2}$, and $m_{3}$. While all move actions start and end at the same vertices, their trajectories (by $m_{1,\varphi}$, $m_{2,\varphi}$, $m_{3,\varphi}$) and durations ($m_{1,D}$, $m_{2,D}$, $m_{3,D}$) may be distinct.
Figure 1: An edge $(v,v^{\prime})\in\mathcal{E}$ and its associated move actions $m_{1}$, $m_{2}$, and $m_{3}$. Each action specifies a trajectory in $\mathcal{M}$ from $v$ to $v^{\prime}$, starting at time $t=0$ and ending after potentially different durations ($m_{1,D}$, $m_{2,D}$, $m_{3,D}$).
In addition to move actions, there are also wait actions not included in $\mathcal{A}$. A wait action $w=(w_{\varphi},w_{D})$ maps to a single vertex $v\in\mathcal{V}$ such that its motion function $w_{\varphi}:[0,w_{D}]\rightarrow\mathit{coord}(v)$ with $w_{D}\in\Re\cup\{\infty\}$. We denote $v(w)=\mathit{from}(w)=\mathit{to}(w)=v$. Wait actions allow for agents to wait at vertices, which may be necessary to e.g. avoid colliding with other agents.
A timed action $\langle a,t\rangle$ (with $a$ either a move or wait action) specifies that an agent executes $a$ starting at time $t\in\Re$, such that the motion function $a_{\varphi}$ is shifted by $t$. That is, the location in $\mathcal{M}$ of an agent executing $\langle a,t\rangle$ is $a_{\varphi}(t^{\prime}-t)$ during $t^{\prime}\in[t,t+a_{D}]$. In other words, an action specifies how an agent moves and the duration to do so (always starting at time 0), while a timed action additionally specifies when the movement begins.
An agent $i$ can execute multiple timed actions, one after the other, to thereby traverse the graph. A plan $\pi_{i}$ for an agent $i$ is a finite sequence of timed actions,
$\pi_{i}=\langle\langle a_{1}^{i},t_{1}^{i}\rangle,\langle a_{2}^{i},t_{2}^{i}\rangle,\dots,\langle a_{n}^{i},t_{n}^{i}\rangle\rangle.$ (1)
For a plan $\pi_{i}$ to be valid, it must start at time 0 ($t_{1}^{i}=0$), start and end at $i$’s start and goal vertices ($\mathit{from}(a_{1}^{i})=S(i)$ and $\mathit{to}(a_{n}^{i})=G(i)$), and every successive timed action must start when and where the previous ends ($\forall k=2,..,n$: $t_{k}^{i}=t_{k-1}^{i}+a_{k-1,D}^{i}$ and $\mathit{from}(a_{k}^{i})=\mathit{to}(a_{k-1}^{i})$). Additionally, the final action $a_{n}^{i}$ is an infinite wait action with duration $a_{n,D}^{i}=\infty$ to capture that agents always remain idle indefinitely at their last location. The duration of $\pi_{i}$ is $\pi_{i,D}=\sum_{k=1}^{n-1}a_{k,D}^{i}$, naturally excluding the final infinite wait action.
Figure 2 illustrates an example of a valid plan $\pi_{i}$ starting at $S(i)$ and ending at $G(i)$. The first timed move action is $\langle m_{1}^{i}=\langle m_{1,\varphi}^{i},1.5\rangle,0\rangle$, which specifies that agent $i$ begins at $t=0$ to move according to the motion function $m_{1,\varphi}^{i}$ which has duration $1.5$. This is followed by the timed wait action $\langle w_{2}^{i}=\langle w_{2,\varphi}^{i},0.5\rangle,1.5\rangle$ specifying that $i$ waits at $v_{2}$ from $t=1.5$ for a duration of $0.5$, that is until $t=2$. Thereafter, $i$ executes $\langle m_{3}^{i}=\langle m_{3,\varphi}^{i},1\rangle,2\rangle$ to arrive at $v_{3}$ at $t=3$, and then executes $\langle m_{4}^{i}=\langle m_{4,\varphi}^{i},1.3\rangle,3\rangle$ to arrive at $v_{4}$ at $t=4.3$. Finally, the timed wait action $\langle w_{5}^{i}=\langle w_{5,\varphi}^{i},\infty\rangle,4.3\rangle$ specifies that $i$ waits at $v_{4}$ indefinitely from $t=4.3$. Naturally, the plan $\pi_{i}$ illustrated in the example represents merely one among countless valid plans; alternative wait actions could be employed at any of the vertices, different move actions could be selected for the same edges, or entirely different paths through the graph could be traversed.
Figure 2: An illustration of a valid plan $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ with action durations $m_{1,D}^{i}=1.5$, $w_{2,D}^{i}=0.5$, $m_{3,D}^{i}=1$, $m_{4,D}^{i}=1.3$, $w_{5,D}^{i}=\infty$.
Each agent occupies a non-zero volume; a collision occurs when the volumes of two agents overlap. A conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ arises when agents $i$ and $j$ collide at some time while respectively executing $\langle a^{i},t^{i}\rangle$ and $\langle a^{j},t^{j}\rangle$. A conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$ with two move actions is a move-move conflict, and a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with a move and a wait action is a move-wait conflict. By convention, the move action in a move-wait conflict is stored in the first position. Note that a wait-wait conflict can only arise after a move-move or move-wait (assuming agents start in collision-free positions). Therefore, we sufficiently consider only move-move and move-wait conflicts. Two plans $\pi_{i}$ and $\pi_{j}$ conflict if there exists a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$.
A joint plan $\Pi$ is a set containing one valid plan for each agent. A solution is a joint plan where no two plans conflict. An optimal solution $\Pi^{*}$ is a solution that minimizes an objective function $\sigma$ over all solutions. The sum-of-costs (sum over plan durations) and makespan (maximum over plan durations) are the two most common objective functions in the MAPF context. Here, we assume that $\sigma(\Pi)$ is strictly monotonically increasing with respect to the maximum duration over all agent plans, $\max_{\pi\in\Pi}\pi_{D}$, which both the sum-of-costs and makespan satisfy.
II-B Algorithmic Properties
In this article, we say that a solution algorithm to the MAPFR problem is
  • sound if it only returns optimal solutions;
  • solution complete if it is guaranteed to terminate with a solution on any solvable MAPFR problem.
In [20], a sound algorithm returns only conflict-free solutions, while an exact algorithm [^1] In [20], the term optimal is used synonymously with exact. However, an algorithm’s optimality typically refers to its time and space complexity and not to the quality of the solutions that it returns. returns only optimal solutions. These terms are never addressed separately in [20] or here. Therefore, we denote both by sound for simplicity. Also note that completeness guarantees the return of a solution if one exists or the report of no solution otherwise, while the weaker solution completeness only guarantees the former.
II-C Continuous-time Conflict Based Search
CCBS is argued to be a sound and solution-complete MAPFR solver for minimal sum-of-costs [20]. CCBS closely follows CBS [14] by simultaneously building and searching a binary constraint tree (CT).
To explain the CT and how it is searched, we begin by briefly introducing CCBS’s underlying planner Constrained Safe Interval Path Planning (CSIPP) which is covered in more detail in Section V-C2. CSIPP is a variant of SIPP [29] that computes a valid plan for a single agent with minimal duration and that satisfies a given set of constraints $C$. In SIPP, an $A^{*}$ [30] search is performed in the vertex-safe interval state space, with edges in $\mathcal{E}$ providing transitions from one state to the next. In CSIPP, it is instead move actions that provide transitions. Constraints in this context (discussed later in this section) constrain the execution of timed actions. Therefore, CSIPP finds the shortest-duration plan which does not include any timed actions forbidden by constraints in $C$.
In the CT, a node $N$ is associated with a set of constraints $N_{C}$ and a joint plan $N_{\Pi}$. The plans in $N_{\Pi}$ are computed using CSIPP given the constraint set $N_{C}$. Therefore, every plan $\pi_{i}\in N_{\Pi}$ is valid and has minimal duration under the constraints in $N_{C}$. The CCBS search starts at the CT root node $N^{R}$, with constraint set $N^{R}_{C}=\varnothing$. Since $N^{R}_{C}$ is empty, every plan $\pi_{i}\in N^{R}_{\Pi}$ is unconstrained (apart from being valid) and therefore has the shortest possible duration over all valid plans for the corresponding agent $i$. At every iteration of CCBS, a CT node $N$ minimizing $\sigma(N_{\Pi})$ over all unexpanded nodes is selected for expansion. If the joint plan $N_{\Pi}$ is a solution (i.e., conflict-free), then $N_{\Pi}$ is returned as an optimal solution. Otherwise, a branching rule is applied to $N$ to spawn two child nodes. To do so, an arbitrary conflict in $N_{\Pi}$ is selected (say between agents $i$ and $j$) and used to construct a pair of constraints $\langle c_{i},c_{j}\rangle$. These constraints $c_{i}$ and $c_{j}$ respectively constrain the plans of agents $i$ and $j$ to avoid this specific conflict. The child nodes, $N^{i}$ and $N^{j}$, each inherit one of the constraints in addition to all constraints at $N$: $N^{i}_{C}=N_{C}\cup\{c^{i}\}$ and $N^{j}_{C}=N_{C}\cup\{c^{j}\}$. The plans in $N_{\Pi}^{i}$ and $N_{\Pi}^{j}$ are then computed using CSIPP to satisfy $N_{C}^{i}$ and $N_{C}^{j}$, respectively.
We clarify the above by introducing an example in Figure 3, which will later be revisited in more detail. In this example, we consider only two agents $i$ and $j$. At the CT node $N$ (Figure 3(a)) a conflict is detected in the joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$, illustrated in Figure 3(b). Specifically, $i$ and $j$ collide while respectively executing the timed actions $\langle m^{i}_{2},t^{i}_{2}\rangle$ and $\langle m^{j}_{3},t^{j}_{3}\rangle$. Therefore, $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ is a conflict. Based on this conflict and a branching rule, a pair of constraints $\langle c^{i},c^{j}\rangle$ is created. We will explore these constraints in more detail later; for now it is sufficient to know that $c^{i}$ constrains when $i$ may execute $m^{i}_{2}$, and $c^{j}$ constrains when $j$ may execute $m^{j}_{3}$, to avoid this specific conflict. Consider the first child node $N^{i}$ (Figure 3(a)) with constraints $N^{i}_{C}=N_{C}\cup\{c^{i}\}$. Using CSIPP, the plans in $N^{i}_{\Pi}$ are computed to satisfy $N^{i}_{C}$. In practice, it is sufficient to only compute $i$’s plan and let all other agents inherit their plans from the parent node, since only $i$ has additional constraints. Figure 3(c) illustrates $N^{i}_{\Pi}$, where agent $i$ now takes a different path through the graph. Similarly, the other child node $N^{j}$ has constraints $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, and a new plan for $j$ is computed using CSIPP. The joint plan $N^{j}_{\Pi}$ is illustrated in Figure 3(d), where $j$ now waits for some time at a vertex before continuing. The resulting joint plans at the child nodes happen to be conflict-free and are therefore both solutions, in this particular example. In general, however, the child nodes’ joint plans may contain conflicts, requiring further CCBS iterations until a solution is found.
(a) The CT node NN and its two children $N^{i}$ and $N^{j}$.
(b) The joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$ contains a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$.
(c) The joint plan $N_{\Pi}^{i}$ satisfying $N^{i}_{C}$.
(d) The joint plan $N_{\Pi}^{j}$ satisfying $N^{j}_{C}$.
Figure 3: An example of a branching rule applied to a CT node $N$. Since $N_{\Pi}$ contains a conflict, two child nodes $N^{i}$ and $N^{j}$ are spawned with additional constraints and recomputed plans.
The soundness and solution-completeness proofs of CCBS [20] rely on the CT satisfying the following properties, where $N$ is a parent to $N^{i}$ and $N^{j}$ and $\mathcal{S}(N)$ is the set of all solutions satisfying $N_{C}$:
$\mathcal{S}(N)$ $=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j}),$ (2)
$\sigma(N_{\Pi})$ $\leq\min\left(\sigma(N^{i}_{\Pi}),\sigma(N^{j}_{\Pi})\right).$ (3)
Property (2) ensures that no solutions are removed when branching a node. At the CT root $N^{R}$, since $N^{R}_{C}=\varnothing$ it holds that all solutions to a MAPFR problem are contained in $\mathcal{S}(N^{R})$. If property (2) holds then $\mathcal{S}(N^{R})=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j})$ (with $N^{i}$ and $N^{j}$ being children of $N^{R}$) which recursively implies that all solutions are reachable from the CT root. Since the search starts at $N^{R}$, all solutions to a MAPFR problem can be found. Since it is the branching rule that determines $\mathcal{S}(N^{i})$ and $\mathcal{S}(N^{j})$ (by the constraint pair that it constructs), it is the branching rule that determines if property (2) is satisfied.
Property (3) is automatically satisfied by CSIPP. Since plans in $N_{\Pi}$ are computed using CSIPP given constraints $N_{C}$, every plan $\pi_{i}\in N_{\Pi}$ is the shortest-duration valid plan over all valid plans satisfying $N_{C}$. For all nodes $N^{\prime}$ below $N$ we know that $N^{\prime}_{C}\supset N_{C}$ since constraints are only added during a branching, never removed. Adding constraints can never result in shorter-duration plans. Therefore, it holds for all agents $i$ that $\pi_{i,D}\leq\pi_{i,D}^{\prime}$ where $\pi_{i,D}\in N_{\Pi}$ and $\pi_{i,D}^{\prime}\in N_{\Pi}^{\prime}$. Since $\sigma(\Pi)$ is strictly monotonically increasing with the maximum plan duration over all plans $\pi\in\Pi$, it follows that $\sigma(N_{\pi})\leq\sigma(N_{\pi}^{\prime})$, thereby satisfying property (3).
In this article, we focus our attention on the theoretically described branching rule in [20] (the theoretical branching rule) and the publicly available implementation of CCBS [^2] https://github.com/PathPlanning/Continuous-CBS. (the implemented branching rule). Note that details regarding how to implement the theoretical branching rule are provided in [20], however, these details do not match the implemented branching rule and are therefore not considered here.
II-C1 The theoretical branching rule
Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ with $\pi_{i},\pi_{j}\in N_{\Pi}$, the earliest time $t^{i}_{u}>t^{i}$ where agent $i$ executing $\langle a_{i},t^{i}_{u}\rangle$ does not conflict with $j$ executing $\langle a^{j},t^{j}\rangle$ is found. This forms the unsafe interval $t^{i},t^{i}_{u})$. From this, a motion constraint $c_{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$ is created, specifying that $i$ is forbidden from executing $\langle a^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$. Similarly for $j$, the unsafe interval $t^{j},t^{j}_{u})$ is found, forming the motion constraint $c_{j}=\langle j,a_{j},[t^{j},t^{j}_{u})\rangle$. Therefore the theoretical branching rule uses the constraint pair $\langle\langle i,a_{i},[t^{i},t^{i}_{u})\rangle,\langle j,a_{j},[t^{j},t^{j}_{u})\rangle\rangle$ to branch on $N$. The CT under the theoretical branching rule is shown in [20] to satisfy properties (2) and (3), by which they then conclude that CCBS is sound and solution complete.
Let us now revisit the example in Figure 3, where the conflict $\langle\langle m_{2}^{i},t_{2}^{i}\rangle,\langle m_{3}^{j},t_{3}^{j}\rangle\rangle$ was detected. From this conflict, the motion constraints $c^{i}=\langle i,m_{2}^{i},[t_{2}^{i},t^{i}_{u})\rangle$ and $c^{j}=\langle j,m_{3}^{j},[t_{3}^{j},t^{j}_{u})\rangle$ are formed. At node $N^{i}$ with $N^{i}_{C}=N_{C}\cup\{c^{i}\}$, agent $i$ is no longer permitted to execute $\langle m_{2}^{i},t\rangle$ for $t\in[t_{2}^{i},t^{i}_{u})$. Therefore, an alternative plan is found where $i$ avoids executing $m_{2}^{i}$ entirely. Similarly at node $N^{j}$ where $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, agent $j$ is no longer permitted to execute $\langle m_{3}^{j},t\rangle$ for $t\in[t_{3}^{j},t^{j}_{u})$. Therefore, an alternative plan is found where $j$ waits long enough to eventually execute $m^{j}_{3}$ some time after $t_{u}^{j}$.
II-C2 The implemented branching rule
The implemented branching rule differs from the theoretical implementation by how timed wait actions $\langle w^{j},t^{j}\rangle$ in move-wait conflicts $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ are handled. For such a conflict, the collision interval $I^{c}$ is found. A collision interval is defined as the first contiguous time interval during which $i$ and $j$ are in collision. Then, a vertex constraint $c^{j}=\langle j,v(w^{j}),I^{c}\rangle$ is created, forbidding $j$ from occupying vertex $v(w^{j})$ at any time during $I^{c}$. Besides this, timed move actions in both move-move and move-wait conflicts are handled in the same way as in the theoretical branching rule, resulting in the motion constraint $c^{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$.
Therefore, the implemented branching rule uses the constraint pair $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,a^{j},[t^{j},t^{j}_{u})\rangle\rangle$ for move-move conflicts and $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$ for move-wait conflicts. No proofs of soundness and solution completeness are provided for CCBS under the implemented branching rule.
II-D Revisiting CCBS
The work in [28] calls the algorithm properties of CCBS into question. The claims are two-fold:
  1. 1.
    under the theoretical branching rule, CCBS does not guarantee termination;
  2. 2.
    under the implemented branching rule, CCBS violates property (2).
At its core, the identified issue lies in how wait actions are handled when resolving conflicts.
For Claim 1[28] highlights that under the theoretical branching rule on a wait action $\langle w^{i},t^{i}\rangle$ (where the motion constraint $\langle i,w^{i},[t^{i},t^{i}_{u})\rangle$ forbidding the execution of $w^{i}$ starting at any time in $t^{i},t^{i}_{u})$ is created), only one wait duration $w^{i}_{D}$ out of infinitely many real-valued durations is removed from the search. E.g., a wait action $w=(w^{i}_{\varphi},w^{i}_{D}+\epsilon)$ with $\epsilon$ arbitrarily small is still permitted during $t^{i},t^{i}_{u})$. Therefore, despite satisfying property (2) such that an optimal solution is reachable from the CT root, there may exist infinitely many nodes between them. Hence, CCBS under the theoretical branching rule does not always terminate after finite iterations, despite the existence of a solution. Therefore, CCBS under the theoretical branching rule is not solution complete. In other words, [28] identifies that properties (2) and (3) are not sufficient to guarantee solution completeness.
For Claim 2, [28] provides two counter-examples showing how the implemented branching rule on a move-wait conflict removes valid solutions from the search (violating property (2)), and consequently returning sub-optimal solutions. Thus, CCBS under the implemented branching rule does not guarantee to only return optimal solutions and is therefore not sound.
Furthermore, [28] introduces shifting constraints. They show how a branching rule using shifting constraints, which relies on a parameter $\delta\geq 0$, does not remove collision-free solutions from the search and is therefore an improvement on the implemented branching rule. However, they conclude that if $\delta>0$ then there will always remain joint plans with conflicts, leading to potentially infinitely many nodes being searched. If instead $\delta=0$ then a similar issue as for the theoretical branching rule arises, also potentially requiring the expansion of infinitely many nodes. That is, CCBS using shifting constraints has the same properties as using the theoretical branching rule: sound but without termination guarantees.
III Preliminaries
This section establishes the theoretical foundations for analyzing branching rules in CCBS. We first introduce a decomposed form of CCBS constraints and extend the notion of sound constraint pairs to constraint sets. We then define a sound branching rule as one that preserves all solutions during node expansion and prove that a branching rule is sound if and only if its underlying constraint pair is sound. Using these definitions, we show that a sound branching rule is sufficient to ensure that CCBS is sound, and is also solution complete if CCBS is guaranteed to terminate. Finally, we introduce the intersection interval which captures the complete temporal overlap between a moving agent and a stationary one — useful for designing sound branching rules that handle move-wait conflicts.
III-A Sound Branching
A definition for a sound pair of constraints $\langle c_{1},c_{2}\rangle$ is provided in [20, 31]. Let the set $\mathbf{}{S}$ contain all solutions to a given MAPFR problem, and let $c(\Pi)$ be true if a joint plan $\Pi$ satisfies a constraint $c$, false otherwise.
Definition III.1 (Sound pair of constraints [20, 31]).
For a given MAPFR problem with solutions $\mathbf{}{S}$, a pair of constraints $\langle c_{1},c_{2}\rangle$ is sound if
$c_{1}(\Pi)\vee c_{2}(\Pi),\quad\forall\Pi\in\mathbf{}{S}.$ (4)
Throughout this article, we shall consider a decomposed form of the motion and vertex constraints by describing them as constraint sets comprised of forbidding constraints. These various constraint forms are equivalent, however, the decomposed form allows us to analyze constraint pairs used in branching rules more explicitly than when using motion and vertex constraints. A forbidding constraint $\langle i,a,t\rangle$ forbids $i$ from executing the specific timed action $\langle a,t\rangle$. A motion constraint $c=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$, which forbids $i$ from executing action $a_{i}$ starting at any time $t\in[t^{i},t^{i}_{u})$, is equivalent to the dense constraint set
$C=\left\{\langle i,a^{i},t\rangle\mid t\in[t^{i},t^{i}_{u})\right\}$
such that $c(\Pi)\Leftrightarrow C(\Pi)$ where $C(\Pi)=\bigwedge_{c^{\prime}\in C}c^{\prime}(\Pi)$. We denote this equivalence by $c\Leftrightarrow C$. Similarly, a vertex constraint $c=\langle i,v,I^{c}\rangle$ which forbids $i$ from occupying $v$ at any time $t\in I^{c}$ is equivalent to $C=\left\{\langle i,w,t\rangle\mid v(w)=v,[t,t+w_{D}]\cap I^{c}\neq\varnothing\right\}$. That is, $c\Leftrightarrow C$.
We now provide a corresponding soundness definition for a pair of constraint sets, and then define soundness of a branching rule. Then, in Lemma III.1, we show that a branching rule using a pair of constraint sets is sound if and only if the constraint set pair is sound.
Definition III.2 (Sound pair of constraint sets).
For a given MAPFR problem, a pair of constraint sets $\langle C_{1},C_{2}\rangle$ is sound if $\forall\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$, $\langle c_{1},c_{2}\rangle$ is sound.
This definition follows quite naturally from extrapolating Definition III.1 to sets of constraints: given a pair of constraints $\langle c_{1},c_{2}\rangle$ and their equivalent constraint sets $C_{1}\Leftrightarrow c_{1}$ and $C_{2}\Leftrightarrow c_{2}$, it holds for all solutions $\Pi\in\mathbf{}{S}$ that
$\quad c_{1}(\Pi)\vee c_{2}(\Pi)$
$\Leftrightarrow$ $\quad C_{1}(\Pi)\vee C_{2}(\Pi)$
$\Leftrightarrow$ $\quad\left(\bigwedge_{c_{1}^{\prime}\in C_{1}}c_{1}^{\prime}(\Pi)\right)\vee\left(\bigwedge_{c_{2}^{\prime}\in C_{2}}c_{2}^{\prime}(\Pi)\right)$
$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:c_{1}^{\prime}(\Pi)\vee c_{2}^{\prime}(\Pi)$
$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\text{ is sound}.$
Definition III.3 (Sound branching rule).
A branching rule applied to a CT node $N$ to spawn children $N^{1}$ and $N^{2}$ is sound if property (2), $S(N)=S(N^{1})\cup S(N^{2})$, always holds.
Lemma III.1 (Equivalence of sound branching and sound constraint set pair).
Let a branching rule applied to a CT node $N$ use a constraint set pair $\langle C_{1},C_{2}\rangle$ to spawn two children $N^{1}$ and $N^{2}$, such that $N^{1}_{C}=N_{C}\cup\left\{C_{1}\right\}$ and $N^{2}_{C}=N_{C}\cup\left\{C_{2}\right\}$. The branching rule is sound iff $\langle C_{1},C_{2}\rangle$ is sound.
Proof.
To prove the branching rule’s soundness, we must show that property (2), $\mathcal{S}(N)=\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$, is always satisfied. For each child $N^{i}$, $i=1,2$, the solutions permitted under the constraints $N^{i}_{C}$ are
$\mathcal{S}(N^{i})$ $=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N^{i}_{C}\right\}$
$=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N_{C}\cup\left\{C_{i}\right\}\right\}$
$=\left\{\Pi\in\mathcal{S}(N)\mid C_{i}(\Pi)\right\}.$
Therefore, for the branching rule to be sound it must hold for every solution $\Pi\in\mathcal{S}(N)$ that $C_{1}(\Pi)\vee C_{2}(\Pi)$. For every $\Pi\in\mathcal{S}(N)$, consider two cases:
  1. 1.
    $C_{1}(\Pi)\vee C_{2}(\Pi)$ holds, then we immediately know that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$;
  2. 2.
    otherwise, assume without loss of generality that $\exists c_{1}\in C_{1}:\neg c_{1}(\Pi)$.
For the latter case, we start by assuming that $\langle C_{1},C_{2}\rangle$ is sound. By Definition III.2, every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$ is sound, meaning $c_{1}(\Pi)\vee c_{2}(\Pi)$ holds. Since we assumed $\neg c_{1}(\Pi)$, it follows that $c_{2}(\Pi)$ must be true. This reasoning applies to every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$, such that $\forall c_{2}\in C_{2}:c_{2}(\Pi)$ holds. This means that $C_{2}(\Pi)$ is true and therefore that $\Pi\in\mathcal{S}(N^{2})$. Therefore, if $\langle C_{1},C_{2}\rangle$ is sound then it holds for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is sound then the branching rule is sound.
Assume instead that $\langle C_{1},C_{2}\rangle$ is unsound: then by Definition III.2 there exists an unsound pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$. Since $\langle c_{1},c_{2}\rangle$ is unsound, there exists a valid solution $\Pi$ such that $\neg(c_{1}(\Pi)\vee c_{2}(\Pi))$, implying $\neg C_{1}(\Pi)$ and $\neg C_{2}(\Pi)$. Therefore, $\Pi\not\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Since $\mathcal{S}(N)$ can contain such a solution $\Pi$, it does not always hold for all $\Pi\in\mathcal{S}(N)$ that $\Pi\in\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$. Thus, if $\langle C_{1},C_{2}\rangle$ is unsound then the branching rule is unsound.
Finally, we put branching rules back into the context of the CCBS algorithm by showing in Theorem III.2 that CCBS is sound when using a sound branching rule. Additionally, we show in Theorem III.3 that CCBS is solution complete if it is both sound and guaranteed to terminate.
Theorem III.2 (Soundness).
CCBS is sound when using a sound branching rule.
Proof.
Property (2) is satisfied since the branching rule is sound, such that all solutions are reachable from the CT root. CCBS performs a best-first search starting at the root, always expanding the CT node with minimum objective value $\sigma(N_{\Pi})$. Since property (3) is satisfied (by CSIPP), expanding a node cannot spawn a new node with a better objective value. Therefore, the objective value of encountered CT nodes increases monotonically. This means that CCBS will encounter any optimal solution before encountering any sub-optimal solution.
CCBS finds and returns the first solution that it encounters, which must be the optimal solution among all reachable solutions (which includes all valid solutions). Therefore, if the branching rule is sound then CCBS is sound.
Theorem III.3 (Solution completeness).
CCBS is solution complete if it is sound and guaranteed to terminate.
Proof.
CCBS’s soundness ensures that all solutions are reachable from the CT root. CCBS has exactly one termination condition: when it selects a CT node $N$ such that $N_{\Pi}$ is a solution, $N_{\Pi}$ is returned. CCBS continues expanding nodes until this condition is met. Since all solutions are reachable from the CT root where CCBS begins its search, and CCBS is guaranteed to terminate, it must eventually encounter and return a solution. Therefore, CCBS is solution complete if it is sound and guaranteed to terminate. ∎
Note that Theorem III.2 shows that CCBS is sound if the branching rule is sound. However, it does not show that CCBS is unsound if the branching rule is unsound. Thus, the branching rule being sound is a sufficient but not necessary condition for CCBS to be sound.
III-B Intersection Interval
We now define the intersection interval which will be useful in Section V when introducing a our sound branching rule. Informally, an intersection interval is the time interval when an agent executing a timed move action is in collision with a stationary agent waiting indefinitely at a vertex.
Let a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ be given, illustrated in Figure 4. Suppose now that agent $j$ waits at $v(w^{j})$ for all time instead of only during $t^{j},t^{j}+w_{D}]$. Let all maximally connected time intervals when the agents are in collision be collected in the set of intersection intervals $\bar{\mathbb{I}}$. Since we know that the agents collide, during $^{c}, 𝕀¯\bar{\mathbb{I}}$ contains at least one interval super set of $I^{c}$.
Definition III.4 (Intersection interval).
Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$, let the intersection interval be the interval $\bar{I^{c}}\in\bar{\mathbb{I}}$ such that $I^{c}\subseteq\bar{I^{c}}$.
Figure 4 illustrates the relation between $I^{c}$, $\bar{I^{c}}$ and $\bar{\mathbb{I}}$. It is useful to note that $I^{c}$ is the intersection between $j$’s occupancy of vertex $v(w^{j})$ and the intersection interval $\bar{I^{c}}$,
$I^{c}=[t^{j},t^{j}+w^{j}_{D}]\cap\bar{I^{c}}.$ (5)
In the continuous-time setting, we assume all collision intervals, unsafe intervals, and intersection intervals are non-degenerate. Singular intervals (where overlap occurs only at a single time instant) have no practical meaning in the continuous-time setting as such situations can ambiguously be interpreted as either a collision or not a collision. In other words, we adopt the interpretation of no collision occurring when agent volumes overlap during a singular time interval.
Figure 4: A timeline of conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with conflict interval $I^{c}$ (red), the set of intersection intervals $\bar{\mathbb{I}}$ (black bars along the move action’s timeline), and the intersection interval $\bar{I^{c}}$.
IV Analyzing CCBS’s implemented branching rule
We use the framework from Section III to show that the implemented branching rule is not sound. Therefore, CCBS under the implemented branching rule does not satisfy the sufficient but not necessary condition for soundness from Theorem IV.2: a sound branching rule. Additionally, since the proofs in [20] for CCBS soundness rely on property (2) being satisfied, which is not fulfilled under an unsound branching rule, those proofs are not applicable either. Thus, this section provides complementary evidence to the counter-examples in [28] and Section VI showing that CCBS using the implemented branching rule is indeed not sound.
Lemma IV.1 provides intermediate results that are used in Theorem IV.2 to show that the implemented branching rule is not sound.
Lemma IV.1 (Collision timing for time-shifted move actions).
Let no collision occur when agent $i$ executes $\langle m^{i},t^{i}\rangle$ and $j$ executes $\langle w^{j},t^{j}\rangle$. If $i$ instead executes $\langle m^{i},t^{i}-\epsilon\rangle$ for some $\epsilon>0$, then a collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur within the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
Proof.
Since $j$ is stationary at $v(w_{j})$ throughout $t^{j},t^{j}+w^{j}_{D}]$ and the move action $m^{i}$ is identical in both $\langle m^{i},t^{i}\rangle$ and $\langle m^{i},t^{i}-\epsilon\rangle$, we can analyze when a collision occurs by comparing agent $i$’s motion in both cases. When agent $i$ executes $\langle m^{i},t^{i}-\epsilon\rangle$, its motion during the interval $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ corresponds exactly to its motion during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$. Since no collision occurred during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$, and agent $j$ remains stationary at the same location $v(w^{j})$, no collision can occur during $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ in the shifted case. Therefore, any collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ must occur during the remaining time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. ∎
Figure 5: Timeline of a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ (red). The resulting constraints forbid $i$ from executing $m^{i}$ starting at $t^{i},t^{i}_{u})$ (orange) and $j$ from occupying $v(w^{j})$ during $I^{c}$ (purple). The timed action $\langle m^{i},t^{i}_{u}\rangle$ and two forbidden timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ are shown.
Theorem IV.2 (Unsoundness of the implemented branching rule).
The implemented branching rule is not sound.
Proof.
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$. From this, the implemented branching rule uses the constraint pair $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$. Letting $C_{i}\Leftrightarrow\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $C_{j}\Leftrightarrow\langle j,v(w^{j}),I^{c}\rangle$, the branching rule equivalently uses the constraint set pair $\langle C_{i},C_{j}\rangle$. Figure 5 illustrates a timeline of this conflict and the constrained intervals $t^{i},t^{i}_{u})$ and $I^{c}$.
By construction of $C_{i}$, we know that $c_{i}=\langle i,m^{i},t^{i}_{u}-\epsilon\rangle\in C_{i}$ for an arbitrarily small $\epsilon>0$. By definition of $t^{i}_{u}$, $\langle m^{i},t^{i}_{u}\rangle$ and $\langle w^{j},t^{j}\rangle$ do not conflict. It then follows by Lemma IV.1 that any collision between agent $i$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $j$ executing $\langle w^{j},t^{j}\rangle$ must occur during the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. See Figure 5 for the timelines of the timed move actions $\langle m^{i},t^{i}_{u}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$.
The constraint set $C_{j}$ contains one forbidding constraint for each timed action that requires occupying $v(w^{j})$ during $I^{c}$. We now aim to find a forbidden timed wait action $\langle w^{\prime},t^{\prime}\rangle$ ($\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$) that does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$. Let
$t^{\prime},t^{\prime}+w^{\prime}_{D}]\subseteq[t_{j},t_{j}+w^{j}_{D}],$
that is, we restrict our search to timed wait actions that execute during a subset time interval of when $\langle w^{j},t^{j}\rangle$ is executed. Recall that a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. Therefore, the same is true when executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$, a collision can only occur during the same interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ (assuming $\epsilon<w^{\prime}_{D}$). From this, we now constrain $t^{\prime}+w^{\prime}_{D}<t^{j}+w^{j}_{D}-\epsilon$ such that $\langle w^{\prime},t^{\prime}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ do not collide:
$t^{\prime},t^{\prime}+w^{\prime}_{D}]\subseteq[t^{j},t^{j}+w^{j}_{D}-\epsilon).$ (6)
Finally, for $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$ we require $t^{\prime},t^{\prime}+w^{\prime}_{D}]$ to intersect with $I^{c}$. Since $\epsilon$ is arbitrarily small, let $\epsilon<|I^{c}|$. It follows that the interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ of size $\epsilon$ can only span at most a subset of $I^{c}$. Additionally, by (5), $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ such that $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ must be a suffix of $I^{c}$. The interval $t_{1}^{c},t^{j}+w^{j}_{D}-\epsilon)$ is therefore non-degenerate. Thus, we require (6) and $t_{1}^{c}\leq t^{\prime}+w^{\prime}_{D}$ to hold. Since any $t^{\prime}\in\Re$ and $w^{j}_{D}\in\Re$ can be selected, we know that there exists a forbidding constraint $c_{j}=\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$. See Figure 5 for an example of such a timed wait action $\langle w^{\prime},t^{\prime}\rangle$.
In conclusion, the timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ do not conflict; $i$ collides with $j$ stationary at $v(w^{j})$ only at some time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$, while $j$ executing $\langle w^{\prime},t^{\prime}\rangle$ does not occupy $v(w^{j})$ at any time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$. Therefore, a solution $\Pi$ can contain both of these timed actions. However, such a solution does not satisfy $c_{i}(\Pi)\vee c_{j}(\Pi)$. Thus, $\langle c_{i},c_{j}\rangle$ is not sound, hence neither is $\langle C_{i},C_{j}\rangle$. By Lemma III.1, the implemented branching rule is not sound. ∎
V A Branching Rule for Sound and Solution-Complete CCBS
In this section, we present our novel branching rule and prove that CCBS using this branching rule is sound and solution complete. To do so, we first show in Section V-A that our branching rule is sound. To provide termination guarantees, we formulate a non-termination condition in Section V-B and then show in Section V-C that CCBS under our branching rule does satisfy this condition. Finally, we use these results to conclude in Section V-D that CCBS under our branching rule is sound and solution complete.
Our branching rule is based on shifting constraints [28] can be understood by considering a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. By construction of $\bar{I^{c}}$, when $i$ executes $\langle m^{i},t^{i}\rangle$ then a collision occurs if $j$ occupies $v(w^{j})$ at any time in $\bar{I^{c}}$. Suppose now that $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ with $\Delta t\in\Re$ instead. Then, the intersection interval will also shift by $\Delta t$, such that a collision occurs if $j$ occupies $v(w^{j})$ at anytime in $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$. Fig. 6 illustrates the timeline of $m^{i}$ starting at $t^{i}$ and incrementally later times. Now, let some $0\leq\delta<|\bar{I^{c}}|$ be given. If $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, with shifted intersection interval $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$, then we know for any value of $\Delta t$ that the interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ will be included in the resulting intersection interval. In other words, if $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, then we know that a collision will arise if agent $j$ occupies $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. This result is useful as it allows us to form a pair of sound constraints from a move-wait conflict.
Figure 6: If move action $m^{i}$ is executed at any time in the (orange) interval $t^{i},t^{i}+\delta)$, then $i$ will necessarily intersect with $j$ waiting at $v(w)$ during the (blue) interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
We now define our branching rule.
Definition V.1 (Our novel branching rule).
Move-move and move-wait conflicts are handled differently:
  • Given a move-move conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$, the theoretical branching rule is applied. That is, the unsafe intervals $t^{i},t^{i}_{u})$ and $t^{j},t^{j}_{u})$ are found. Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$, and $j$ is forbidden from executing $\langle m^{j},t\rangle$ for any $t\in[t^{j},t^{j}_{u})$.
  • Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$, let
    $\delta=\min\left(\gamma|\bar{I^{c}}|,\;t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$ (7)
    where $0<\gamma<1$ is a fixed constant. Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}+\delta)$. Agent $j$ is forbidden from occupying the vertex $v(w_{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (vertex constraint) and executing any timed move action $\langle m,t\rangle$ where $\mathit{from}(m)=v(w^{j})$ and $t\in[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
The fixed constant $\gamma\in(0,1)$ dictates which agent in a move-wait conflict is more heavily constrained: setting $\gamma$ near $1$ increases the constrained interval on agent $i$ while setting it closer to 0 increases the constrained interval on $j$. Figure 7 illustrates three different cases where our branching rule is applied to a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. In the first case $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$ while in the second and third case $\delta=\gamma|\bar{I^{c}}|$.
Figure 7: $\delta$ for three different move-wait conflicts, with $\gamma=0.6$. In the top cases, $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$, in the bottom two cases $\delta=\gamma(\bar{t_{2}^{c}}-\bar{t_{1}^{c}})$. Orange interval is the restriction on the move action, purple is the restriction on the wait action.
V-A Soundness
We now show that our branching rule is sound. The branching rule applies the same rule as in CCBS on move-move conflicts, which is shown in Section 4.2 of [20] to satisfy property (2) and is therefore sound. Hence, we refer the reader to there for more details. What remains is to prove the soundness of our branching rule on move-wait conflicts.
Theorem V.1 (Soundness of our branching rule).
Our branching rule is sound when applied to a move-wait conflict.
Proof.
Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ and intersection interval $\bar{I^{c}}$. For some $\delta<|\bar{I^{c}}|$, agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for $t\in[t^{i},t^{i}+\delta)$. The equivalent constraint set of forbidding constraints is $C_{i}=\left\{\langle i,m^{i},t\rangle\mid t\in[t^{i},t^{i}+\delta)\right\}$. As this is a shifting constraint, we know that a collision occurs with $j$ located at $v(w_{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. On the other hand, agent $j$ is forbidden from executing any wait action that requires occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, or any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. All of these forbidden timed actions require being located at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. Thus, a collision will occur if $i$ and $j$ each execute any of their forbidden actions. Let $C_{j}$ contain one forbidding constraint for each of $j$’s forbidden actions. Then, it holds for any $\langle c_{i},c_{j}\rangle\in\langle C_{i},C_{j}\rangle$ that a valid solution must satisfy at least $c_{i}$ or $c_{j}$. Therefore, $\langle C_{i},C_{j}\rangle$ is sound and by Lemma III.1 so is our branching rule. ∎
To conclude, our branching rule is shown to be sound when applied to move-move conflicts (in [20]) and move-wait conflicts (in Theorem V.1). Therefore, the branching rule is sound.
V-B A Requirement for Non-termination
To prove that CCBS under our branching rule is solution complete and therefore terminates on a solvable MAPFR problem, we employ a proof by contradiction. We begin by assuming the opposite: that CCBS never terminates on some solvable MAPFR problem, despite the existence of a solution, and arrive at a condition which must hold if CCBS never terminates.
We will establish in this section that if CCBS runs indefinitely, then the CT must contain an infinite sequence of nodes that are continuously selected and branched upon. By analyzing this infinite sequence, we derive a specific requirement that any non-terminating execution must satisfy. It will later be show in Section V-C that CCBS under our branching rule does not satisfy this requirement, thus proving termination. Specifically, what we establish here is:
  1. V-B1) Infinite CT node sequence: There exists an infinite descending path in the constraint tree where each node contains conflicts and has objective value $\leq\sigma^{*}$, where $\sigma^{*}$ is the optimal.
  2. V-B2) Concentration on $\tau^{\infty}$: Infinitely many nodes in the sequence map to the same trajectory $\tau^{\infty}$.
  3. V-B3) Unbounded Constraint Accumulation: Constraints which determine how nodes map to $\tau^{\infty}$ accumulate unboundedly along the infinite node sequence, yet there must always exist another node mapping to $\tau^{\infty}$ and still maintaining an objective value lower than the optimal value.
The final point formulates the requirement on the branching rule for non-termination.
We begin by defining trajectories and how plans map to them.
Definition V.2 (Trajectory).
A trajectory $\tau$ is a finite sequence of move actions, $\tau=\langle m_{1},m_{2}\dots,m_{n}\rangle\in\mathcal{A}^{n}$.
Each move action $m_{i}\in\tau$ has a duration $m_{i,D}$. Thus, the duration of $\tau$ is $\tau_{D}=\sum_{k=1}^{n}m_{k,D}$.
Definition V.3 (Joint Trajectory).
A joint trajectory $\mathcal{T}$ is a set containing one trajectory $\tau_{i}$ for each agent $i$.
Definition V.4 (Plan-Trajectory Mapping).
A plan $\pi$ maps to a trajectory $\tau$, denoted $\pi\sim\tau$, if $\tau$ contains exactly every move action in $\pi$ in the order that they appear. Similarly, $\Pi\sim\mathcal{T}$ if it holds for every agent $i$ that $\pi_{i}\sim\tau_{i}$ with $\pi_{i}\in\Pi,\tau_{i}\in\mathcal{T}$.
Informally, a trajectory $\tau$ describes the sequence of moves an agent takes when following a plan $\pi\sim\tau$, but without timing details. Multiple plans can therefore map to the same trajectory, as plans can contain various wait actions. For example, the following two plans
  • $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ from Figure 2, and
  • $\pi_{i}^{\prime}=\langle\langle m_{1}^{i},0\rangle,\langle m_{3}^{i},1.5\rangle,\langle m_{4}^{i},2.5\rangle\rangle$
both map to the same trajectory $\tau_{i}=\langle m_{1}^{i},m_{3}^{i},m_{4}^{i}\rangle$. Since each trajectory $\tau$ has a duration $\tau_{D}$, a joint trajectory $\mathcal{T}$ has an objective value $\sigma(\mathcal{T})$.
Observe that $\tau_{D}\leq\pi_{D}$ if $\pi\sim\tau$ since $\pi$ contains at least all the move actions in $\tau$ but possibly additional wait actions. From this, and the fact that $\sigma$ is strictly monotonically increasing with the maximum plan duration, if $\Pi\sim\mathcal{T}$ then
$\sigma\left(\mathcal{T}\right)\leq\sigma\left(\Pi\right).$ (8)
Finally, we observe that the set $\mathbb{T}$ of all possible joint trajectories $\mathcal{T}$ is countable. This follows from the trajectories $\tau$ being finite sequences from the countable set $\mathcal{A}$, meaning that the set of all $\tau$ is countable. Since the set of all $\tau$ is countable, and a joint trajectory $\mathcal{T}$ comprises a discrete number of trajectories (exactly one per agent), the set of all possible joint trajectories $\mathbb{T}$ is countable. We define the subset $\mathbb{T}^{c}\subset\mathbb{T}$ to contain all joint trajectories with an objective value at or below some fixed constant $c\in\Re$:
$\mathbb{T}^{c}=\left\{\mathcal{T}\in\mathbb{T}\mid\sigma\left(\mathcal{T}\right)\leq c\right\}.$ (9)
It holds that $\mathbb{T}^{c}$ is finite for any $c<\infty$.
V-B1 Infinite CT node sequence
Since our branching rule is sound (Section V-A), all solutions (including the optimal solution) is reachable from the CT root. Non-termination implies that at every iteration, a non-solution CT node $N$ is expanded instead of a node representing an optimal solution with objective value $\sigma^{*}$. Since the search is best-first, it must hold that $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$ otherwise it would not be selected for expansion. Additionally, $N_{\Pi}$ must contain a conflict, otherwise $N_{\Pi}$ would constitute a valid solution. For non-termination to persist indefinitely, such nodes must be available at every iteration. This leads to the following recursive requirement: at every iteration there must exist at least one node $N$ satisfying the above two conditions, with a child node also satisfying the above two conditions, otherwise eventually all nodes satisfying these properties will be depleted and the algorithm will terminate with an optimal solution. In summary, non-termination implies that at every iteration exists at least one CT node $N$ where
  1. (I)
    $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$,
  2. (II)
    $N_{\Pi}$ contains a conflict, and
  3. (III)
    a child node of NN satisfies (I) and (II).
From this it follows the existence of an infinite descending path
$\varsigma=\langle N^{1},N^{2},\dots\rangle$ (10)
in the CT, where for every $k=1,2,\dots$ it holds that $N^{k}$ satisfies (I)-(III) and $N^{k+1}$ is a child of $N^{k}$.
V-B2 Concentration on $\tau^{\infty}$
Consider a node $N^{k}\in\varsigma$. When $N^{k}$ is expanded, CCBS detects a conflict between two plans in $N^{k}_{\Pi}$ and applies the branching rule to create two child nodes. By definition of sequence $\varsigma$, exactly one of these children continues the sequence as $N^{k+1}$. For one of the plans involved in the conflict, say $\pi_{i}\in N^{k}_{\Pi}$, a constraint is created for agent $i$ and added to the constraint set of $N^{k+1}$. In the context of our analysis, we say that $\pi_{i}$ is “branched on” at $N^{k}$.
Definition V.5 (Node-trajectory mapping).
A node $N\in\varsigma$, where $\pi$ is branched on, maps to the trajectory $\tau$ for which $\pi\sim\tau$.
With Definition V.5, we can map every one of the infinitely many nodes in $\varsigma$ to a specific trajectory. We now establish that infinitely many nodes in $\varsigma$ map to a specific trajectory, denoted $\tau^{\infty}$. Consider $N\in\varsigma$. The set of plans $N_{\Pi}$ satisfies $N_{\Pi}\sim\mathcal{T}$ for some $\mathcal{T}\in\mathbb{T}$, and $N$ maps to a trajectory $\tau\in\mathcal{T}$. By inequality (8), $\sigma\left(\mathcal{T}\right)\leq\sigma\left(N_{\Pi}\right)$, and by condition (I), $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$. Therefore, it follows that $\sigma\left(\mathcal{T}\right)\leq\sigma^{*}$, implying $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$. Since $\mathbb{T}^{\sigma^{*}}$ is finite and each $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$ contains a finite number of trajectories, the set of all possible trajectories that nodes in $\varsigma$ map to is finite. However, $\varsigma$ contains infinitely many nodes. By the Pigeonhole principle, infinitely many nodes in $\varsigma$ must map to at least one specific trajectory $\tau^{\infty}$.
V-B3 Unbounded Constraint Accumulation
For every branching at a node $N^{k}\in\varsigma$, a constraint is created and propagated to all subsequent nodes $N^{h}\in\varsigma$ for $h>k$. Since infinitely many nodes in $\varsigma$ map to $\tau^{\infty}$, constraints from branching on plans $\pi\sim\tau^{\infty}$ accumulate unboundedly, yet there must always exist some $\pi^{\prime}\sim\tau^{\infty}$ permitted under these constraints. On the other hand, by condition (I), every $N^{k}\in\varsigma$ must satisfy $\sigma\left(N^{k}_{\Pi}\right)\leq\sigma^{*}$. Since $\sigma$ is strictly monotonically increasing with the maximum plan duration, there exists a fixed constant $c<\infty$ (determined by $\sigma^{*}$) such that all plans in $N^{k}_{\Pi}$ have duration $\leq c$. Therefore, it must hold for all $N^{k}\in\varsigma$ that $\forall\pi\in N^{k}_{\Pi}:\pi_{D}\leq c$, otherwise $\sigma\left(N^{k}_{\Pi}\right)$ grows unboundedly to eventually violate condition (I).
Finally, we have arrived at our requirement for non-termination: for some fixed $c<\infty$ and trajectory $\tau^{\infty}$, infinitely many CCBS branching steps on a plan $\pi\sim\tau^{\infty}$ must preserve the existence of at least one valid plan $\pi^{\prime}\sim\tau^{\infty}$ with duration $\pi_{D}\leq c$.
V-C Solution Completeness
In this section, we prove that CCBS under our branching rule terminates within finite iterations by showing that the non-termination requirement established in Section V-B cannot be satisfied. This completes our proof by contradiction.
We prove that CCBS under our branching rule does not satisfy the non-termination requirement by showing that each branching operation reduces the set of feasible execution times for actions in $\tau^{\infty}$ by non-degenerate amounts, eventually making it impossible to construct any plan $\pi\sim\tau^{\infty}$ with a bounded duration.
The proof is structured in five stages:
  1. V-C1) Permitted Execution Times: We formalize how constraints restrict move execution and vertex occupancy times, establishing the framework for analyzing constraint accumulation.
  2. V-C2) CSIPP Planning Behavior: We explain how CCBS’s underlying path planner (CSIPP) constructs plans, focusing on how it selects execution times from feasible intervals.
  3. V-C3) Non-degenerate Reductions For Move Actions: We show how our branching rule applied to a timed move action reduces feasible move execution times by a non-degenerate amount.
  4. V-C4) Non-degenerate Reductions For Wait Actions: We show how our branching rule applied to a timed wait action reduces feasible move execution or feasible vertex occupancy times by a non-degenerate amount.
  5. V-C5) Bounded Plan Impossibility: Using the results from Sections V-C3 and V-C4, and the fact that all timed actions in a bounded-duration plan must end before a fixed time, we show that the feasible move execution and vertex occupancy times are exhausted after finite iterations of the CCBS search. Therefore, we show that CCBS under our branching rule does not satisfy the non-termination requirement and therefore is guaranteed to terminate.
V-C1 Permitted Execution Times
To prove that the non-termination requirement cannot be satisfied, we establish a framework to analyze how constraints accumulate and restrict agent movements. At the CT root $N^{R}$ where the constraint set is empty, $N^{R}_{C}=\varnothing$, all agents are permitted to execute all actions at any time $t\geq 0$. However, at nodes further down in the CT with non-empty constraint sets, any number of constraints may restrict when agents can execute moves and occupy vertices. Thus, under the constraints $N_{C}$ at CT node $N$, we define for a specific agent $i$
  • $\mathcal{T}_{m}\subseteq[0,\infty)$ as the set of all times when $i$ is permitted to execute a move action $m\in\mathcal{A}$, and
  • $\mathcal{T}_{v}\subseteq[0,\infty)$ as the set of all times when $i$ is permitted to occupy a vertex $v\in\mathcal{V}$.
Since constraints under our branching rule restrict move execution and vertex occupancy over time intervals (Definition V.1), every set $\mathcal{T}_{x}$ where $x$ ranges over all moves $m\in\mathcal{A}$ and vertices $v\in\mathcal{V}$ can be described as the union of a finite number maximally connected time intervals:
$\mathcal{T}_{x}=\bigcup_{h=1}^{N_{x}}S_{x}^{h},\quad S_{x}^{h}=[s_{x}^{h},e_{x}^{h})$ (11)
We refer to $S_{x}^{h}$ as a safe interval for $i$ to execute a move (when $x=m$) or occupy a vertex (when $x=v$).
V-C2 CSIPP Planning Behavior
Understanding how CSIPP selects execution times is crucial for proving that constraint additions cause non-degenerate reductions in feasible time intervals. The key insight is that CSIPP, like SIPP [29] which CSIPP is based on, systematically chooses the earliest possible time within a safe interval for each action. This property prevents CSIPP from making arbitrarily small timing adjustments when constraints are added and thereby postponing termination indefinitely.
CSIPP performs an A [30] search in the vertex-safe interval space. Each search state is represented by a tuple $(v,S_{v})$ where agent $i$ is located at $v\in\mathcal{V}$ during the safe interval $S_{v}=[s_{v},e_{v})\subseteq\mathcal{T}_{v}$. Let $t^{a}\in S_{v}$ be the time when $i$ arrives $v$. Unless state $(v,S_{v})$ is a goal state and $e_{v}=\infty$, agent $i$ must execute a move action to leave $v$ before the end of the safe interval at $e_{v}$. For each move action $m$ with $\mathit{from}(m)=v$, agent $i$ can execute $m$ at any time $t$ satisfying
  • $t\in[t^{a},e_{v})$ (after arriving and before it is unsafe),
  • $t\in\mathcal{T}_{m}$ (it is safe to execute the move action), and
  • $t+m_{D}\in\mathcal{T}_{v^{\prime}}$ (it is safe to occupancy $v^{\prime}$),
where $v^{\prime}=\mathit{to}(m)$. In other words, $t\in P_{m}$ where
$P_{m}=[t^{a},e_{v})\cap\mathcal{T}_{m}\cap\mathcal{T}_{v^{\prime}}^{m_{D}}$ (12)
and $\mathcal{T}_{v^{\prime}}^{m_{D}}=\{t-m_{D}\mid t\in\mathcal{T}_{v^{\prime}}\}$ to compensate for the arrival at $v^{\prime}$ being $m_{D}$ after $t$. By (11), $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$ are unions of maximally connected intervals, such that $P$ is also a union of maximally connected intervals. Fig. 8 illustrates how $P$ is formed from the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
Figure 8: Illustration of the construction of $P$ as the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
For every maximally connected interval $s_{m},e_{m})\subseteq P_{m}$, an edge representing $i$ executing $\langle m,s_{m}\rangle$ is created, leading to the next state $(v^{\prime},S_{v^{\prime}})$ where $S_{v^{\prime}}\subseteq\mathcal{T}_{v^{\prime}}$ is the safe interval at $v^{\prime}$ such that $s_{m}+m_{D}\in S_{v^{\prime}}$.
CSIPP always selects the earliest time $s_{m}$ for $i$ to execute $m$. This earliest-time selection property is essential for our proof: since the earliest time $s_{m}$ in $s_{m},e_{m})$ is selected, it is not possible to select an earlier time $t^{\prime}<s_{m}$ arbitrarily close to $s_{m}$ for the move action $m$ to be executed.
V-C3 Non-degenerate Reductions For Move Actions
When our branching rule is applied to a move action $\langle m^{i},t^{i}\rangle\in\pi_{i}$ in a move-move or move-wait conflict, a constraint is created which forbids the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$. We assume that $\pi_{i}$ was returned from CSIPP, such that a safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t_{i}$ was identified during the search. Consequently, the new constraint removes either a prefix or the entirety of $s_{m},e_{m})$ from $P_{m}$.
Starting with a move-move conflict, our branching rule forbids the execution of $m^{i}$ during the unsafe interval $t^{i},t^{i}_{u})$, which is non-degenerate by construction.
Consider instead a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$. Our branching rule creates a constraint forbidding the execution of $m^{i}$ during $t^{i},t^{i}+\delta)$. If $\delta$ is bounded away from 0, then $t^{i},t^{i}+\delta)$ is non-degenerate. By (7), $\delta=\min\left(\gamma|\bar{I^{c}}|,t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$ is the minimum of two positive quantities:
  • $\delta=\gamma|\bar{I^{c}}|$: It holds that $\delta$ is bounded away from zero since $\bar{I^{c}}$ is non-degenerate and $\gamma\in(0,1)$ is a fixed, non-zero value.
  • $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$: From (5) we know that $I^{c}\subseteq\bar{I^{c}}$ and $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$. The former implies that $\bar{t_{1}^{c}}\leq t_{1}^{c}<t_{2}^{c}\leq\bar{t_{2}^{c}}$ where strict inequality holds since $I^{c}$ is non-degenerate. The latter implies that $t^{j}\leq t_{1}^{c}<t_{2}^{c}<{t_{2}^{c}}^{\prime}\leq t^{j}+t^{j}_{D}$ where ${t_{2}^{c}}^{\prime}>t_{2}^{c}$ is arbitrarily close to $t_{2}^{c}$ (to correct for $I^{c}=[t_{1}^{c},t_{2}^{c})$ not including $t_{2}^{c}$ while $t^{j},t^{j}+t^{j}_{D}]$ does include $t^{j}+t^{j}_{D}$). These together give
    $\bar{t_{1}^{c}}$ $\leq t_{1}^{c}<t_{2}^{c}<{t_{2}^{c}}^{\prime}\leq t^{j}+w^{j}_{D},$
    $\bar{t_{1}^{c}}$ $<t^{j}+w^{j}_{D},$
    $0$ $<t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}=\delta,$
    such that $\delta$ is bounded away from 0.
Hence, $\delta$ is bounded away from zero, implying that the constrained interval $t^{i},t^{i}+\delta)$ is non-degenerate.
In both cases, where $\langle m^{i},t^{i}\rangle\in\pi_{i}$ is contained in a move-move or move-wait conflict, a constraint is created forbidding the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$. That is, $t^{i},t^{\prime})$ is removed from $\mathcal{T}_{m}$. Since $s_{m}=t^{i}$ for some $s_{m},e_{m})\subseteq P_{m}$, by (12), either the entire safe interval $s_{m},e_{m})$ is removed (if $t^{\prime}\geq e_{m}$) or a non-degenerate prefix $s_{m},t^{\prime})$ is removed from $P_{m}$ (if $t^{\prime}<e_{m}$).
V-C4 Non-degenerate Reductions For Wait Actions
In this section, we consider the constraints created when applying our branching rule to a wait action $\langle w^{j},t^{j}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$. It will be shown that this leads to two cases. In the first case, we show that there exists a timed move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ following $\langle w^{j},t^{j}\rangle$, and a constraint is created that forbids the execution of $m$ during a non-degenerate interval $t^{j}+w^{j}_{D},t^{\prime})$. Therefore, the same conclusions as in Section V-C3 are drawn. In the second case, a constraint is created forbidding the use of $v(w^{j})$ during a non-degenerate interval $t^{\prime},t^{\prime\prime})$, necessarily leading to a non-degenerate reduction in $\mathcal{T}_{v(w^{j})}$.
When our branching rule is applied to $\langle w^{j},t^{j}\rangle$, two constraints are created (Definition V.1). One constraint forbids $j$ from occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, and the other forbids the execution of any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$. We begin by showing that the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate by expanding $\bar{t_{1}^{c}}+\delta$ using (7):
$\bar{t_{1}^{c}}+\delta$ $=\bar{t_{1}^{c}}+\min\left(\gamma|\bar{I^{c}}|,\;t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$
$=\min\left(\bar{t_{1}^{c}}+\gamma|\bar{I^{c}}|,\;\bar{t_{1}^{c}}+t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$
$=\min\left(\bar{t_{1}^{c}}+\gamma(\bar{t_{2}^{c}}-\bar{t_{1}^{c}}),\;t^{j}+w^{j}_{D}\right)$
$=\min\left((1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}},\;t^{j}+w^{j}_{D}\right)$ (13)
which shows that $\bar{t_{1}^{c}}+\delta\leq(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}$. Since $0<\gamma<1$ is fixed to some value less than $1$, it follows that $(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}<\bar{t_{2}^{c}}$ and therefore that $\bar{t_{1}^{c}}+\delta<\bar{t_{2}^{c}}$. Thus, $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate.
Next, we show that the constrained $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ necessarily overlap with the time interval $t^{j},t^{j}+w^{j}_{D}]$ when $j$ occupies vertex $v(w^{j})$ while executing $\langle w^{j},t^{j}\rangle$. To do so, we combine two conditions. First, under the assumption made in Section III-B that $I^{c}$ is non-degenerate together with $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ by (5), it follows that $t^{j}<\bar{t_{2}^{c}}$ meaning that $j$’s occupancy of $v(w^{j})$ starts before the constrained interval ends. Second, (13) shows that $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$, implying that $j$’s occupancy of $v(w^{j})$ ends at or after the constrained interval begins. Together, these conditions are sufficient to guarantee overlap by at least a singular value.
There exists two cases: the overlap between $t^{j},t^{j}+w^{j}_{D}]$ and $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is either singular or non-degenerate.
  • Singular: this implies that $\bar{t_{1}^{c}}+\delta=t^{j}+w^{j}_{D}$. In this case, $w^{j}$ cannot be the final infinite wait action in $j$’s plan: $\bar{t_{1}^{c}}<\infty$ (since collisions occur at finite times) and $\delta<\infty$ (since $|\bar{I^{c}}|<\infty$, see (7)). Thus, the left-hand-side is $\infty$ implying that the right-hand side containing $w^{j}_{D}$ is also $\infty$. Since $w^{j}$ does not have infinite duration and is therefore not the final infinite wait action, there is a subsequent move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ with $\mathit{from}(m)=v(w^{j})$ in $j$’s plan. Our branching rule creates a constraint forbidding $m$ from being executed during the non-degenerate interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w^{j}_{D},\bar{t_{2}^{c}})$. The same reasoning and conclusions as in Section V-C3 for move actions in move-move and move-wait conflicts applies: a non-degenerate prefix or the entirety of some safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t^{j}+w^{j}_{D}$ identified during the CSIPP search is removed from $P_{m}$.
  • Non-degenerate: we know that $t^{j},t^{j}+w^{j}_{D}]\subseteq\mathcal{T}_{v(w^{j})}$ else the timed wait action $\langle w^{j},t^{j}\rangle$ would not have been returned by CSIPP. This means that the non-degenerate overlap $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\cap[t_{j},t^{j}+w^{j}_{D}]$ is also a subset of $\mathcal{T}_{v(w^{j})}$. In other words, this tells us that a non-degenerate part of $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ exists in $\mathcal{T}_{v(w^{j})}$. Our branching rule creates a constraint forbidding the occupancy of $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, therefore removing $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ from $\mathcal{T}_{v(w^{j})}$. Thus, a non-degenerate interval is removed from $\mathcal{T}_{v(w^{j})}$.
V-C5 Bounded Plan Impossibility
Consider a CT node $N\in\varsigma$ mapping to $\tau^{\infty}$. At $N$, a plan $\pi\in N_{\Pi}$ with $\pi\sim\tau^{\infty}$ is branched on. To satisfy the non-termination condition (I), $\pi_{D}\leq c$ for some fixed $c<\infty$ which requires that every timed action in $\pi$ (except the final infinite wait action, implicitly omitted in the following) ends before $c$. For simplicity, we suffice with the less restrictive requirement that all timed move actions must begin before $c$. In other words, under the constraints $N_{C}$ and $\pi_{D}\leq c$ it holds that
$\forall m\in\tau^{\infty}:P_{m}\subseteq[0,c)$ (14)
and
$\forall v\in\mathbf{v}:\mathcal{T}_{v}\subseteq[0,c)$ (15)
where $\mathbf{v}=\bigcup_{m\in\tau^{\infty}}\left\{\mathit{from}(m),\mathit{to}(m)\right\}$ contains all vertices visited along $\tau^{\infty}$.
In Section V-C3, we showed that when our branching rule is applied to a move action $\langle m,t\rangle\in\pi$, then either a non-degenerate prefix or the entirety of a maximally connected interval $s_{m},e_{m})\subseteq P_{m}$ is removed from $P_{m}$. Similarly, it was shown in Section V-C4 that when our branching rule is applied to a wait action $\langle w,t\rangle\in\pi$, then one of two cases occur. In the first case, the same constraining as when our branching rule is applied to a move action occurs for the move action following $\langle w,t\rangle$ in $\pi$. In the second case, a non-degenerate reduction of $\mathcal{T}_{v(w)}$ occurs.
Before proceeding, we analyze the one case which has yet to be addressed: when the action $\langle w_{n},t_{n}\rangle$ in $\pi\sim\tau^{\infty}$ that is branched on at $N$ is an infinite wait action, $w_{n,D}=\infty$. When our branching rule is applied to $\langle w_{n},t_{n}\rangle$, a constraint is created forbidding the use of $v(w_{n})$ during some interval $t^{\prime},t^{\prime\prime})$. From Section V-C4 we know that $t_{n}<t^{\prime\prime}$ by some non-degenerate amount. Thus, the earliest time the agent can arrive at $v(w_{n})$ to remain there indefinitely is at the later time $t^{\prime\prime}$ than $t_{n}$. Consequently, the timed move action $\langle m_{n-1},t_{n-1}\rangle\in\pi$ preceding $\langle w_{n},t_{n}\rangle$ is no longer valid since it leads to the agent arriving at $v(w_{n})$ before $t^{\prime\prime}$. Instead, $m_{n-1}$ can only occur at some time later than $t^{\prime\prime}-m_{n,D}$. Since $m_{n-1}$ is the last move action in $\tau^{\infty}$, these constraint additions lead to a non-degenerate increase in the duration of all plans $\pi^{\prime}\sim\tau^{\infty}$.
From the above, at every branching there occurs a non-degenerate reduction in one of the finite sets $P_{m}$ or $\mathcal{T}_{v}$ (with $m\in\tau^{\infty}$ and $v\in\mathbf{v}$) or a non-degenerate increase the in duration of all valid plans $\pi^{\prime}\sim\tau^{\infty}$. Therefore, after a finite many branchings either one of the sets will be depleted or the duration of valid plans mapping to $\tau^{\infty}$ will grow beyond the fixed $c<\infty$. In either case, eventually after a finite many branchings no plan $\pi\sim\tau^{\infty}$ satisfying the accumulated constraints $N_{C}$ and $\pi_{D}<c$ will exist.
To conclude, the requirement of non-termination which was established in Section V-B cannot be satisfied by CCBS under our branching rule. Therefore, termination is guaranteed after a finite number of iterations.
V-D Soundness and Solution Completeness
Theorem V.2 (Sound and solution complete).
CCBS under our branching rule is sound and solution complete.
Proof.
Our branching rule was shown in Section V-A to be sound. By Theorem III.2, CCBS is therefore sound when using a sound branching rule. Additionally, Section V-C proved that CCBS under our branching rule terminates within finite iterations. Since CCBS under our branching rule is both sound and guaranteed to terminate, Theorem III.3 establishes that it is solution complete. ∎
VI Experimental Evaluation
We solve the MAPFR problem in Figure 9 for minimal sum-of-costs using two variants of the publicly available CCBS implementation  [^3] https://github.com/PathPlanning/Continuous-CBS.: the unmodified CCBS release using the implemented branching rule, and an otherwise identical version using our branching rule  [^4] https://github.com/Adcombrink/S-and-SC-CCBS.. Experiments are run on a Mac Studio, 16-core M4 Max CPU, 64 GB RAM, macOS Sequoia (15.3). We use $\gamma=0.5$ when calculating $\delta$.
Figure 9: A MAPFR problem with four agents, each agent $i$ starts at vertex $S(i)$ and must reach its goal vertex $G(i)$.
The example in Figure 9 contains four agents, $i=1,..,4$, each with a start vertex $S(i)$ and goal vertex $G(i)$. All agents are circular with radius $r=\sqrt{2}/4$ and traverse edges at constant speed $1$ in a straight line. Agent $1$ starts at vertex $S(1)=E$ and must pass vertex $F$ to reach its goal $G(1)=G$, therefore, agent $2$ with $S(2)=G(2)=F$ must move out of the way. The solution found by both methods is for agent $2$ to move to $C$ to allow agent $1$ to pass. However, agent $2$ must avoid a collision with agent $3$ which passes vertex $C$ on its way to its goal vertex $D$. So, there are two options for the interaction between agents $2$ and $3$ at $C$:
  1. 1.
    agent $2$ waits for agent $3$, in turn meaning that agent $1$ must wait for agent $2$, or
  2. 2.
    agent $3$ waits for agent $2$, meaning that agent $4$ must wait for agent $3$.
Due to the difference in length of edges $EF$ and $AB$, the time that agent $4$ must wait for agent $3$ is larger than the time that agent $1$ must wait for agent $2$. Therefore, these two possible solutions have different sum-of-costs, only the one where agent $2$ waits for agent $3$ is optimal.
We now examine how CCBS under the implemented branching rule handles this problem. At the CT root $N^{R}$, with $N^{R}_{C}=\varnothing$, each agent’s initial plan is generated without collision avoidance:
$N^{R}_{\Pi}=\left\{\begin{split}\pi_{1}&=\langle\langle EF,0\rangle\langle FG,1.5\rangle\langle G,2.5,\infty\rangle\rangle\\ \pi_{2}&=\langle\langle F,0,\infty\rangle\rangle\\ \pi_{3}&=\langle\langle BC,0\rangle\langle CD,1\rangle\langle D,2,\infty\rangle\rangle\\ \pi_{4}&=\langle\langle AB,0\rangle\langle B,1,\infty\rangle\rangle\end{split}\right\}$
where $\langle VU,t\rangle$ denotes a timed move action along edge $(V,U)$ starting at time $t$, and $\langle V,t_{1},t_{2}\rangle$ is a timed wait action at vertex $V$ from $t_{1}$ to $t_{2}$. Two conflicts appear in $N^{R}_{\Pi}$: $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ and $\langle\langle FG,1.5\rangle,\langle F,0,\infty\rangle\rangle$ (between agents $1$ and $2$). CCBS branches on the first conflict, with collision interval $I^{c}=[0.793,1.5)$. This yields the constraints $\langle 1,EF,[0,\infty)\rangle$ (forbidding agent $1$ from ever traversing $EF$) and $\langle 2,F,I^{c}\rangle$ (forbidding agent $2$ from occupying $F$ during $I^{c}$). The first constraint is infeasible, as agent $1$ must traverse $EF$ to reach its goal $G$. Thus, agent $2$ must vacate $F$ before $t=0.793$. However, the earliest time that agent $2$ can traverse $FC$ while allowing agent $3$ to pass first is at $t=1$, which is not possible due to the constraint. Therefore, all valid solutions where agent $2$ waits for agent $3$ are removed from the search. Consequently, agent $3$ must wait for agent $2$ to complete its detour via $C$. The resulting solution is:
$\Pi=\left\{\begin{split}\pi_{1}&=\langle\langle EF,0\rangle\langle FG,1.5\rangle\langle G,2.5,\infty\rangle\rangle\\ \pi_{2}&=\langle\langle FC,0\rangle\langle C,1,1.5\rangle\langle CF,1.5\rangle\langle F,0,\infty\rangle\rangle\\ \pi_{3}&=\langle\langle B,0,1.5\rangle\langle BC,1.5\rangle\langle CD,2.5\rangle\langle D,3.5,\infty\rangle\rangle\\ \pi_{4}&=\langle\langle A,0,1.2\rangle\langle AB,1.2\rangle\langle B,2.2,\infty\rangle\rangle\end{split}\right\}.$
We now examine how CCBS under our branching rule handles the same conflict. The conflict $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ with $I^{c}=[0.793,1.5)$ has an intersection interval, i.e., the time during which agent $1$ executing $\langle EF,0\rangle$ is in collision with agent $2$ waiting at vertex $F$, of $\bar{I^{c}}=I^{c}=[0.793,1.5)$). This yields
$\delta=\min\left(0.5\,(1.5-0.793),\infty\right)=0.354$
and the following set of constraints:
  • Agent $1$: motion constraint $\langle 1,EF,[0,0.354)\rangle$,
  • Agent $2$: vertex constraint $\langle 2,F,[1.17,1.5)\rangle$ and motion constraints $\langle 2,FE,[1.17,1.5)\rangle$, $\langle 2,FC,[1.17,1.5)\rangle$, $\langle 2,FG,[1.17,1.5)\rangle$.
Unlike the constraints produced by standard CCBS, these restrictions exclude only infeasible behaviors while preserving feasible solutions — including the case where agent $2$ waits until after agent $3$ passes vertex $C$. The resulting solution is:
$\Pi=\left\{\begin{split}\pi_{1}&=\langle\langle R,0,0.5\rangle\langle EF,0.5\rangle\langle FG,2\rangle\langle G,2.5,\infty\rangle\rangle\\ \pi_{2}&=\langle\langle F,0,1\rangle\langle FC,1\rangle\langle CF,2\rangle\langle F,3,\infty\rangle\rangle\\ \pi_{3}&=\langle\langle BC,0\rangle\langle CD,1\rangle\langle D,2,\infty\rangle\rangle\\ \pi_{4}&=\langle\langle AB,0\rangle\langle B,1,\infty\rangle\rangle\end{split}\right\}$
where agent $2$ waits at $F$ until agent $3$ has traversed $C$.
The performance using each branching rule is reported in Table I. There, we see that our branching rule yields a solution with lower sum-of-costs, makespan, and computation time than standard CCBS. Consequently, this provides additional evidence to that in [28] that CCBS is not sound. Additionally, with the support of the theoretical results in Section V, the solution found by CCBS under our branching rule has optimal sum-of-costs.
TABLE I: Performance comparison between standard CCBS and CCBS using our branching rule.
Branching rule Sum-of-Costs Makespan Computation Time
Standard $10.707$ $3.500$ $8.1$ ms
Ours $9.000$ $3.000$ $1.4$ ms
VII Conclusions
This work set out to resolve the recently revealed long-standing gap between the theory and practice of Continuous-time Conflict Based Search (CCBS) for optimal Multi-Agent Path Finding in continuous time (MAPFR). We revisited the MAPFR formulation and analyzed the two existing branching schemes — the theoretical rule and the publicly available implementation — through a new analytical framework. This analysis yielded simple, sufficient conditions for soundness and solution completeness and showed that the implemented branching rule violates them, thereby supporting the counter-examples that cast doubt on CCBS.
We then presented a our novel branching rule and proved that it meets all conditions: it never excludes solutions from the search and guarantees termination when used in CCBS. Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of discrete-time CBS for the first time in the continuous domain. We validated these results by presenting an example MAPFR problem and applying standard CCBS and CCBS using our branching rule. Results shows that our branching rule produces a solution with lower sum-of-costs than standard CCBS, thereby showing definitively that standard CCBS is not sound.
Practically, our rule can be adopted as a drop-in replacement in existing CCBS code-bases to obtain certificates of optimality without risking non-termination. Conceptually, the constraint-set perspective and non-termination criterion offer tools for analyzing other CCBS-like MAPFR solvers and extensions, laying a solid foundation for the next generation of multi-agent path finding algorithms.

### v2

@@
-Preparation of Papers for IEEE Sponsored Conferences & Symposia*
+Optimal Multi-agent Path Finding in Continuous Time
 Abstract
-Continuous-time Conflict Based-Search (CCBS) has long been viewed as the de-facto optimal solver for multi-agent path finding in continuous time (MAPFR).
-Recent findings, however, show that the original theoretical variant of CCBS can suffer from non-termination, while the widely used implementation can return sub-optimal solutions.
-We introduce an analytical framework that yields simple and sufficient conditions under which any CCBS-style algorithm is both sound, i.e., returns only optimal solutions, and solution complete, i.e., terminates on every solvable MAPFR instance.
-Investigating the publicly available implementation of CCBS reveals that it violates these conditions.
-Though this merely indicates that CCBS might be unsound, this indication is supported by counter-examples.
-Leveraging the analytical framework, we propose a novel branching rule and prove that it satisfies the sufficient conditions, thereby restoring soundness and termination guarantees.
-Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of the discrete-time CBS for the first time in the continuous domain.
-We experimentally apply standard CCBS and CCBS under our branching rule to an example problem, with our branching rule returning a solution with lower sum-of-costs than standard CCBS.
-Because the branching rule largely only affects the branching step, it can be adopted as a drop-in replacement in existing code-bases, as we show in our provided implementation.
-Beyond CCBS, the analytical framework and termination criterion provide a systematic way to evaluate other CCBS-like MAPFR solvers and future extensions.
-I INTRODUCTION
+Continuous-time Conflict Based-Search (CCBS) has long been viewed as the standard optimal baseline for multi-agent path finding in continuous time (MAPFR), yet recent critiques show that the theoretically described CCBS can fail to terminate on solvable MAPFR problems while the publicly available reference implementation can return sub-optimal solutions.
+This work presents an analytical framework that yields simple and sufficient conditions under which any CCBS-style algorithm is both sound (returns only optimal solutions) and solution complete (terminates on every solvable MAPFR problem).
+Investigating the reference CCBS implementation reveals that it violates our sufficient conditions for soundness, with counterexamples demonstrating sub-optimality.
+Leveraging the framework, we introduce a branching rule ($\delta$-BR) and prove it restores soundness and termination guarantees.
+Consequently, the resulting CCBS variant is both sound and solution complete.
+To our knowledge, this is the first MAPFR solver matching the guarantees of the discrete-time CBS.
+On a constructed example, CCBS with $\delta$-BR improves sum-of-costs from $10.707$ to $9.000$ ($\approx 16\%$ lower) compared to the reference CCBS implementation.
+Across benchmarks, the reference CCBS implementation is generally able to find solutions faster than CCBS with $\delta$-BR due to its more aggressive pruning.
+However, this comes at the cost of occasional sub-optimality and potential non-termination when all solutions are pruned, whereas $\delta$-BR preserves optimality and guarantees termination by design.
+Because $\delta$-BR largely only affects the branching step, it can be adopted as a drop-in replacement in existing codebases, as we show in our provided implementation.
+Beyond CCBS, the analytical framework and termination criterion provide a systematic way to evaluate other CCBS-like MAPFR solvers and future extensions, thereby offering tools for rigorous analysis of next-generation MAPFR algorithms.
+keywords:
+Multi-agent Path Finding , Continuous-time Conflict-based Search , Algorithm soundness and solution completeness , Optimal Path planning
+journal: Artificial Intelligence
+[…]
+[chalmers]organization=Division of Systems and Control, Department of Electrical Engineering, Chalmers University of Technology, city=Göteborg, country=Sweden
+1 Introduction
 As robotics and automation take an increasingly larger role in society, there is a growing need to move agents from where they are to where they need to be without collision.
-This is the problem of Multi-agent Path Finding (MAPF) which finds application across diverse domains, from computer games [1] to warehouses [2, 3], factories [4] to airports [5].
+This is the problem of Multi-agent Path Finding (MAPF), which finds application across diverse domains, from computer games [1] to warehouses [2, 3], factories [4] to airports [5].
 Progress toward fast and exact solution methods has been steady for decades, targeting the many variants of MAPF.
@@
 Agents are represented as points and can wait at vertices and traverse edges, with time being discrete, edges taking unit-time to traverse, and an objective function based on arrival times at goal vertices being minimized.
-This problem is NP-hard [6] and has many variations, comprehensively detailed in e.g.
-[7, 8, 9].
+This problem is NP-hard [6] and has many variations, comprehensively detailed in, e.g., [7, 8, 9].
 Early focus was placed on reducing computation and searching for sub-optimal solutions by, e.g., establishing a priority order among agents or restricting the search to a limited time-window [10, 11].
 Shortly thereafter, exact methods such as $M^{*}$ [12], Increasing Cost Tree Search [13] and Conflict Based-Search (CBS) [14] were introduced.
-These methods cleverly avoid searching the entire solution space — which grows exponentially with the number of agents — by adopting a “lazy” approach of planning agents individually and addressing conflicts only as they occur.
-Of the three, CBS has enjoyed the largest adoption, with multiple continuations.
-To name a few, Meta-Agent CBS [15] and Improved CBS [16] for performance enhancements, Enhanced CBS [17] for bounded sub-optimal solutions, Rolling Horizon Collision Resolution [18] for lifelong MAPF, and Multi-Goal CBS [19] for MAPF where agents are assigned multiple goal vertices.
-The simplifying assumptions in MAPF that time is discretized and edges take unit-time to traverse bring significant drawbacks to its practical applicability; in real-world settings, environments do not always follow grid-like structures, agent speeds and action times vary, and artificial wait times to ensure that agents move in lockstep inflates the objective value.
+These methods cleverly avoid searching the entire solution space — which grows exponentially with the number of agents — by adopting a “lazy” approach of planning agents individually and addressing collisions only as they occur.
+Of the three, CBS has enjoyed the largest adoption with multiple continuations.
+To name a few, Meta-Agent CBS [15] and Improved CBS [16] for performance enhancements, Enhanced CBS [17] and CBS-Budget [18] for bounded sub-optimal solutions, Rolling Horizon Collision Resolution [19] for lifelong MAPF, and Multi-Goal CBS [20] for MAPF where agents are assigned multiple goal vertices.
+The simplifying assumptions of discrete time and unit-length edges in MAPF bring significant drawbacks to its practical applicability; in real-world settings, environments do not always follow grid-like structures, agent speeds and action times vary, and artificial wait times to ensure that agents move in lockstep inflate the objective value.
 The continuous-time MAPF variant, MAPFR [21], removes such discrete-time assumptions.
-In MAPFR, agents can traverse edges at any real-valued time by following a motion function that defines a continuous trajectory between two connected vertices.
+In MAPFR, agents can traverse edges at any real-valued time by following one of potentially several motion functions that define a continuous trajectory between two connected vertices.
 Additionally, agents have volumes such that two agents collide when their volumes overlap.
 The work in [22] addresses MAPFR for sub-optimal solutions, while [23] finds optimal solutions to a simplified MAPFR problem where agents traverse edges in straight lines at constant speeds.
-Currently, the only known method to address MAPFR in its entirety for optimal solutions is Continuous-time CBS (CCBS) [20] which claims to be sound and solution complete, that is, it guarantees to return an optimal solution if one exists.
+To our knowledge, the only known method to address MAPFR in its entirety for optimal solutions is Continuous-time CBS (CCBS) [21] which claims to be sound (only returns optimal solutions) and solution complete (guaranteed to terminate on any solvable MAPFR problem).
 Since the introduction of CCBS, numerous continuation works have been published, such as [24] for performance enhancements, $T$-Robust CCBS [25] for solutions that are robust to delays, [26] for Any-Angle MAPF, Continuous-time Prioritize Lifelong Planner [27] for lifelong MAPF, and [28] where CCBS is modified and applied to quadcopter drones.
-Despite CCBS’s wide-spread adoption, only recently has a critical assessment in [28] of the claims made in [20] established that CCBS does not fulfill its guarantees.
+Despite CCBS’s wide-spread adoption, only recently has a critical assessment in [29] established that CCBS does not fulfill its guarantees.
 The theoretical description of CCBS does not guarantee termination under the existence of an optimal solution (a condition for solution completeness) and the publicly available implementation does not guarantee that a returned solution is optimal (a condition for soundness).
 This is shown in [29] theoretically and with counter-examples.
-Therefore, these results imply that currently there does not exist a MAPFR solver for optimal solutions.
+Therefore, these results reveal a significant gap between the theory and practice of CCBS and any continuation works that rely on the soundness and solution complete guarantees.
 In this work, we develop an analytical framework to aid the analysis of CCBS and find sufficient but not necessary requirements for CCBS to be sound and solution complete.
-We then show that the branching rule in CCBS does not satisfy this requirement, which by itself does not mean that CCBS is not sound and solution complete, but it does complement the results in [28] and our experiments.
-Using the analytical framework, we then present a novel branching rule and prove that CCBS is sound and solution complete when using it.
-With this work, we therefore provide the first provably sound and solution-complete method to the MAPFR problem.
-The outline of this article is the following: In Section II, we formally define the MAPFR problem, soundness, and solution completeness.
+We then show that the publicly available CCBS implementation does not satisfy these requirements, which by itself does not mean that CCBS is not sound and solution complete, but it does complement the results in [29] and our experiments showing unsoundness.
+Leveraging this framework, we introduce a new branching rule ($\delta$-BR) and prove that CCBS using $\delta$-BR is sound and solution complete.
+That is, CCBS with $\delta$-BR is guaranteed to return an optimal solution within a finite number of iterations on any solvable MAPFR problem.
+Our only assumption on the objective function is that it is strictly monotonically increasing with respect to the maximum agent arrival time, thus, we guarantee optimality for sum-of-costs, makespan, and any other objective function with this property.
+Therefore, to our knowledge, we provide the first provably sound and solution-complete method to the MAPFR problem.
+With our provided implementation, we show that $\delta$-BR can be adopted as a drop-in replacement in existing CCBS implementations.
+These results are empirically validated on a constructed counterexample and benchmark problems, showing that the existing CCBS implementation is generally able to find solutions faster than CCBS with $\delta$-BR due to its aggressive search space pruning, however, occasionally returns sub-optimal solutions after removing all optimal solutions.
+Furthermore, the introduced framework and non-termination criteria provide analytical tools for reasoning about CCBS-like solvers and their extensions, thereby laying a foundation for the next generation of sound and solution complete MAPFR algorithms.
+The outline of this article is the following: In Section 2, we formally define the MAPFR problem, soundness, and solution completeness.
 CCBS is also introduced, along with details surrounding the findings in [29].
-The framework for analyzing branching rules and CCBS’s algorithmic properties is presented in Section III.
-In Section IV, we apply the framework on CCBS’s implemented branching rule.
-Our branching rule is presented in Section V, along with soundness and solution-completeness proofs.
-Section VI presents experimental results when applying standard CCBS and CCBS using our branching rule to an example problem.
-Finally, conclusions can be found in VII.
-II Background
+The framework for analyzing branching rules and CCBS’s algorithmic properties is presented in Section 3.
+In Section 4, we apply the framework on CCBS’s implemented branching rule.
+Our branching rule is presented in Section 5, along with soundness and solution completeness proofs.
+Section 6 presents empirical results.
+Finally, conclusions are found in 7.
+2 Background
 In this section, we formally introduce the MAPFR problem and define what we mean by soundness and solution completeness.
 Thereafter, CCBS [21] and the recent findings on CCBS’s guarantees in [29] are presented.
-II-A Problem Formulation
+2.1 Problem Formulation
 A MAPFR problem is defined in [21] by a tuple $\langle\mathcal{G},\mathcal{M},S,G,\mathit{coord},\mathcal{A}\rangle$, where $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ represents a graph with vertices $\mathcal{V}$ and edges $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$, $\mathcal{M}$ denotes a 2D metric space, $S$ and $G$ respectively map agents to start and goal vertices, $\mathit{coord}:\mathcal{V}\rightarrow\mathcal{M}$ maps vertices to coordinates in $\mathcal{M}$, and $\mathcal{A}$ comprises a finite set of move actions.
@@
 Each action specifies a trajectory in $\mathcal{M}$ from $v$ to $v^{\prime}$, starting at time $t=0$ and ending after potentially different durations ($m_{1,D}$, $m_{2,D}$, $m_{3,D}$).
+Figure 2: An illustration of a valid plan $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ with action durations $m_{1,D}^{i}=1.5$, $w_{2,D}^{i}=0.5$, $m_{3,D}^{i}=1$, $m_{4,D}^{i}=1.3$, $w_{5,D}^{i}=\infty$.
 In addition to move actions, there are also wait actions not included in $\mathcal{A}$.
@@
 That is, the location in $\mathcal{M}$ of an agent executing $\langle a,t\rangle$ is $a_{\varphi}(t^{\prime}-t)$ during $t^{\prime}\in[t,t+a_{D}]$.
-In other words, an action specifies how an agent moves and the duration to do so (always starting at time 0), while a timed action additionally specifies when the movement begins.
+In other words, an action specifies how an agent moves and the duration to do so, while a timed action additionally specifies when the movement begins.
 An agent $i$ can execute multiple timed actions, one after the other, to thereby traverse the graph. A plan $\pi_{i}$ for an agent $i$ is a finite sequence of timed actions,
 $\pi_{i}=\langle\langle a_{1}^{i},t_{1}^{i}\rangle,\langle a_{2}^{i},t_{2}^{i}\rangle,\dots,\langle a_{n}^{i},t_{n}^{i}\rangle\rangle.$ (1)
-For a plan $\pi_{i}$ to be valid, it must start at time 0 ($t_{1}^{i}=0$), start and end at $i$’s start and goal vertices ($\mathit{from}(a_{1}^{i})=S(i)$ and $\mathit{to}(a_{n}^{i})=G(i)$), and every successive timed action must start when and where the previous ends ($\forall k=2,..,n$: $t_{k}^{i}=t_{k-1}^{i}+a_{k-1,D}^{i}$ and $\mathit{from}(a_{k}^{i})=\mathit{to}(a_{k-1}^{i})$).
+For a plan $\pi_{i}$ to be valid, it must start at time 0 ($t_{1}^{i}=0$), start and end at $i$’s start and goal vertices ($\mathit{from}(a_{1}^{i})=S(i)$ and $\mathit{to}(a_{n}^{i})=G(i)$), and every successive timed action must start when and where the previous ends,
+$\left(t_{k}^{i}=t_{k-1}^{i}+a_{k-1,D}^{i}\right)\wedge\left(\mathit{to}(a_{k-1}^{i})=\mathit{from}(a_{k}^{i})\right)\quad\forall k=2,..,n.$
 Additionally, the final action $a_{n}^{i}$ is an infinite wait action with duration $a_{n,D}^{i}=\infty$ to capture that agents always remain idle indefinitely at their last location.
@@
 The first timed move action is $\langle m_{1}^{i}=\langle m_{1,\varphi}^{i},1.5\rangle,0\rangle$, which specifies that agent $i$ begins at $t=0$ to move according to the motion function $m_{1,\varphi}^{i}$ which has duration $1.5$.
-This is followed by the timed wait action $\langle w_{2}^{i}=\langle w_{2,\varphi}^{i},0.5\rangle,1.5\rangle$ specifying that $i$ waits at $v_{2}$ from $t=1.5$ for a duration of $0.5$, that is until $t=2$.
+This is followed by the timed wait action $\langle w_{2}^{i}=\langle w_{2,\varphi}^{i},0.5\rangle,1.5\rangle$ specifying that $i$ waits at $v(w_{2}^{i})=v_{2}$ from $t=1.5$ for a duration of $0.5$, that is until $t=2$.
 Thereafter, $i$ executes $\langle m_{3}^{i}=\langle m_{3,\varphi}^{i},1\rangle,2\rangle$ to arrive at $v_{3}$ at $t=3$, and then executes $\langle m_{4}^{i}=\langle m_{4,\varphi}^{i},1.3\rangle,3\rangle$ to arrive at $v_{4}$ at $t=4.3$.
 Finally, the timed wait action $\langle w_{5}^{i}=\langle w_{5,\varphi}^{i},\infty\rangle,4.3\rangle$ specifies that $i$ waits at $v_{4}$ indefinitely from $t=4.3$.
-Naturally, the plan $\pi_{i}$ illustrated in the example represents merely one among countless valid plans; alternative wait actions could be employed at any of the vertices, different move actions could be selected for the same edges, or entirely different paths through the graph could be traversed.
-Figure 2: An illustration of a valid plan $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ with action durations $m_{1,D}^{i}=1.5$, $w_{2,D}^{i}=0.5$, $m_{3,D}^{i}=1$, $m_{4,D}^{i}=1.3$, $w_{5,D}^{i}=\infty$.
+Naturally, the plan $\pi_{i}$ illustrated in the example represents merely one among countless valid plans; alternative wait actions could be employed at any of the vertices, different move actions could be selected for the same edges, or an entirely different path through the graph could be traversed.
 Each agent occupies a non-zero volume; a collision occurs when the volumes of two agents overlap.
@@
 Note that a wait-wait conflict can only arise after a move-move or move-wait (assuming agents start in collision-free positions).
-Therefore, we sufficiently consider only move-move and move-wait conflicts.
+Therefore, we sufficiently consider only move-move and move-wait conflicts, as a wait-wait conflict cannot exist without one of the other conflict types.
 Two plans $\pi_{i}$ and $\pi_{j}$ conflict if there exists a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$.
-A joint plan $\Pi$ is a set containing one valid plan for each agent.
-A solution is a joint plan where no two plans conflict.
-An optimal solution $\Pi^{*}$ is a solution that minimizes an objective function $\sigma$ over all solutions.
-The sum-of-costs (sum over plan durations) and makespan (maximum over plan durations) are the two most common objective functions in the MAPF context.
-Here, we assume that $\sigma(\Pi)$ is strictly monotonically increasing with respect to the maximum duration over all agent plans, $\max_{\pi\in\Pi}\pi_{D}$, which both the sum-of-costs and makespan satisfy.
-II-B Algorithmic Properties
+A joint plan $\Pi$ is a set containing one valid plan for each agent, a solution is a joint plan that does not contain any two conflicting plans, and an optimal solution $\Pi^{*}$ is a solution that minimizes an objective function $\sigma$ over all solutions.
+The sum-of-costs $\sigma(\Pi)=\sum_{\pi\in\Pi}\pi_{D}$ (sum over plan durations) and makespan $\sigma(\Pi)=\max_{\pi\in\Pi}\pi_{D}$ (maximum over plan durations) are the two most common objective functions in the MAPF context.
+Here, we assume that $\sigma(\Pi)$ is strictly monotonically increasing with respect to the maximum duration over all agent plans in $\Pi$, $\max_{\pi\in\Pi}\pi_{D}$, which both the sum-of-costs and makespan satisfy.
+Therefore, all proofs and guarantees provided in this work are valid for any objective function satisfying this assumption.
+2.2 Algorithmic Properties
 In this article, we say that a solution algorithm to the MAPFR problem is
@@
   • solution complete if it is guaranteed to terminate with a solution on any solvable MAPFR problem.
-In [20], a sound algorithm returns only conflict-free solutions, while an exact algorithm [^1] In [20], the term optimal is used synonymously with exact.
-However, an algorithm’s optimality typically refers to its time and space complexity and not to the quality of the solutions that it returns.
+In [21], a sound algorithm returns only solutions, while an exact algorithm [^1] In [21], the term optimal is used synonymously with exact.
+However, an algorithm’s optimality typically refers to its time and space complexity and not to the quality of the solutions that it produces.
 returns only optimal solutions.
@@
 Also note that completeness guarantees the return of a solution if one exists or the report of no solution otherwise, while the weaker solution completeness only guarantees the former.
-II-C Continuous-time Conflict Based Search
-CCBS is argued to be a sound and solution-complete MAPFR solver for minimal sum-of-costs [20]. CCBS closely follows CBS [14] by simultaneously building and searching a binary constraint tree (CT).
-To explain the CT and how it is searched, we begin by briefly introducing CCBS’s underlying planner Constrained Safe Interval Path Planning (CSIPP) which is covered in more detail in Section V-C2.
-CSIPP is a variant of SIPP [29] that computes a valid plan for a single agent with minimal duration and that satisfies a given set of constraints $C$.
-In SIPP, an $A^{*}$ [30] search is performed in the vertex-safe interval state space, with edges in $\mathcal{E}$ providing transitions from one state to the next.
-In CSIPP, it is instead move actions that provide transitions.
-Constraints in this context (discussed later in this section) constrain the execution of timed actions.
-Therefore, CSIPP finds the shortest-duration plan which does not include any timed actions forbidden by constraints in $C$.
+2.3 Continuous-time Conflict Based Search
+CCBS is argued to be a sound and solution-complete MAPFR solver for minimal sum-of-costs [21]. CCBS closely follows CBS by simultaneously building and searching a binary constraint tree (CT).
+To explain the CT and how it is searched, we begin by briefly introducing CCBS’s underlying planner Constrained Safe Interval Path Planning (CSIPP) which is covered in more detail in Section 5.3.2.
+CSIPP — a variant of Safe Interval Path Planning (SIPP) [30] — computes a minimal-duration valid plan for a single agent that satisfies a given set of constraints $C$.
+In SIPP, an $A^{*}$ [31] search is performed in the vertex-safe interval state space, with edges in $\mathcal{E}$ providing transitions from one state to the next, and safe intervals defining when an edge may be traversed or a vertex may be occupied.
+In CSIPP, constraints define the safe intervals and move actions provide transitions between states.
 In the CT, a node $N$ is associated with a set of constraints $N_{C}$ and a joint plan $N_{\Pi}$.
-The plans in $N_{\Pi}$ are computed using CSIPP given the constraint set $N_{C}$.
-Therefore, every plan $\pi_{i}\in N_{\Pi}$ is valid and has minimal duration under the constraints in $N_{C}$.
-The CCBS search starts at the CT root node $N^{R}$, with constraint set $N^{R}_{C}=\varnothing$.
-Since $N^{R}_{C}$ is empty, every plan $\pi_{i}\in N^{R}_{\Pi}$ is unconstrained (apart from being valid) and therefore has the shortest possible duration over all valid plans for the corresponding agent $i$.
-At every iteration of CCBS, a CT node $N$ minimizing $\sigma(N_{\Pi})$ over all unexpanded nodes is selected for expansion.
-If the joint plan $N_{\Pi}$ is a solution (i.e., conflict-free), then $N_{\Pi}$ is returned as an optimal solution.
+The plans in $N_{\Pi}$ are computed using CSIPP given the constraint set $N_{C}$, such that every plan $\pi_{i}\in N_{\Pi}$ is valid and has minimal duration under the constraints in $N_{C}$.
+The CCBS search starts at the CT root node $N^{R}$ with the empty constraint set $N^{R}_{C}=\varnothing$.
+Since $N^{R}_{C}$ is empty, every plan $\pi_{i}\in N^{R}_{\Pi}$ is unconstrained (apart from being valid) and therefore has the shortest possible duration over all valid plans for agent $i$.
+At every iteration of CCBS, a node $N$ minimizing $\sigma(N_{\Pi})$ over all unexpanded nodes is selected for expansion.
+If $N_{\Pi}$ is a solution (i.e., conflict-free), then $N_{\Pi}$ is returned as an optimal solution.
 Otherwise, a branching rule is applied to $N$ to spawn two child nodes.
@@
 These constraints $c_{i}$ and $c_{j}$ respectively constrain the plans of agents $i$ and $j$ to avoid this specific conflict.
-The child nodes, $N^{i}$ and $N^{j}$, each inherit one of the constraints in addition to all constraints at $N$: $N^{i}_{C}=N_{C}\cup\{c^{i}\}$ and $N^{j}_{C}=N_{C}\cup\{c^{j}\}$.
+The child nodes $N^{i}$ and $N^{j}$ each inherit one of the constraints in addition to all constraints at $N$: $N^{i}_{C}=N_{C}\cup\{c^{i}\}$ and $N^{j}_{C}=N_{C}\cup\{c^{j}\}$.
 The plans in $N_{\Pi}^{i}$ and $N_{\Pi}^{j}$ are then computed using CSIPP to satisfy $N_{C}^{i}$ and $N_{C}^{j}$, respectively.
@@
 Therefore, $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ is a conflict.
-Based on this conflict and a branching rule, a pair of constraints $\langle c^{i},c^{j}\rangle$ is created.
+Based on this conflict, a branching rule is used to create a pair of constraints $\langle c^{i},c^{j}\rangle$.
 We will explore these constraints in more detail later; for now it is sufficient to know that $c^{i}$ constrains when $i$ may execute $m^{i}_{2}$, and $c^{j}$ constrains when $j$ may execute $m^{j}_{3}$, to avoid this specific conflict.
@@
 Using CSIPP, the plans in $N^{i}_{\Pi}$ are computed to satisfy $N^{i}_{C}$.
-In practice, it is sufficient to only compute $i$’s plan and let all other agents inherit their plans from the parent node, since only $i$ has additional constraints.
+In practice, it is sufficient to only compute $i$’s plan since only the constraints on $i$ have changed, while all other agents can inherit their plans from $N$.
 Figure 3(c) illustrates $N^{i}_{\Pi}$, where agent $i$ now takes a different path through the graph.
@@
 The joint plan $N^{j}_{\Pi}$ is illustrated in Figure 3(d), where $j$ now waits for some time at a vertex before continuing.
-The resulting joint plans at the child nodes happen to be conflict-free and are therefore both solutions, in this particular example.
-In general, however, the child nodes’ joint plans may contain conflicts, requiring further CCBS iterations until a solution is found.
-(a) The CT node NN and its two children $N^{i}$ and $N^{j}$.
-(b) The joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$ contains a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$.
-(c) The joint plan $N_{\Pi}^{i}$ satisfying $N^{i}_{C}$.
-(d) The joint plan $N_{\Pi}^{j}$ satisfying $N^{j}_{C}$.
-Figure 3: An example of a branching rule applied to a CT node $N$.
-Since $N_{\Pi}$ contains a conflict, two child nodes $N^{i}$ and $N^{j}$ are spawned with additional constraints and recomputed plans.
-The soundness and solution-completeness proofs of CCBS [20] rely on the CT satisfying the following properties, where $N$ is a parent to $N^{i}$ and $N^{j}$ and $\mathcal{S}(N)$ is the set of all solutions satisfying $N_{C}$:
+In this particular example, the resulting joint plans at the child nodes happen to be conflict-free and are therefore both solutions.
+In general, however, the child nodes’ joint plans may still contain conflicts, requiring further CCBS iterations until a solution is found.
+(a) The CT node $N$ where a conflict is detected in $N_{\Pi}$, and its two children $N^{i}$ and $N^{j}$.
+(b) The joint plan $N_{\Pi}=\left\{\pi_{i},\pi_{j}\right\}$ contains a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ between agents $i$ and $j$.
+(c) The joint plan $N^{i}_{\Pi}$.
+Additional constraints on agent $i$ are added to $N^{i}_{C}$ to specifically avoid the collision detected at $N$.
+A new plan $\pi_{i}\in N^{i}_{\Pi}$ satisfying $N^{i}_{C}$ is then computed, where $i$ traverses another path in the graph.
+(d) The joint plan $N^{j}_{\Pi}$.
+Additional constraints on agent $j$ are added to $N^{j}_{C}$ to specifically avoid the collision detected at $N$.
+A new plan $\pi_{j}\in N^{j}_{\Pi}$ satisfying $N^{j}_{C}$ is then computed, where $j$ waits for some time at a node.
+Figure 3: An example of a branching rule applied to a conflict $\langle\langle m^{i}_{2},t^{i}_{2}\rangle,\langle m^{j}_{3},t^{j}_{3}\rangle\rangle$ at CT node $N$.
+For agent $i$ ($j$), a child node $N^{i}$ ($N^{j}$) is spawned with additional constraints on $i$ ($j$) to avoid this specific conflict.
+A new plan is then computed which satisfies these constraints.
+The soundness and solution completeness proofs of CCBS in [21] rely on the CT satisfying the following properties, where $N$ is a parent to $N^{i}$ and $N^{j}$ and $\mathcal{S}(N)$ is the set of all solutions satisfying $N_{C}$:
 $\mathcal{S}(N)$ $=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j}),$ (2)
@@
 If property (2) holds then $\mathcal{S}(N^{R})=\mathcal{S}(N^{i})\cup\mathcal{S}(N^{j})$ (with $N^{i}$ and $N^{j}$ being children of $N^{R}$) which recursively implies that all solutions are reachable from the CT root.
-Since the search starts at $N^{R}$, all solutions to a MAPFR problem can be found.
+Since the search starts at $N^{R}$, all solutions to a MAPFR problem are included in the search and therefore can be found.
 Since it is the branching rule that determines $\mathcal{S}(N^{i})$ and $\mathcal{S}(N^{j})$ (by the constraint pair that it constructs), it is the branching rule that determines if property (2) is satisfied.
-Property (3) is automatically satisfied by CSIPP.
-Since plans in $N_{\Pi}$ are computed using CSIPP given constraints $N_{C}$, every plan $\pi_{i}\in N_{\Pi}$ is the shortest-duration valid plan over all valid plans satisfying $N_{C}$.
-For all nodes $N^{\prime}$ below $N$ we know that $N^{\prime}_{C}\supset N_{C}$ since constraints are only added during a branching, never removed.
+Property (3) is automatically satisfied by CSIPP: since plans in $N_{\Pi}$ are computed using CSIPP given constraints $N_{C}$, every plan $\pi_{i}\in N_{\Pi}$ is the shortest-duration valid plan over all valid plans satisfying $N_{C}$.
+For all nodes $N^{\prime}$ below $N$ we know that $N_{C}\subset N^{\prime}_{C}$ since constraints are only added during a branching, never removed.
 Adding constraints can never result in shorter-duration plans.
@@
 Since $\sigma(\Pi)$ is strictly monotonically increasing with the maximum plan duration over all plans $\pi\in\Pi$, it follows that $\sigma(N_{\pi})\leq\sigma(N_{\pi}^{\prime})$, thereby satisfying property (3).
-In this article, we focus our attention on the theoretically described branching rule in [20] (the theoretical branching rule) and the publicly available implementation of CCBS [^2] https://github.com/PathPlanning/Continuous-CBS.
-(the implemented branching rule).
-Note that details regarding how to implement the theoretical branching rule are provided in [20], however, these details do not match the implemented branching rule and are therefore not considered here.
-II-C1 The theoretical branching rule
-Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ with $\pi_{i},\pi_{j}\in N_{\Pi}$, the earliest time $t^{i}_{u}>t^{i}$ where agent $i$ executing $\langle a_{i},t^{i}_{u}\rangle$ does not conflict with $j$ executing $\langle a^{j},t^{j}\rangle$ is found.
-This forms the unsafe interval $t^{i},t^{i}_{u})$.
-From this, a motion constraint $c_{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$ is created, specifying that $i$ is forbidden from executing $\langle a^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$.
-Similarly for $j$, the unsafe interval $t^{j},t^{j}_{u})$ is found, forming the motion constraint $c_{j}=\langle j,a_{j},[t^{j},t^{j}_{u})\rangle$.
-Therefore the theoretical branching rule uses the constraint pair $\langle\langle i,a_{i},[t^{i},t^{i}_{u})\rangle,\langle j,a_{j},[t^{j},t^{j}_{u})\rangle\rangle$ to branch on $N$.
-The CT under the theoretical branching rule is shown in [20] to satisfy properties (2) and (3), by which they then conclude that CCBS is sound and solution complete.
-Let us now revisit the example in Figure 3, where the conflict $\langle\langle m_{2}^{i},t_{2}^{i}\rangle,\langle m_{3}^{j},t_{3}^{j}\rangle\rangle$ was detected.
+In this article, we focus our attention on two branching rules: the theoretical branching rule (TBR) which is used in the soundness and solution completeness proofs in [21], and the implemented branching rule (IBR) found in the publicly available reference CCBS implementation [^2] https://github.com/PathPlanning/Continuous-CBS..
+2.3.1 The Theoretical Branching Rule
+The TBR uses the abstract definition of actions, encompassing move and wait actions alike, such that move-move and move-wait conflicts are addressed using the same procedure.
+Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$, the TBR finds the earliest time for each agent to execute its action without colliding with the other agent.
+That is, the earliest time $t^{i}_{u}>t^{i}$ for agent $i$ to execute $\langle a_{i},t^{i}_{u}\rangle$ without colliding with $j$ executing $\langle a^{j},t^{j}\rangle$ is found.
+The interval $t^{i},t^{i}_{u})$ is denoted the unsafe interval since $\langle a_{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$ certainly conflicts with $\langle a^{j},t^{j}\rangle$.
+This unsafe interval is used to define the motion constraint $c_{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$ specifying that $i$ is forbidden from executing $\langle a^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$.
+Likewise for agent $j$, the unsafe interval $t^{j},t^{j}_{u})$ is found and the motion constraint $c_{j}=\langle j,a_{j},[t^{j},t^{j}_{u})\rangle$ is defined.
+Letting the conflict be detected at CT node $N$, such that $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle\in\pi_{i}\times\pi_{j}$ with $\pi_{i},\pi_{j}\in N_{\Pi}$, the TBR uses the constraint pair $\langle c_{i},c_{j}\rangle$ to branch on $N$.
+The CT using the TBR is shown in [21] to satisfy properties (2) and (3), by which they then conclude that CCBS is sound and solution complete.
+Let us now revisit the example in Figure 3 where the conflict $\langle\langle m_{2}^{i},t_{2}^{i}\rangle,\langle m_{3}^{j},t_{3}^{j}\rangle\rangle$ was detected.
 From this conflict, the motion constraints $c^{i}=\langle i,m_{2}^{i},[t_{2}^{i},t^{i}_{u})\rangle$ and $c^{j}=\langle j,m_{3}^{j},[t_{3}^{j},t^{j}_{u})\rangle$ are formed.
@@
 Similarly at node $N^{j}$ where $N^{j}_{C}=N_{C}\cup\{c^{j}\}$, agent $j$ is no longer permitted to execute $\langle m_{3}^{j},t\rangle$ for $t\in[t_{3}^{j},t^{j}_{u})$.
-Therefore, an alternative plan is found where $j$ waits long enough to eventually execute $m^{j}_{3}$ some time after $t_{u}^{j}$.
-II-C2 The implemented branching rule
-The implemented branching rule differs from the theoretical implementation by how timed wait actions $\langle w^{j},t^{j}\rangle$ in move-wait conflicts $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ are handled.
-For such a conflict, the collision interval $I^{c}$ is found.
-A collision interval is defined as the first contiguous time interval during which $i$ and $j$ are in collision.
-Then, a vertex constraint $c^{j}=\langle j,v(w^{j}),I^{c}\rangle$ is created, forbidding $j$ from occupying vertex $v(w^{j})$ at any time during $I^{c}$.
-Besides this, timed move actions in both move-move and move-wait conflicts are handled in the same way as in the theoretical branching rule, resulting in the motion constraint $c^{i}=\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$.
-Therefore, the implemented branching rule uses the constraint pair $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,a^{j},[t^{j},t^{j}_{u})\rangle\rangle$ for move-move conflicts and $\langle\langle i,a^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$ for move-wait conflicts.
-No proofs of soundness and solution completeness are provided for CCBS under the implemented branching rule.
-II-D Revisiting CCBS
-The work in [28] calls the algorithm properties of CCBS into question. The claims are two-fold:
+Therefore, an alternative plan is found where $j$ waits long enough to eventually execute $m^{j}_{3}$ at $t_{u}^{j}$.
+2.3.2 The Implemented Branching Rule
+The IBR differs from the TBR by how move-wait conflicts $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ are branched on.
+To avoid confusion between the IBR and the described implementation in [21], we first describe the latter and then how the IBR differs.
+With the TBR, the motion constraint $\langle j,w^{j},[t^{j},t^{j}_{u})\rangle$ is created for the move-wait conflict above.
+Such a motion constraint forbids $j$ from starting the execution of $w^{j}$ at any time in $t^{j},t^{j}_{u})$.
+That is, $\langle w^{j},t\rangle$ is forbidden for all $t\in[t^{j},t^{j}_{u})$.
+However, in [21], CSIPP is described to implement such a motion constraint as a vertex constraint $\langle j,v(w^{j}),[t^{j},t^{j}_{u})\rangle$, forbidding $j$ from occupying the vertex $v(w^{j})$ during $t^{j},t^{j}_{u})$.
+Observe that such a vertex constraint is more constraining than the motion constraint; the motion constraint forbids specifically $w^{j}$ from being executed starting at any time in $t^{j},t^{j}_{u})$, while the vertex constraint forbids every wait action $w$ at the same vertex ($v(w)=v(w^{j})$) from being executed if it overlaps with $t^{j},t^{j}_{u})$, i.e., if $t,t+w_{D}]\cap[t^{j},t^{j}_{u})\neq\varnothing$.
+Therefore, already here we see a dissonance between the TBR and the described implementation.
+Similar to the described implementation, the IBR also implements vertex constraints.
+However, the IBR uses the intersection interval $\bar{I^{c}}$ instead of the unsafe interval $t^{j},t^{j}_{u})$.
+Given the conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, $\bar{I^{c}}$ (formally defined in Definition 3.4) is the time interval during which agent $i$ while executing $\langle m^{i},t^{i}\rangle$ would collide with agent $j$ waiting indefinitely at $v(w^{j})$.
+Thus, the IBR creates a vertex constraint $\langle j,v(w^{j}),\bar{I^{c}}\rangle$.
+The intersection interval $\bar{I^{c}}$ is not necessarily the same as the unsafe interval $t^{j},t^{j}_{u})$ since it does not take into account when $j$ begins executing its wait action.
+For example, if for $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ we have $\bar{t_{1}^{c}}<t^{j}<\bar{t_{2}^{c}}$, then $j$ can only begin waiting at $v(w^{j})$ from $\bar{t_{2}^{c}}$ else a collision occurs, resulting in the unsafe interval $t^{j},\bar{t_{2}^{c}})\neq\bar{I^{c}}$.
+Besides this, timed move actions in both move-move and move-wait conflicts are handled in the same way as in the TBR, resulting in the motion constraint $\langle i,a^{i},[t^{i},t^{i}_{u})\rangle$.
+Therefore, the IBR uses the constraint pair $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,m^{j},[t^{j},t^{j}_{u})\rangle\rangle$ for move-move conflicts and $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),\bar{I^{c}}\rangle\rangle$ for move-wait conflicts.
+No proofs of soundness and solution completeness are provided for CCBS under this branching rule.
+2.4 A Critique on CCBS
+The work in [29] questions the algorithm properties of CCBS. The claims are two-fold:
   1. 1.
-    under the theoretical branching rule, CCBS does not guarantee termination;
+    Under the TBR, CCBS does not guarantee termination;
   2. 2.
-    under the implemented branching rule, CCBS violates property (2).
+    Under the IBR, CCBS violates property (2).
 At its core, the identified issue lies in how wait actions are handled when resolving conflicts.
-For Claim 1[28] highlights that under the theoretical branching rule on a wait action $\langle w^{i},t^{i}\rangle$ (where the motion constraint $\langle i,w^{i},[t^{i},t^{i}_{u})\rangle$ forbidding the execution of $w^{i}$ starting at any time in $t^{i},t^{i}_{u})$ is created), only one wait duration $w^{i}_{D}$ out of infinitely many real-valued durations is removed from the search.
+For Claim 1[29] highlights that under the TBR on a wait action $\langle w^{i},t^{i}\rangle$ (giving rise to a motion constraint $\langle i,w^{i},[t^{i},t^{i}_{u})\rangle$), only one wait duration $w^{i}_{D}$ out of infinitely many real-valued durations is removed from the search.
 E.g., a wait action $w=(w^{i}_{\varphi},w^{i}_{D}+\epsilon)$ with $\epsilon$ arbitrarily small is still permitted during $t^{i},t^{i}_{u})$.
-Therefore, despite satisfying property (2) such that an optimal solution is reachable from the CT root, there may exist infinitely many nodes between them.
-Hence, CCBS under the theoretical branching rule does not always terminate after finite iterations, despite the existence of a solution.
-Therefore, CCBS under the theoretical branching rule is not solution complete.
-In other words, [28] identifies that properties (2) and (3) are not sufficient to guarantee solution completeness.
-For Claim 2, [28] provides two counter-examples showing how the implemented branching rule on a move-wait conflict removes valid solutions from the search (violating property (2)), and consequently returning sub-optimal solutions.
-Thus, CCBS under the implemented branching rule does not guarantee to only return optimal solutions and is therefore not sound.
+Therefore, despite satisfying property (2) ensuring that any optimal solution is reachable from the CT root, there may exist infinitely many nodes between such a solution and the root.
+Hence, CCBS under the TBR does not always terminate after finite iterations, despite the existence of a solution.
+Therefore, CCBS under the TBR is not solution complete.
+Essentially, [29] identifies that properties (2) and (3) are not sufficient to guarantee solution completeness.
+For Claim 2, [29] provides two counterexamples showing how the IBR on a move-wait conflict removes valid solutions from the search (violating property (2)), and consequently returning sub-optimal solutions.
+Thus, CCBS under the IBR does not guarantee to only return optimal solutions and is therefore not sound.
 Furthermore, [29] introduces shifting constraints.
-They show how a branching rule using shifting constraints, which relies on a parameter $\delta\geq 0$, does not remove collision-free solutions from the search and is therefore an improvement on the implemented branching rule.
-However, they conclude that if $\delta>0$ then there will always remain joint plans with conflicts, leading to potentially infinitely many nodes being searched.
-If instead $\delta=0$ then a similar issue as for the theoretical branching rule arises, also potentially requiring the expansion of infinitely many nodes.
-That is, CCBS using shifting constraints has the same properties as using the theoretical branching rule: sound but without termination guarantees.
-III Preliminaries
+They show how a branching rule using shifting constraints, which relies on a parameter $\delta\geq 0$, does not remove collision-free solutions from the search and is therefore an improvement on the IBR.
+However, they conclude that if $\delta>0$ then in the CT there will always remain permitted joint plans containing conflicts, leading to potentially infinitely many nodes being searched.
+If instead $\delta=0$, then a similar issue as with the TBR arises, also potentially requiring the expansion of infinitely many nodes.
+That is, CCBS using shifting constraints has the same properties as using the TBR: sound but without termination guarantees.
+3 Preliminaries
 This section establishes the theoretical foundations for analyzing branching rules in CCBS.
-We first introduce a decomposed form of CCBS constraints and extend the notion of sound constraint pairs to constraint sets.
+We first introduce a decomposed form of CCBS constraints and extend to this the notion of sound constraint pairs.
 We then define a sound branching rule as one that preserves all solutions during node expansion and prove that a branching rule is sound if and only if its underlying constraint pair is sound.
 Using these definitions, we show that a sound branching rule is sufficient to ensure that CCBS is sound, and is also solution complete if CCBS is guaranteed to terminate.
-Finally, we introduce the intersection interval which captures the complete temporal overlap between a moving agent and a stationary one — useful for designing sound branching rules that handle move-wait conflicts.
-III-A Sound Branching
+Finally, we formally introduce the collision interval and intersection interval.
+3.1 Sound Branching
 A definition for a sound pair of constraints $\langle c_{1},c_{2}\rangle$ is provided in [21, 32].
-Let the set $\mathbf{}{S}$ contain all solutions to a given MAPFR problem, and let $c(\Pi)$ be true if a joint plan $\Pi$ satisfies a constraint $c$, false otherwise.
-Definition III.1 (Sound pair of constraints [20, 31]).
-For a given MAPFR problem with solutions $\mathbf{}{S}$, a pair of constraints $\langle c_{1},c_{2}\rangle$ is sound if
-$c_{1}(\Pi)\vee c_{2}(\Pi),\quad\forall\Pi\in\mathbf{}{S}.$ (4)
+Let the set $\mathcal{S}$ contain all solutions to a given MAPFR problem, and let $c(\Pi)$ be true if a joint plan $\Pi$ satisfies a constraint $c$, false otherwise.
+Definition 3.1 (Sound pair of constraints [21, 32]).
+For a given MAPFR problem with solutions $\mathcal{S}$, a pair of constraints $\langle c_{1},c_{2}\rangle$ is sound if
+$c_{1}(\Pi)\vee c_{2}(\Pi),\quad\forall\Pi\in\mathcal{S}.$ (4)
 Throughout this article, we shall consider a decomposed form of the motion and vertex constraints by describing them as constraint sets comprised of forbidding constraints.
-These various constraint forms are equivalent, however, the decomposed form allows us to analyze constraint pairs used in branching rules more explicitly than when using motion and vertex constraints.
+These constraint forms are equivalent, however, the decomposed form allows us to analyze constraint pairs used in branching rules more explicitly than when using motion and vertex constraints.
 A forbidding constraint $\langle i,a,t\rangle$ forbids $i$ from executing the specific timed action $\langle a,t\rangle$.
@@
 We denote this equivalence by $c\Leftrightarrow C$.
-Similarly, a vertex constraint $c=\langle i,v,I^{c}\rangle$ which forbids $i$ from occupying $v$ at any time $t\in I^{c}$ is equivalent to $C=\left\{\langle i,w,t\rangle\mid v(w)=v,[t,t+w_{D}]\cap I^{c}\neq\varnothing\right\}$.
-That is, $c\Leftrightarrow C$.
+Similarly, a vertex constraint $c=\langle i,v,\bar{I^{c}}\rangle$ forbidding $i$ from occupying $v$ at any time $t\in\bar{I^{c}}$ is equivalent to $C=\left\{\langle i,w,t\rangle\mid v(w)=v,[t,t+w_{D}]\cap\bar{I^{c}}\neq\varnothing\right\}$.
 We now provide a corresponding soundness definition for a pair of constraint sets, and then define soundness of a branching rule.
-Then, in Lemma III.1, we show that a branching rule using a pair of constraint sets is sound if and only if the constraint set pair is sound.
-Definition III.2 (Sound pair of constraint sets).
+Then, in Lemma 3.1, we show that a branching rule using a pair of constraint sets is sound if and only if the constraint set pair is sound.
+Definition 3.2 (Sound pair of constraint sets).
 For a given MAPFR problem, a pair of constraint sets $\langle C_{1},C_{2}\rangle$ is sound if $\forall\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$, $\langle c_{1},c_{2}\rangle$ is sound.
-This definition follows quite naturally from extrapolating Definition III.1 to sets of constraints: given a pair of constraints $\langle c_{1},c_{2}\rangle$ and their equivalent constraint sets $C_{1}\Leftrightarrow c_{1}$ and $C_{2}\Leftrightarrow c_{2}$, it holds for all solutions $\Pi\in\mathbf{}{S}$ that
-$\quad c_{1}(\Pi)\vee c_{2}(\Pi)$
-$\Leftrightarrow$ $\quad C_{1}(\Pi)\vee C_{2}(\Pi)$
-$\Leftrightarrow$ $\quad\left(\bigwedge_{c_{1}^{\prime}\in C_{1}}c_{1}^{\prime}(\Pi)\right)\vee\left(\bigwedge_{c_{2}^{\prime}\in C_{2}}c_{2}^{\prime}(\Pi)\right)$
-$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:c_{1}^{\prime}(\Pi)\vee c_{2}^{\prime}(\Pi)$
-$\Leftrightarrow$ $\quad\forall\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\in C_{1}\times C_{2}:\langle c_{1}^{\prime},c_{2}^{\prime}\rangle\text{ is sound}.$
-Definition III.3 (Sound branching rule).
+Definition 3.3 (Sound branching rule).
 A branching rule applied to a CT node $N$ to spawn children $N^{1}$ and $N^{2}$ is sound if property (2), $S(N)=S(N^{1})\cup S(N^{2})$, always holds.
-Lemma III.1 (Equivalence of sound branching and sound constraint set pair).
+Lemma 3.1 (Equivalence of sound branching and sound constraint set pair).
 Let a branching rule applied to a CT node $N$ use a constraint set pair $\langle C_{1},C_{2}\rangle$ to spawn two children $N^{1}$ and $N^{2}$, such that $N^{1}_{C}=N_{C}\cup\left\{C_{1}\right\}$ and $N^{2}_{C}=N_{C}\cup\left\{C_{2}\right\}$.
@@
 For each child $N^{i}$, $i=1,2$, the solutions permitted under the constraints $N^{i}_{C}$ are
-$\mathcal{S}(N^{i})$ $=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N^{i}_{C}\right\}$
-$=\left\{\Pi\in\mathbf{}{S}\mid c(\Pi),c\in N_{C}\cup\left\{C_{i}\right\}\right\}$
+$\mathcal{S}(N^{i})$ $=\left\{\Pi\in\mathcal{S}\mid c(\Pi),c\in N^{i}_{C}\right\}$
+$=\left\{\Pi\in\mathcal{S}\mid c(\Pi),c\in N_{C}\cup\left\{C_{i}\right\}\right\}$
 $=\left\{\Pi\in\mathcal{S}(N)\mid C_{i}(\Pi)\right\}.$
-Therefore, for the branching rule to be sound it must hold for every solution $\Pi\in\mathcal{S}(N)$ that $C_{1}(\Pi)\vee C_{2}(\Pi)$. For every $\Pi\in\mathcal{S}(N)$, consider two cases:
+Therefore, for the branching rule to be sound it must hold for every solution $\Pi\in\mathcal{S}(N)$ that $C_{1}(\Pi)\vee C_{2}(\Pi)$ since then $\mathcal{S}(N)=\mathcal{S}(N^{1})\cup\mathcal{S}(N^{2})$.
+For every $\Pi\in\mathcal{S}(N)$, consider two cases:
   1. 1.
@@
 For the latter case, we start by assuming that $\langle C_{1},C_{2}\rangle$ is sound.
-By Definition III.2, every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$ is sound, meaning $c_{1}(\Pi)\vee c_{2}(\Pi)$ holds.
+By Definition 3.2, every pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$ is sound, meaning $c_{1}(\Pi)\vee c_{2}(\Pi)$ holds.
 Since we assumed $\neg c_{1}(\Pi)$, it follows that $c_{2}(\Pi)$ must be true.
@@
 Thus, if $\langle C_{1},C_{2}\rangle$ is sound then the branching rule is sound.
-Assume instead that $\langle C_{1},C_{2}\rangle$ is unsound: then by Definition III.2 there exists an unsound pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$.
+Assume instead that $\langle C_{1},C_{2}\rangle$ is unsound, then by Definition 3.2 there exists an unsound pair $\langle c_{1},c_{2}\rangle\in C_{1}\times C_{2}$.
 Since $\langle c_{1},c_{2}\rangle$ is unsound, there exists a valid solution $\Pi$ such that $\neg(c_{1}(\Pi)\vee c_{2}(\Pi))$, implying $\neg C_{1}(\Pi)$ and $\neg C_{2}(\Pi)$.
@@
 Thus, if $\langle C_{1},C_{2}\rangle$ is unsound then the branching rule is unsound.
-Finally, we put branching rules back into the context of the CCBS algorithm by showing in Theorem III.2 that CCBS is sound when using a sound branching rule.
-Additionally, we show in Theorem III.3 that CCBS is solution complete if it is both sound and guaranteed to terminate.
-Theorem III.2 (Soundness).
+Finally, we put branching rules back into the context of the CCBS algorithm by showing in Theorem 3.2 that CCBS is sound when using a sound branching rule.
+Additionally, we show in Theorem 3.3 that CCBS is solution complete if it is both sound and guaranteed to terminate.
+Theorem 3.2 (Soundness).
 CCBS is sound when using a sound branching rule.
@@
 Therefore, if the branching rule is sound then CCBS is sound.
-Theorem III.3 (Solution completeness).
+Theorem 3.3 (Solution completeness).
 CCBS is solution complete if it is sound and guaranteed to terminate.
@@
 CCBS’s soundness ensures that all solutions are reachable from the CT root.
-CCBS has exactly one termination condition: when it selects a CT node $N$ such that $N_{\Pi}$ is a solution, $N_{\Pi}$ is returned.
+CCBS has exactly one termination condition: when it selects a CT node $N$ such that $N_{\Pi}$ is a solution, then $N_{\Pi}$ is returned.
 CCBS continues expanding nodes until this condition is met.
@@
 ∎
-Note that Theorem III.2 shows that CCBS is sound if the branching rule is sound.
+Note that Theorem 3.2 shows that CCBS is sound if the branching rule is sound.
 However, it does not show that CCBS is unsound if the branching rule is unsound.
 Thus, the branching rule being sound is a sufficient but not necessary condition for CCBS to be sound.
-III-B Intersection Interval
-We now define the intersection interval which will be useful in Section V when introducing a our sound branching rule.
-Informally, an intersection interval is the time interval when an agent executing a timed move action is in collision with a stationary agent waiting indefinitely at a vertex.
-Let a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ be given, illustrated in Figure 4.
+3.2 Collision Interval and Intersection Interval
+We now formally define the collision interval and intersection interval, which will be useful in Section 5 when introducing $\delta$-BR.
+Given a conflict $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$, the collision interval $I^{c}$ is the first contiguous time interval during which agents $i$ and $j$ are in collision with each other while executing their respective timed actions.
+Such a collision interval must exist for any move-move or move-wait conflict, otherwise it is not a conflict.
+On the other hand, an intersection interval $\bar{I^{c}}$ exists only for a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$.
+Informally, for such a move-wait conflict, $\bar{I^{c}}$ is the time interval during which $i$ and $j$ are in collision if $j$ remains stationary at vertex $v(w^{j})$ for all time instead of only during the time when executing $\langle w^{j},t^{j}\rangle$.
+If multiple such intervals exist, then $\bar{I^{c}}$ is the interval containing $I^{c}$.
+More concretely, let a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ be given (illustrated in Figure 4).
 Suppose now that agent $j$ waits at $v(w^{j})$ for all time instead of only during $t^{j},t^{j}+w_{D}]$.
@@
 Since we know that the agents collide, during $^{c}, 𝕀¯\bar{\mathbb{I}}$ contains at least one interval super set of $I^{c}$.
-Definition III.4 (Intersection interval).
+Definition 3.4 (Intersection interval).
 Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$, let the intersection interval be the interval $\bar{I^{c}}\in\bar{\mathbb{I}}$ such that $I^{c}\subseteq\bar{I^{c}}$.
 Figure 4 illustrates the relation between $I^{c}$, $\bar{I^{c}}$ and $\bar{\mathbb{I}}$.
-It is useful to note that $I^{c}$ is the intersection between $j$’s occupancy of vertex $v(w^{j})$ and the intersection interval $\bar{I^{c}}$,
+It is useful to note that the collision interval $I^{c}$ is the intersection between $j$’s occupancy of vertex $v(w^{j})$ and the intersection interval $\bar{I^{c}}$,
 $I^{c}=[t^{j},t^{j}+w^{j}_{D}]\cap\bar{I^{c}}.$ (5)
@@
 Figure 4: A timeline of conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with conflict interval $I^{c}$ (red), the set of intersection intervals $\bar{\mathbb{I}}$ (black bars along the move action’s timeline), and the intersection interval $\bar{I^{c}}$.
-IV Analyzing CCBS’s implemented branching rule
-We use the framework from Section III to show that the implemented branching rule is not sound.
-Therefore, CCBS under the implemented branching rule does not satisfy the sufficient but not necessary condition for soundness from Theorem IV.2: a sound branching rule.
-Additionally, since the proofs in [20] for CCBS soundness rely on property (2) being satisfied, which is not fulfilled under an unsound branching rule, those proofs are not applicable either.
-Thus, this section provides complementary evidence to the counter-examples in [28] and Section VI showing that CCBS using the implemented branching rule is indeed not sound.
-Lemma IV.1 provides intermediate results that are used in Theorem IV.2 to show that the implemented branching rule is not sound.
-Lemma IV.1 (Collision timing for time-shifted move actions).
-Let no collision occur when agent $i$ executes $\langle m^{i},t^{i}\rangle$ and $j$ executes $\langle w^{j},t^{j}\rangle$.
-If $i$ instead executes $\langle m^{i},t^{i}-\epsilon\rangle$ for some $\epsilon>0$, then a collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur within the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
+4 Analyzing CCBS’s Implemented Branching Rule
+We use the framework from Section 3 to show that the IBR is not sound.
+Therefore, CCBS using the IBR does not satisfy the sufficient but not necessary condition for soundness from Theorem 4.2: a sound branching rule.
+Additionally, since the proofs in [21] for CCBS’s soundness rely on property (2) being satisfied, which is not fulfilled under the IBR, those proofs do not apply.
+Thus, this section provides complementary evidence to the counterexamples in [29] and Section 6 showing that CCBS using the IBR is indeed not sound.
+Lemma 4.1 provides intermediate results that are used in Theorem 4.2 to show that the IBR is not sound.
+Lemma 4.1 (Collision timing for time-shifted move actions).
+Let $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ not be a conflict.
+If $i$ instead executes $\langle m^{i},t^{i}-\epsilon\rangle$ for some $\epsilon>0$, then a collision with $j$ executing $\langle w^{j},t^{j}\rangle$ can only occur within the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
 Proof.
@@
 When agent $i$ executes $\langle m^{i},t^{i}-\epsilon\rangle$, its motion during the interval $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ corresponds exactly to its motion during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$.
-Since no collision occurred during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$, and agent $j$ remains stationary at the same location $v(w^{j})$, no collision can occur during $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ in the shifted case.
+Since no collision occurred during $t^{j}+\epsilon,t^{j}+w^{j}_{D}]$ when executing $\langle m^{i},t^{i}\rangle$, and agent $j$ remains stationary at the same location $v(w^{j})$, no collision occurs during $t^{j},t^{j}+w^{j}_{D}-\epsilon]$ in the shifted case.
 Therefore, any collision between $i$ and $j$ executing $\langle m^{i},t^{i}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ must occur during the remaining time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
 ∎
-Figure 5: Timeline of a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ (red).
-The resulting constraints forbid $i$ from executing $m^{i}$ starting at $t^{i},t^{i}_{u})$ (orange) and $j$ from occupying $v(w^{j})$ during $I^{c}$ (purple).
-The timed action $\langle m^{i},t^{i}_{u}\rangle$ and two forbidden timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ are shown.
-Theorem IV.2 (Unsoundness of the implemented branching rule).
-The implemented branching rule is not sound.
+Figure 5: Timeline of a conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval, $\bar{I^{c}}$ and the collision interval in red.
+The resulting constraints forbid $i$ from executing $m^{i}$ starting at $t^{i},t^{i}_{u})$ (orange) and $j$ from occupying $v(w^{j})$ during $\bar{I^{c}}$ (purple).
+The timed action $\langle m^{i},t^{i}_{u}\rangle$ and two non-conflicting forbidden timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ are shown.
+Theorem 4.2 (Unsoundness of the IBR).
+CCBS’s IBR is not sound.
 Proof.
-Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$.
-From this, the implemented branching rule uses the constraint pair $\langle\langle i,m^{i},[t^{i},t^{i}_{u})\rangle,\langle j,v(w^{j}),I^{c}\rangle\rangle$.
-Letting $C_{i}\Leftrightarrow\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $C_{j}\Leftrightarrow\langle j,v(w^{j}),I^{c}\rangle$, the branching rule equivalently uses the constraint set pair $\langle C_{i},C_{j}\rangle$.
-Figure 5 illustrates a timeline of this conflict and the constrained intervals $t^{i},t^{i}_{u})$ and $I^{c}$.
-By construction of $C_{i}$, we know that $c_{i}=\langle i,m^{i},t^{i}_{u}-\epsilon\rangle\in C_{i}$ for an arbitrarily small $\epsilon>0$.
+Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$.
+From this, the IBR creates the constraints $\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $\langle j,v(w^{j}),\bar{I^{c}}\rangle$.
+Letting $C_{i}\Leftrightarrow\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $C_{j}\Leftrightarrow\langle j,v(w^{j}),\bar{I^{c}}\rangle$, the IBR branches using the constraint set pair $\langle C_{i},C_{j}\rangle$.
+The move-wait conflict and the constraint intervals $t^{i},t^{i}_{u})$ and $\bar{I^{c}}$ are illustrated in the upper part of Figure 5.
+By construction of $C_{i}$, for $0<\epsilon<t^{i}_{u}-t^{i}$ we know that $\langle i,m^{i},t^{i}_{u}-\epsilon\rangle\in C_{i}$.
 By definition of $t^{i}_{u}$, $\langle m^{i},t^{i}_{u}\rangle$ and $\langle w^{j},t^{j}\rangle$ do not conflict.
-It then follows by Lemma IV.1 that any collision between agent $i$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $j$ executing $\langle w^{j},t^{j}\rangle$ must occur during the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
-See Figure 5 for the timelines of the timed move actions $\langle m^{i},t^{i}_{u}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$.
-The constraint set $C_{j}$ contains one forbidding constraint for each timed action that requires occupying $v(w^{j})$ during $I^{c}$.
-We now aim to find a forbidden timed wait action $\langle w^{\prime},t^{\prime}\rangle$ ($\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$) that does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$.
+It then follows from Lemma 4.1 that any collision between agent $i$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $j$ executing $\langle w^{j},t^{j}\rangle$ must occur during the time interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
+The timelines of $\langle m^{i},t^{i}_{u}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ are illustrated in the lower part of Figure 5.
+The constraint set $C_{j}$ contains one forbidding constraint for each timed action that requires occupying $v(w^{j})$ during $\bar{I^{c}}$.
+We now aim to find a forbidden timed wait action $\langle w^{\prime},t^{\prime}\rangle$ (i.e., $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$) that does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$.
 Let
@@
 Recall that a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{j},t^{j}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
-Therefore, the same is true when executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$, a collision can only occur during the same interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ (assuming $\epsilon<w^{\prime}_{D}$).
-From this, we now constrain $t^{\prime}+w^{\prime}_{D}<t^{j}+w^{j}_{D}-\epsilon$ such that $\langle w^{\prime},t^{\prime}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ do not collide:
+The same also holds for $\langle w^{\prime},t^{\prime}\rangle$ since it entails $j$ waiting stationary at the same vertex; a collision between $i$ and $j$ executing $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ can only occur during $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
+We can use this by constraining $\langle w^{\prime},t^{\prime}\rangle$ to end before the start of this interval, that is $t^{\prime}+w^{\prime}_{D}<t^{j}+w^{j}_{D}-\epsilon$, and by doing so ensuring that $\langle w^{\prime},t^{\prime}\rangle$ and $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ do not collide:
 $t^{\prime},t^{\prime}+w^{\prime}_{D}]\subseteq[t^{j},t^{j}+w^{j}_{D}-\epsilon).$ (6)
-Finally, for $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$ we require $t^{\prime},t^{\prime}+w^{\prime}_{D}]$ to intersect with $I^{c}$.
-Since $\epsilon$ is arbitrarily small, let $\epsilon<|I^{c}|$.
-It follows that the interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ of size $\epsilon$ can only span at most a subset of $I^{c}$.
-Additionally, by (5), $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ such that $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ must be a suffix of $I^{c}$.
-The interval $t_{1}^{c},t^{j}+w^{j}_{D}-\epsilon)$ is therefore non-degenerate.
-Thus, we require (6) and $t_{1}^{c}\leq t^{\prime}+w^{\prime}_{D}$ to hold.
-Since any $t^{\prime}\in\Re$ and $w^{j}_{D}\in\Re$ can be selected, we know that there exists a forbidding constraint $c_{j}=\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$.
-See Figure 5 for an example of such a timed wait action $\langle w^{\prime},t^{\prime}\rangle$.
-In conclusion, the timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ do not conflict; $i$ collides with $j$ stationary at $v(w^{j})$ only at some time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$, while $j$ executing $\langle w^{\prime},t^{\prime}\rangle$ does not occupy $v(w^{j})$ at any time in $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$.
+Finally, to ensure that $\langle j,w^{\prime},t^{\prime}\rangle$ is indeed a member of $C_{j}$, we require $t^{\prime},t^{\prime}+w^{\prime}_{D}]$ to intersect with $\bar{I^{c}}$.
+Since $\epsilon$ can be arbitrarily small, let $\epsilon<|\bar{I^{c}}|$.
+It follows that the interval $t^{j}+w^{j}_{D}-\epsilon,t^{j}+w^{j}_{D}]$ of size $\epsilon$ can only span at most a subset of $\bar{I^{c}}$.
+Thus, a non-degenerate subset of $t^{j},t^{j}+w^{j}_{D}-\epsilon)$ is also included in $\bar{I^{c}}$.
+Therefore, there exists a $\langle w^{\prime},t^{\prime}\rangle$ such that (6) is satisfied and $t^{\prime},t^{\prime}+w_{D}^{\prime})\cap\bar{I^{c}}\neq\varnothing$.
+These properties ensure that $\langle w^{\prime},t^{\prime}\rangle$ does not conflict with $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and that $\langle j,w^{\prime},t^{\prime}\rangle\in C_{j}$.
+An example of such a timed wait action is illustrated at the bottom of Figure 5.
+In conclusion, the timed actions $\langle m^{i},t^{i}_{u}-\epsilon\rangle$ and $\langle w^{\prime},t^{\prime}\rangle$ do not conflict.
 Therefore, a solution $\Pi$ can contain both of these timed actions.
-However, such a solution does not satisfy $c_{i}(\Pi)\vee c_{j}(\Pi)$.
-Thus, $\langle c_{i},c_{j}\rangle$ is not sound, hence neither is $\langle C_{i},C_{j}\rangle$.
-By Lemma III.1, the implemented branching rule is not sound.
+However, letting $c_{i}=\langle i,m^{i},t^{i}_{u}-\epsilon\rangle$ and $c_{j}=\langle j,w^{\prime},t^{\prime}\rangle$, $c_{i}(\Pi)\vee c_{j}(\Pi)$ is not satisfied.
+Thus, $\langle c_{i},c_{j}\rangle$ is not sound, and since $c_{i}\in C_{i}$ and $c_{j}\in C_{j}$, $\langle C_{i},C_{j}\rangle$ is not sound either.
+By Lemma 3.1, the IBR is therefore not sound.
 ∎
-V A Branching Rule for Sound and Solution-Complete CCBS
-In this section, we present our novel branching rule and prove that CCBS using this branching rule is sound and solution complete.
-To do so, we first show in Section V-A that our branching rule is sound.
-To provide termination guarantees, we formulate a non-termination condition in Section V-B and then show in Section V-C that CCBS under our branching rule does satisfy this condition.
-Finally, we use these results to conclude in Section V-D that CCBS under our branching rule is sound and solution complete.
-Our branching rule is based on shifting constraints [28] can be understood by considering a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$.
-By construction of $\bar{I^{c}}$, when $i$ executes $\langle m^{i},t^{i}\rangle$ then a collision occurs if $j$ occupies $v(w^{j})$ at any time in $\bar{I^{c}}$.
+5 A Branching Rule for Sound and Solution-Complete CCBS
+In this section, we present $\delta$-BR and prove that CCBS using $\delta$-BR is sound and solution complete.
+To do so, we first show in Section 5.1 that $\delta$-BR is sound.
+Termination guarantees are provided by formulating a non-termination condition in Section 5.2 and then showing in Section 5.3 that CCBS under $\delta$-BR does not satisfy this condition.
+Finally, these results are used in Section 5.4 to conclude that CCBS under $\delta$-BR is sound and solution complete.
+$\delta$-BR is based on shifting constraints [29] which can be understood by considering a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$.
+By definition of $\bar{I^{c}}$, when agent $i$ executes $\langle m^{i},t^{i}\rangle$ then a collision occurs if agent $j$ occupies $v(w^{j})$ at any time in $\bar{I^{c}}$.
 Suppose now that $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ with $\Delta t\in\Re$ instead.
-Then, the intersection interval will also shift by $\Delta t$, such that a collision occurs if $j$ occupies $v(w^{j})$ at anytime in $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$.
-Fig.
-6 illustrates the timeline of $m^{i}$ starting at $t^{i}$ and incrementally later times.
-Now, let some $0\leq\delta<|\bar{I^{c}}|$ be given.
-If $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, with shifted intersection interval $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$, then we know for any value of $\Delta t$ that the interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ will be included in the resulting intersection interval.
-In other words, if $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, then we know that a collision will arise if agent $j$ occupies $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+Then, the intersection interval will also shift by $\Delta t$ such that a collision occurs if $j$ occupies $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$.
+Figure 6 illustrates the timeline of $m^{i}$ starting at $t^{i}$ and incrementally later times.
+Now, let some $\delta$ be given such that $0\leq\delta<|\bar{I^{c}}|$.
+Then, if agent $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, we know that the shifted intersection interval $\bar{t_{1}^{c}}+\Delta t,\bar{t_{2}^{c}}+\Delta t)$ will at least contain the interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (shown in Figure 6).
+Therefore, we know that if agent $i$ executes $\langle m^{i},t^{i}+\Delta t\rangle$ for any $0\leq\Delta t\leq\delta$, then agent $j$ cannot occupy vertex $v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ without a collision occurring.
 This result is useful as it allows us to form a pair of sound constraints from a move-wait conflict.
-Figure 6: If move action $m^{i}$ is executed at any time in the (orange) interval $t^{i},t^{i}+\delta)$, then $i$ will necessarily intersect with $j$ waiting at $v(w)$ during the (blue) interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
-We now define our branching rule.
-Definition V.1 (Our novel branching rule).
-Move-move and move-wait conflicts are handled differently:
-  • Given a move-move conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$, the theoretical branching rule is applied.
-That is, the unsafe intervals $t^{i},t^{i}_{u})$ and $t^{j},t^{j}_{u})$ are found.
-Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$, and $j$ is forbidden from executing $\langle m^{j},t\rangle$ for any $t\in[t^{j},t^{j}_{u})$.
-  • Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$, let
+Figure 6: Shifting the start of the timed move action $\langle m^{i},t^{i}\rangle$ also shifts the intersection interval $\bar{I^{c}}$.
+Executing $m^{i}$ at any time in the (orange) interval $t^{i},t^{i}+\delta)$ will result in a shifted intersection interval which necessarily includes the (purple) interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+Definition 5.1 ($\delta$-BR).
+$\delta$-BR handles move-move and move-wait conflicts differently:
+  • Given a move-move conflict $\langle\langle m^{i},t^{i}\rangle,\langle m^{j},t^{j}\rangle\rangle$, the TBR/IBR is applied: the unsafe intervals $t^{i},t^{i}_{u})$ and $t^{j},t^{j}_{u})$ are found and the pair of motion constraints $\langle i,m^{i},[t^{i},t^{i}_{u})\rangle$ and $\langle j,m^{j},[t^{j},t^{j}_{u})\rangle$ are used to branch on.
+That is, agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}_{u})$ and agent $j$ is forbidden from executing $\langle m^{j},t\rangle$ for any $t\in[t^{j},t^{j}_{u})$.
+  • Given a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$, let
     $\delta=\min\left(\gamma|\bar{I^{c}}|,\;t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$ (7)
     where $0<\gamma<1$ is a fixed constant.
-Agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for any $t\in[t^{i},t^{i}+\delta)$.
-Agent $j$ is forbidden from occupying the vertex $v(w_{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (vertex constraint) and executing any timed move action $\langle m,t\rangle$ where $\mathit{from}(m)=v(w^{j})$ and $t\in[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
-The fixed constant $\gamma\in(0,1)$ dictates which agent in a move-wait conflict is more heavily constrained: setting $\gamma$ near $1$ increases the constrained interval on agent $i$ while setting it closer to 0 increases the constrained interval on $j$.
-Figure 7 illustrates three different cases where our branching rule is applied to a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$.
-In the first case $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$ while in the second and third case $\delta=\gamma|\bar{I^{c}}|$.
-Figure 7: $\delta$ for three different move-wait conflicts, with $\gamma=0.6$.
-In the top cases, $\delta=t_{2}^{c}-\bar{t_{1}^{c}}$, in the bottom two cases $\delta=\gamma(\bar{t_{2}^{c}}-\bar{t_{1}^{c}})$.
-Orange interval is the restriction on the move action, purple is the restriction on the wait action.
-V-A Soundness
-We now show that our branching rule is sound.
-The branching rule applies the same rule as in CCBS on move-move conflicts, which is shown in Section 4.2 of [20] to satisfy property (2) and is therefore sound.
+For agent $i$, the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$ is created.
+For agent $j$, the vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for every $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$ is created.
+That is, agent $i$ is forbidden from executing $m^{i}$ at any time in $t^{i},t^{i}+\delta)$, and agent $j$ is forbidden from occupying $v(w_{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ or executing any move action from $v(w_{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+The fixed constant $\gamma\in(0,1)$ controls which agent in a move-wait conflict is more constrained: setting $\gamma$ near $1$ increases the constrained interval on the moving agent while setting $\gamma$ closer to 0 increases the constrained interval on waiting agent.
+Figure 7 illustrates four different cases where $\delta$-BR (with $\gamma=0.7$) is applied to a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$.
+In the first two cases (Figures 7(a) and 7(b)) we have $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ since $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}<\gamma|\bar{I^{c}}|$.
+Thus, agent $j$ is forbidden from occupying $v(w^{j})$ from $t^{j}+w^{j}_{D}$ until the end of $\bar{I^{c}}$.
+On the other hand, in the second two cases (Figures 7(c) and 7(d)) we have $\delta=\gamma|\bar{I^{c}}|$.
+Thus, a non-singular part of agent $j$’s occupation of $v(w^{j})$ is included in the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+Furthermore, it is also visible that agent $j$’s constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is smaller than agent $i$’s constrained interval $t^{i},t^{i}+\delta)$ as a result of $\gamma>0.5$.
+(a) $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$
+(b) $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$
+(c) $\delta=\gamma|\bar{I^{c}}|$
+(d) $\delta=\gamma|\bar{I^{c}}|$
+Figure 7: An illustration of $\delta$-BR applied to four different move-wait conflicts, $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$, where the start time $t^{j}$ of wait action $w^{j}$ progressively increases for each sub-figure.
+$\delta=\min(\gamma|\bar{I^{c}}|,\;t^{j}+w^{j}_{D}-\bar{t_{1}^{c}})$ with $\gamma=0.7$ is shown; $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ in Figures 7(a) and 7(b), and $\delta=\gamma|\bar{I^{c}}|$ in Figures 7(c) and 7(d).
+Agent $i$’s constrained interval $t^{i},t^{i}+\delta)$ (orange) and agent $j$’s constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ (purple) are shown.
+5.1 Soundness
+We now show that $\delta$-BR is sound.
+$\delta$-BR applies TBR/IBR on move-move conflicts, which is shown in Section 4.2 of [21] to satisfy property (2) and is therefore sound.
 Hence, we refer the reader to there for more details.
-What remains is to prove the soundness of our branching rule on move-wait conflicts.
-Theorem V.1 (Soundness of our branching rule).
-Our branching rule is sound when applied to a move-wait conflict.
+What remains is to prove the soundness of $\delta$-BR on move-wait conflicts.
+Theorem 5.1 (Soundness of $\delta$-BR).
+$\delta$-BR is sound when applied to a move-wait conflict.
 Proof.
-Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}$ and intersection interval $\bar{I^{c}}$.
-For some $\delta<|\bar{I^{c}}|$, agent $i$ is forbidden from executing $\langle m^{i},t\rangle$ for $t\in[t^{i},t^{i}+\delta)$.
-The equivalent constraint set of forbidding constraints is $C_{i}=\left\{\langle i,m^{i},t\rangle\mid t\in[t^{i},t^{i}+\delta)\right\}$.
-As this is a shifting constraint, we know that a collision occurs with $j$ located at $v(w_{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
-On the other hand, agent $j$ is forbidden from executing any wait action that requires occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, or any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
-All of these forbidden timed actions require being located at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
-Thus, a collision will occur if $i$ and $j$ each execute any of their forbidden actions.
-Let $C_{j}$ contain one forbidding constraint for each of $j$’s forbidden actions.
-Then, it holds for any $\langle c_{i},c_{j}\rangle\in\langle C_{i},C_{j}\rangle$ that a valid solution must satisfy at least $c_{i}$ or $c_{j}$.
-Therefore, $\langle C_{i},C_{j}\rangle$ is sound and by Lemma III.1 so is our branching rule.
+Consider a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}$, on which we apply $\delta$-BR.
+For agent $i$, $\delta$-BR creates the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$.
+Since the motion constraint is in the form of a shifting constraint, and by (7) we know that $\delta<|\bar{I^{c}}|$, we know that a collision occurs if agent $j$ occupies vertex $v(w^{j})$ at any time during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+The motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$ can equivalently be expressed in its decomposed form — a set of forbidding constraints — $C_{i}=\left\{\langle i,m^{i},t\rangle\mid t\in[t^{i},t^{i}+\delta)\right\}$.
+On the other hand, $\delta$-BR creates for agent $j$ a vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and the motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for every $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$.
+These can collectively be represented by the set of forbidding constraints
+$C_{j}$ $=\left\{\langle j,w,t\rangle\mid v(w)=v(w^{j}),[t,t+w_{D}]\cap[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\neq\varnothing\right\}$
+$\cup\left\{\langle j,m,t\rangle\mid\mathit{from}(m)=v(w^{j}),m\in\mathcal{A},t\in[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\right\}$
+For every constraint $\langle j,a,t\rangle\in C_{j}$, the forbidden timed action $\langle a,t\rangle$ requires $j$ being located at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$; the forbidden timed wait actions require $j$ occupying $v(w^{j})$ during some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and the forbidden timed move actions start at $v(w^{j})$ at some time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+Thus, if agent $i$ executes any of the actions forbidden by $C_{i}$ and agent $j$ executes any of the actions forbidden by $C_{j}$, then a collision will occur.
+The above implies that for every $\langle c_{i},c_{j}\rangle=\langle\langle i,a^{i},t^{i}\rangle,\langle j,a^{j},t^{j}\rangle\rangle\in C_{i}\times C_{j}$ the pair of timed actions $\langle\langle a^{i},t^{i}\rangle,\langle a^{j},t^{j}\rangle\rangle$ is a conflict.
+Thus, any solution must satisfy $c_{i}\vee c_{j}$, meaning that $\langle c_{i},c_{j}\rangle$ is sound for every $\langle c_{i},c_{j}\rangle\in C_{i}\times C_{j}$ and therefore that $\langle C_{i},C_{j}\rangle$ is sound.
+By Lemma 3.1, this means that $\delta$-BR is sound on any move-wait conflict.
 ∎
-To conclude, our branching rule is shown to be sound when applied to move-move conflicts (in [20]) and move-wait conflicts (in Theorem V.1). Therefore, the branching rule is sound.
-V-B A Requirement for Non-termination
-To prove that CCBS under our branching rule is solution complete and therefore terminates on a solvable MAPFR problem, we employ a proof by contradiction.
-We begin by assuming the opposite: that CCBS never terminates on some solvable MAPFR problem, despite the existence of a solution, and arrive at a condition which must hold if CCBS never terminates.
+To conclude, $\delta$-BR is shown to be sound when applied to move-move conflicts (in [21]) and move-wait conflicts (in Theorem 5.1).
+As these are the only two conflict types considered, $\delta$-BR is therefore sound.
+5.2 A Requirement for Non-termination
+To prove that CCBS using $\delta$-BR is solution complete and therefore terminates on a solvable MAPFR problem, we employ a proof by contradiction.
+We begin by assuming the opposite: that CCBS never terminates on a solvable MAPFR problem despite the existence of a solution.
+This will result in a condition which must hold if CCBS never terminates.
 We will establish in this section that if CCBS runs indefinitely, then the CT must contain an infinite sequence of nodes that are continuously selected and branched upon.
 By analyzing this infinite sequence, we derive a specific requirement that any non-terminating execution must satisfy.
-It will later be show in Section V-C that CCBS under our branching rule does not satisfy this requirement, thus proving termination.
+It will later be show in Section 5.3 that CCBS using $\delta$-BR does not satisfy this requirement, thus proving termination.
 Specifically, what we establish here is:
   1.
-V-B1) Infinite CT node sequence: There exists an infinite descending path in the constraint tree where each node contains conflicts and has objective value $\leq\sigma^{*}$, where $\sigma^{*}$ is the optimal.
-  2. V-B2) Concentration on $\tau^{\infty}$: Infinitely many nodes in the sequence map to the same trajectory $\tau^{\infty}$.
+5.2.1) Infinite CT node sequence: There exists an infinite descending path in the constraint tree where each node contains conflicts and has objective value $\leq\sigma^{*}$, with $\sigma^{*}$ being the optimal.
+  2.
+5.2.2) Concentration on $\tau^{\infty}$: We define a trajectory in Definition 5.2 and a mapping from a CT node to a trajectory in Definition 5.5, and find that infinitely many nodes in the sequence map to the same trajectory $\tau^{\infty}$.
   3.
-V-B3) Unbounded Constraint Accumulation: Constraints which determine how nodes map to $\tau^{\infty}$ accumulate unboundedly along the infinite node sequence, yet there must always exist another node mapping to $\tau^{\infty}$ and still maintaining an objective value lower than the optimal value.
-The final point formulates the requirement on the branching rule for non-termination.
-We begin by defining trajectories and how plans map to them.
-Definition V.2 (Trajectory).
+5.2.3) Unbounded Constraint Accumulation: Constraints which determine how nodes map to $\tau^{\infty}$ accumulate unboundedly along the infinite node sequence, yet there must always exist another node mapping to $\tau^{\infty}$ which still maintains an objective value lower than the optimal value.
+The final point formulates the requirement on the branching rule for non-termination. We begin by defining trajectories and how plans map to them.
+Definition 5.2 (Trajectory).
 A trajectory $\tau$ is a finite sequence of move actions, $\tau=\langle m_{1},m_{2}\dots,m_{n}\rangle\in\mathcal{A}^{n}$.
 Each move action $m_{i}\in\tau$ has a duration $m_{i,D}$. Thus, the duration of $\tau$ is $\tau_{D}=\sum_{k=1}^{n}m_{k,D}$.
-Definition V.3 (Joint Trajectory).
+Definition 5.3 (Joint Trajectory).
 A joint trajectory $\mathcal{T}$ is a set containing one trajectory $\tau_{i}$ for each agent $i$.
-Definition V.4 (Plan-Trajectory Mapping).
+Definition 5.4 (Plan-Trajectory Mapping).
 A plan $\pi$ maps to a trajectory $\tau$, denoted $\pi\sim\tau$, if $\tau$ contains exactly every move action in $\pi$ in the order that they appear.
-Similarly, $\Pi\sim\mathcal{T}$ if it holds for every agent $i$ that $\pi_{i}\sim\tau_{i}$ with $\pi_{i}\in\Pi,\tau_{i}\in\mathcal{T}$.
-Informally, a trajectory $\tau$ describes the sequence of moves an agent takes when following a plan $\pi\sim\tau$, but without timing details.
+Similarly, a joint plan $\Pi$ maps to a joint trajectory $\mathcal{T}$, $\Pi\sim\mathcal{T}$, if it holds for every agent $i$ that $\pi_{i}\sim\tau_{i}$ with $\pi_{i}\in\Pi,\tau_{i}\in\mathcal{T}$.
+Informally, a trajectory $\tau$ describes the sequence of move actions an agent takes when following a plan $\pi\sim\tau$, but without timing details.
 Multiple plans can therefore map to the same trajectory, as plans can contain various wait actions.
 For example, the following two plans
-  • $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ from Figure 2, and
+  • $\pi_{i}=\langle\langle m_{1}^{i},0\rangle,\langle w_{2}^{i},1.5\rangle,\langle m_{3}^{i},2\rangle,\langle m_{4}^{i},3\rangle,\langle w_{5}^{i},4.3\rangle\rangle$ introduced in Figure 2, and
   • $\pi_{i}^{\prime}=\langle\langle m_{1}^{i},0\rangle,\langle m_{3}^{i},1.5\rangle,\langle m_{4}^{i},2.5\rangle\rangle$
 both map to the same trajectory $\tau_{i}=\langle m_{1}^{i},m_{3}^{i},m_{4}^{i}\rangle$.
-Since each trajectory $\tau$ has a duration $\tau_{D}$, a joint trajectory $\mathcal{T}$ has an objective value $\sigma(\mathcal{T})$.
+Since each trajectory $\tau$ has a duration $\tau_{D}$, a joint trajectory $\mathcal{T}$ can be evaluated on the objective function, $\sigma(\mathcal{T})$.
 Observe that $\tau_{D}\leq\pi_{D}$ if $\pi\sim\tau$ since $\pi$ contains at least all the move actions in $\tau$ but possibly additional wait actions.
@@
 This follows from the trajectories $\tau$ being finite sequences from the countable set $\mathcal{A}$, meaning that the set of all $\tau$ is countable.
-Since the set of all $\tau$ is countable, and a joint trajectory $\mathcal{T}$ comprises a discrete number of trajectories (exactly one per agent), the set of all possible joint trajectories $\mathbb{T}$ is countable.
+Since the set of all $\tau$ is countable, and a joint trajectory $\mathcal{T}$ comprises a discrete number of trajectories (exactly one per agent), the set $\mathbb{T}$ is countable.
 We define the subset $\mathbb{T}^{c}\subset\mathbb{T}$ to contain all joint trajectories with an objective value at or below some fixed constant $c\in\Re$:
@@
 It holds that $\mathbb{T}^{c}$ is finite for any $c<\infty$.
-V-B1 Infinite CT node sequence
-Since our branching rule is sound (Section V-A), all solutions (including the optimal solution) is reachable from the CT root.
-Non-termination implies that at every iteration, a non-solution CT node $N$ is expanded instead of a node representing an optimal solution with objective value $\sigma^{*}$.
+5.2.1 Infinite CT node sequence
+Since $\delta$-BR is sound (Section 5.1), all solutions (including all optimal solutions) are reachable from the CT root.
+Non-termination of CCBS implies that at every iteration, a non-solution CT node $N$ is expanded instead of a node representing an optimal solution with objective value $\sigma^{*}$.
 Since the search is best-first, it must hold that $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$ otherwise it would not be selected for expansion.
@@
 For non-termination to persist indefinitely, such nodes must be available at every iteration.
-This leads to the following recursive requirement: at every iteration there must exist at least one node $N$ satisfying the above two conditions, with a child node also satisfying the above two conditions, otherwise eventually all nodes satisfying these properties will be depleted and the algorithm will terminate with an optimal solution.
-In summary, non-termination implies that at every iteration exists at least one CT node $N$ where
+This leads to the following recursive requirement: at every iteration there must exist at least one node $N$ satisfying the above two conditions, with at least one child node also satisfying the above two conditions, otherwise eventually all nodes satisfying these properties will be depleted and the algorithm will terminate with an optimal solution.
+In summary, non-termination implies that at every iteration there exists at least one CT node $N$ where
   1. (I)
@@
   2. (II)
-    $N_{\Pi}$ contains a conflict, and
+    $N_{\Pi}$ contains conflicts and is therefore not a solution, and
   3. (III)
-    a child node of NN satisfies (I) and (II).
+    at least one child node of NN satisfies (I) and (II).
 From this it follows the existence of an infinite descending path
@@
 in the CT, where for every $k=1,2,\dots$ it holds that $N^{k}$ satisfies (I)-(III) and $N^{k+1}$ is a child of $N^{k}$.
-V-B2 Concentration on $\tau^{\infty}$
+5.2.2 Concentration on $\tau^{\infty}$
 Consider a node $N^{k}\in\varsigma$.
-When $N^{k}$ is expanded, CCBS detects a conflict between two plans in $N^{k}_{\Pi}$ and applies the branching rule to create two child nodes.
+When $N^{k}$ is expanded, CCBS detects a conflict between two plans in $N^{k}_{\Pi}$ and applies a branching rule to create two child nodes.
 By definition of sequence $\varsigma$, exactly one of these children continues the sequence as $N^{k+1}$.
-For one of the plans involved in the conflict, say $\pi_{i}\in N^{k}_{\Pi}$, a constraint is created for agent $i$ and added to the constraint set of $N^{k+1}$.
+For one of the plans involved in the conflict, say $\pi_{i}\in N^{k}_{\Pi}$, a constraint is created for agent $i$ and added to the constraint set $N^{k+1}_{C}$.
 In the context of our analysis, we say that $\pi_{i}$ is “branched on” at $N^{k}$.
-Definition V.5 (Node-trajectory mapping).
+Definition 5.5 (Node-trajectory mapping).
 A node $N\in\varsigma$, where $\pi$ is branched on, maps to the trajectory $\tau$ for which $\pi\sim\tau$.
-With Definition V.5, we can map every one of the infinitely many nodes in $\varsigma$ to a specific trajectory.
+With Definition 5.5, we can map every one of the infinitely many nodes in $\varsigma$ to a specific trajectory.
 We now establish that infinitely many nodes in $\varsigma$ map to a specific trajectory, denoted $\tau^{\infty}$.
 Consider $N\in\varsigma$.
-The set of plans $N_{\Pi}$ satisfies $N_{\Pi}\sim\mathcal{T}$ for some $\mathcal{T}\in\mathbb{T}$, and $N$ maps to a trajectory $\tau\in\mathcal{T}$.
+The joint plan $N_{\Pi}$ satisfies $N_{\Pi}\sim\mathcal{T}$ for some $\mathcal{T}\in\mathbb{T}$, and $N$ maps to a trajectory $\tau\in\mathcal{T}$.
 By inequality (8), $\sigma\left(\mathcal{T}\right)\leq\sigma\left(N_{\Pi}\right)$, and by condition (I), $\sigma\left(N_{\Pi}\right)\leq\sigma^{*}$.
-Therefore, it follows that $\sigma\left(\mathcal{T}\right)\leq\sigma^{*}$, implying $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$.
+Therefore, it follows that $\sigma\left(\mathcal{T}\right)\leq\sigma^{*}$ and therefore that $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$.
 Since $\mathbb{T}^{\sigma^{*}}$ is finite and each $\mathcal{T}\in\mathbb{T}^{\sigma^{*}}$ contains a finite number of trajectories, the set of all possible trajectories that nodes in $\varsigma$ map to is finite.
 However, $\varsigma$ contains infinitely many nodes.
-By the Pigeonhole principle, infinitely many nodes in $\varsigma$ must map to at least one specific trajectory $\tau^{\infty}$.
-V-B3 Unbounded Constraint Accumulation
-For every branching at a node $N^{k}\in\varsigma$, a constraint is created and propagated to all subsequent nodes $N^{h}\in\varsigma$ for $h>k$.
-Since infinitely many nodes in $\varsigma$ map to $\tau^{\infty}$, constraints from branching on plans $\pi\sim\tau^{\infty}$ accumulate unboundedly, yet there must always exist some $\pi^{\prime}\sim\tau^{\infty}$ permitted under these constraints.
+By the Pigeonhole principle, there exists at least one trajectory $\tau^{\infty}$ which infinitely many nodes in $\varsigma$ map to.
+5.2.3 Unbounded Constraint Accumulation
+For every branching at a node $N^{k}\in\varsigma$, a constraint $c$ is created and propagated to all subsequent nodes $N^{h}\in\varsigma$ for $h>k$, such that $c\in N^{h}_{C}$.
+Since infinitely many nodes in $\varsigma$ map to $\tau^{\infty}$, constraints from branching on plans $\pi\sim\tau^{\infty}$ accumulate unboundedly along $\varsigma$, yet there must always exist some $\pi^{\prime}\sim\tau^{\infty}$ permitted under these constraints.
 On the other hand, by condition (I), every $N^{k}\in\varsigma$ must satisfy $\sigma\left(N^{k}_{\Pi}\right)\leq\sigma^{*}$.
 Since $\sigma$ is strictly monotonically increasing with the maximum plan duration, there exists a fixed constant $c<\infty$ (determined by $\sigma^{*}$) such that all plans in $N^{k}_{\Pi}$ have duration $\leq c$.
-Therefore, it must hold for all $N^{k}\in\varsigma$ that $\forall\pi\in N^{k}_{\Pi}:\pi_{D}\leq c$, otherwise $\sigma\left(N^{k}_{\Pi}\right)$ grows unboundedly to eventually violate condition (I).
-Finally, we have arrived at our requirement for non-termination: for some fixed $c<\infty$ and trajectory $\tau^{\infty}$, infinitely many CCBS branching steps on a plan $\pi\sim\tau^{\infty}$ must preserve the existence of at least one valid plan $\pi^{\prime}\sim\tau^{\infty}$ with duration $\pi_{D}\leq c$.
-V-C Solution Completeness
-In this section, we prove that CCBS under our branching rule terminates within finite iterations by showing that the non-termination requirement established in Section V-B cannot be satisfied.
-This completes our proof by contradiction.
-We prove that CCBS under our branching rule does not satisfy the non-termination requirement by showing that each branching operation reduces the set of feasible execution times for actions in $\tau^{\infty}$ by non-degenerate amounts, eventually making it impossible to construct any plan $\pi\sim\tau^{\infty}$ with a bounded duration.
+Therefore, it must hold for all $N^{k}\in\varsigma$ that $\forall\pi\in N^{k}_{\Pi}:\pi_{D}\leq c$, otherwise $\sigma\left(N^{k}_{\Pi}\right)$ necessarily grows to eventually violate condition (I).
+Finally, we have arrived at our requirement for non-termination: for some fixed $c<\infty$ and trajectory $\tau^{\infty}$, there must exists some plan $\pi\sim\tau^{\infty}$ with duration $\pi_{D}\leq c$ that is permitted under the constraints accumulated from infinitely many branchings on plans also mapping to $\tau^{\infty}$.
+5.3 Solution Completeness
+In this section, we prove that CCBS under $\delta$-BR terminates within finite iterations by showing that the non-termination requirement established in Section 5.2 cannot be satisfied, thereby completing our proof by contradiction.
+This is done by showing that each branching operation reduces the set of feasible execution times for actions in $\tau^{\infty}$ by non-degenerate amounts, eventually making it impossible to construct any plan $\pi\sim\tau^{\infty}$ with a bounded duration.
 The proof is structured in five stages:
-  1. V-C1) Permitted Execution Times: We formalize how constraints restrict move execution and vertex occupancy times, establishing the framework for analyzing constraint accumulation.
-  2. V-C2) CSIPP Planning Behavior: We explain how CCBS’s underlying path planner (CSIPP) constructs plans, focusing on how it selects execution times from feasible intervals.
-  3. V-C3) Non-degenerate Reductions For Move Actions: We show how our branching rule applied to a timed move action reduces feasible move execution times by a non-degenerate amount.
+  1.
+5.3.1) Permitted Execution Times: We formalize how constraints produced under $\delta$-BR restrict move execution and vertex occupancy times, establishing the framework for analyzing constraint accumulation.
+  2. 5.3.2) CSIPP Planning Behavior: We explain how CCBS’s underlying path planner, CSIPP, constructs plans, focusing on how it selects execution times from safe intervals.
+  3.
+5.3.3) Non-degenerate Reductions For Move Actions: We show how $\delta$-BR applied to a timed move action (in a move-move or move-wait conflict) reduces feasible move execution times by a non-degenerate amount.
   4.
-V-C4) Non-degenerate Reductions For Wait Actions: We show how our branching rule applied to a timed wait action reduces feasible move execution or feasible vertex occupancy times by a non-degenerate amount.
+5.3.4) Non-degenerate Reductions For Wait Actions: We show how $\delta$-BR applied to a timed wait action reduces feasible move execution or feasible vertex occupancy times by a non-degenerate amount.
   5.
-V-C5) Bounded Plan Impossibility: Using the results from Sections V-C3 and V-C4, and the fact that all timed actions in a bounded-duration plan must end before a fixed time, we show that the feasible move execution and vertex occupancy times are exhausted after finite iterations of the CCBS search.
-Therefore, we show that CCBS under our branching rule does not satisfy the non-termination requirement and therefore is guaranteed to terminate.
-V-C1 Permitted Execution Times
+5.3.5) Bounded Plan Impossibility: Using the results from Sections 5.3.3 and 5.3.4, and the fact that all timed actions in a bounded-duration plan must end before a fixed time, we show that the feasible move execution and vertex occupancy times are exhausted after finite iterations of the CCBS search.
+Therefore, we show that CCBS under $\delta$-BR does not satisfy the non-termination requirement and therefore is guaranteed to terminate.
+5.3.1 Permitted Execution Times
 To prove that the non-termination requirement cannot be satisfied, we establish a framework to analyze how constraints accumulate and restrict agent movements.
 At the CT root $N^{R}$ where the constraint set is empty, $N^{R}_{C}=\varnothing$, all agents are permitted to execute all actions at any time $t\geq 0$.
-However, at nodes further down in the CT with non-empty constraint sets, any number of constraints may restrict when agents can execute moves and occupy vertices.
+However, at nodes further down in the CT with non-empty constraint sets, any number of constraints may restrict when agents can execute move actions and occupy vertices.
 Thus, under the constraints $N_{C}$ at CT node $N$, we define for a specific agent $i$
@@
   • $\mathcal{T}_{v}\subseteq[0,\infty)$ as the set of all times when $i$ is permitted to occupy a vertex $v\in\mathcal{V}$.
-Since constraints under our branching rule restrict move execution and vertex occupancy over time intervals (Definition V.1), every set $\mathcal{T}_{x}$ where $x$ ranges over all moves $m\in\mathcal{A}$ and vertices $v\in\mathcal{V}$ can be described as the union of a finite number maximally connected time intervals:
-$\mathcal{T}_{x}=\bigcup_{h=1}^{N_{x}}S_{x}^{h},\quad S_{x}^{h}=[s_{x}^{h},e_{x}^{h})$ (11)
+Since constraints under the various branching rules restrict move execution and vertex occupancy over time intervals (Definition 5.1), every set $\mathcal{T}_{x}$ — with $x$ ranging over all moves $m\in\mathcal{A}$ and vertices $v\in\mathcal{V}$ — can be described as the union of a finite number of maximally connected time intervals:
+$\mathcal{T}_{x}=\bigcup_{h=1}^{N_{x}}S_{x}^{h},\quad S_{x}^{h}=[s_{x}^{h},e_{x}^{h}).$ (11)
 We refer to $S_{x}^{h}$ as a safe interval for $i$ to execute a move (when $x=m$) or occupy a vertex (when $x=v$).
-V-C2 CSIPP Planning Behavior
+5.3.2 CSIPP Planning Behavior
 Understanding how CSIPP selects execution times is crucial for proving that constraint additions cause non-degenerate reductions in feasible time intervals.
@@
 Each search state is represented by a tuple $(v,S_{v})$ where agent $i$ is located at $v\in\mathcal{V}$ during the safe interval $S_{v}=[s_{v},e_{v})\subseteq\mathcal{T}_{v}$.
-Let $t^{a}\in S_{v}$ be the time when $i$ arrives $v$.
-Unless state $(v,S_{v})$ is a goal state and $e_{v}=\infty$, agent $i$ must execute a move action to leave $v$ before the end of the safe interval at $e_{v}$.
+Let $t^{a}\in S_{v}$ be the time when $i$ arrives at $v$.
+Unless state $(v,S_{v})$ is a goal state and $e_{v}=\infty$, agent $i$ must execute a move action to leave $v$ before the end of the safe interval, $e_{v}$.
 For each move action $m$ with $\mathit{from}(m)=v$, agent $i$ can execute $m$ at any time $t$ satisfying
-  • $t\in[t^{a},e_{v})$ (after arriving and before it is unsafe),
-  • $t\in\mathcal{T}_{m}$ (it is safe to execute the move action), and
-  • $t+m_{D}\in\mathcal{T}_{v^{\prime}}$ (it is safe to occupancy $v^{\prime}$),
-where $v^{\prime}=\mathit{to}(m)$. In other words, $t\in P_{m}$ where
+  • $t\in[t^{a},e_{v})$ (executing $m$ after arriving at $v$ and before it is unsafe to remain there),
+  • $t\in\mathcal{T}_{m}$ (executing $m$ when it safe to do so), and
+  • $t+m_{D}\in\mathcal{T}_{v^{\prime}}$ (executing $m$ to arrive at the next node $v^{\prime}=\mathit{to}(m)$ when it is safe to do so),
+In other words, it is safe to execute $m$ at some time $t\in P_{m}$ where
 $P_{m}=[t^{a},e_{v})\cap\mathcal{T}_{m}\cap\mathcal{T}_{v^{\prime}}^{m_{D}}$ (12)
 and $\mathcal{T}_{v^{\prime}}^{m_{D}}=\{t-m_{D}\mid t\in\mathcal{T}_{v^{\prime}}\}$ to compensate for the arrival at $v^{\prime}$ being $m_{D}$ after $t$.
-By (11), $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$ are unions of maximally connected intervals, such that $P$ is also a union of maximally connected intervals.
-Fig.
-8 illustrates how $P$ is formed from the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
-Figure 8: Illustration of the construction of $P$ as the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
-For every maximally connected interval $s_{m},e_{m})\subseteq P_{m}$, an edge representing $i$ executing $\langle m,s_{m}\rangle$ is created, leading to the next state $(v^{\prime},S_{v^{\prime}})$ where $S_{v^{\prime}}\subseteq\mathcal{T}_{v^{\prime}}$ is the safe interval at $v^{\prime}$ such that $s_{m}+m_{D}\in S_{v^{\prime}}$.
-CSIPP always selects the earliest time $s_{m}$ for $i$ to execute $m$.
-This earliest-time selection property is essential for our proof: since the earliest time $s_{m}$ in $s_{m},e_{m})$ is selected, it is not possible to select an earlier time $t^{\prime}<s_{m}$ arbitrarily close to $s_{m}$ for the move action $m$ to be executed.
-V-C3 Non-degenerate Reductions For Move Actions
-When our branching rule is applied to a move action $\langle m^{i},t^{i}\rangle\in\pi_{i}$ in a move-move or move-wait conflict, a constraint is created which forbids the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$.
-We assume that $\pi_{i}$ was returned from CSIPP, such that a safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t_{i}$ was identified during the search.
-Consequently, the new constraint removes either a prefix or the entirety of $s_{m},e_{m})$ from $P_{m}$.
-Starting with a move-move conflict, our branching rule forbids the execution of $m^{i}$ during the unsafe interval $t^{i},t^{i}_{u})$, which is non-degenerate by construction.
-Consider instead a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$ and intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$.
-Our branching rule creates a constraint forbidding the execution of $m^{i}$ during $t^{i},t^{i}+\delta)$.
-If $\delta$ is bounded away from 0, then $t^{i},t^{i}+\delta)$ is non-degenerate.
-By (7), $\delta=\min\left(\gamma|\bar{I^{c}}|,t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}\right)$ is the minimum of two positive quantities:
-  • $\delta=\gamma|\bar{I^{c}}|$: It holds that $\delta$ is bounded away from zero since $\bar{I^{c}}$ is non-degenerate and $\gamma\in(0,1)$ is a fixed, non-zero value.
-  • $\delta=t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$: From (5) we know that $I^{c}\subseteq\bar{I^{c}}$ and $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$.
+By (11), $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$ are unions of maximally connected intervals, such that $P_{m}$ is also a union of maximally connected intervals.
+Figure 8 illustrates how $P_{m}$ is formed from the intersection between $t^{a},e_{v})$, $\mathcal{T}_{m}$ and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
+Figure 8: Illustration of the construction of $P_{m}$ as the intersection between $t^{a},e_{v})\subseteq\mathcal{T}_{v}$, $\mathcal{T}_{m}$, and $\mathcal{T}_{v^{\prime}}^{m_{D}}$.
+For every maximally connected interval $s_{m},e_{m})\subseteq P_{m}$, CSIPP creates an edge representing that agent $i$ executes $\langle m,s_{m}\rangle$, leading to the next state $(v^{\prime},S_{v^{\prime}})$ where $S_{v^{\prime}}\subseteq\mathcal{T}_{v^{\prime}}$ is the safe interval at $v^{\prime}$ such that $s_{m}+m_{D}\in S_{v^{\prime}}$.
+CSIPP always selects the earliest time $s_{m}$ in a safe interval $s_{m},e_{m})\subseteq P_{m}$ for $i$ to execute $m$.
+This earliest-time selection property is essential for our proof: since the earliest time $s_{m}$ in $s_{m},e_{m})$ is selected, it is not possible to select an earlier time $t^{\prime}<s_{m}$ arbitrarily close to $s_{m}$ for the move action $m$ to be executed, since $t^{\prime}\not\in P_{m}$.
+5.3.3 Non-degenerate Reductions For Move Actions
+When $\delta$-BR is applied to a move action $\langle m^{i},t^{i}\rangle\in\pi_{i}$ in a move-move or move-wait conflict, the motion constraint $\langle i,m^{i},[t^{i},t^{\prime})\rangle$ is created.
+Assuming that $\pi_{i}$ was returned from CSIPP, a safe interval $s_{m^{i}},e_{m^{i}})\subseteq P_{m^{i}}$ with $s_{m^{i}}=t_{i}$ was identified during the search to construct $\pi_{i}$.
+We will first show that the constrained interval $t^{i},t^{\prime})$ is non-degenerate, and then consequently that the new constraint removes either a prefix or the entirety of $s_{m^{i}},e_{m^{i}})$ from $P_{m^{i}}$.
+If $\langle m^{i},t^{i}\rangle$ is contained in a move-move conflict, $\delta$-BR forbids the execution of $m^{i}$ during the unsafe interval $t^{i},t^{i}_{u})$ which is non-degenerate by construction.
+Thus, in this case we know that the constrained interval $t^{i},t^{\prime})=[t^{i},t^{i}_{u})$ is non-degenerate.
+Consider instead that $\langle m^{i},t^{i}\rangle$ is contained in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ and collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$.
+$\delta$-BR creates the motion constraint $\langle i,m^{i},[t^{i},t^{i}+\delta)\rangle$.
+The constrained interval $t^{i},t^{\prime})=[t^{i},t^{i}+\delta)$ is non-degenerate only if $\delta$ is bounded away from 0.
+By (7), $\delta$ is the minimum of two positive quantities, $\gamma|\bar{I^{c}}|$ and $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$:
+  • $\gamma|\bar{I^{c}}|$ is bounded away from zero since $\bar{I^{c}}$ is non-degenerate and $\gamma\in(0,1)$ is a fixed, non-zero value.
+  • From (5) we know that $I^{c}\subseteq\bar{I^{c}}$ and $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$.
 The former implies that $\bar{t_{1}^{c}}\leq t_{1}^{c}<t_{2}^{c}\leq\bar{t_{2}^{c}}$ where strict inequality holds since $I^{c}$ is non-degenerate.
@@
     $\bar{t_{1}^{c}}$ $<t^{j}+w^{j}_{D},$
-    $0$ $<t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}=\delta,$
-    such that $\delta$ is bounded away from 0.
-Hence, $\delta$ is bounded away from zero, implying that the constrained interval $t^{i},t^{i}+\delta)$ is non-degenerate.
-In both cases, where $\langle m^{i},t^{i}\rangle\in\pi_{i}$ is contained in a move-move or move-wait conflict, a constraint is created forbidding the execution of $m^{i}$ during a non-degenerate interval $t^{i},t^{\prime})$.
-That is, $t^{i},t^{\prime})$ is removed from $\mathcal{T}_{m}$.
-Since $s_{m}=t^{i}$ for some $s_{m},e_{m})\subseteq P_{m}$, by (12), either the entire safe interval $s_{m},e_{m})$ is removed (if $t^{\prime}\geq e_{m}$) or a non-degenerate prefix $s_{m},t^{\prime})$ is removed from $P_{m}$ (if $t^{\prime}<e_{m}$).
-V-C4 Non-degenerate Reductions For Wait Actions
-In this section, we consider the constraints created when applying our branching rule to a wait action $\langle w^{j},t^{j}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$.
-It will be shown that this leads to two cases.
-In the first case, we show that there exists a timed move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ following $\langle w^{j},t^{j}\rangle$, and a constraint is created that forbids the execution of $m$ during a non-degenerate interval $t^{j}+w^{j}_{D},t^{\prime})$.
-Therefore, the same conclusions as in Section V-C3 are drawn.
-In the second case, a constraint is created forbidding the use of $v(w^{j})$ during a non-degenerate interval $t^{\prime},t^{\prime\prime})$, necessarily leading to a non-degenerate reduction in $\mathcal{T}_{v(w^{j})}$.
-When our branching rule is applied to $\langle w^{j},t^{j}\rangle$, two constraints are created (Definition V.1).
-One constraint forbids $j$ from occupying $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, and the other forbids the execution of any move action $m$ with $\mathit{from}(m)=v(w^{j})$ at any time in $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$.
+    $0$ $<t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}.$
+    Thus, $t^{j}+w^{j}_{D}-\bar{t_{1}^{c}}$ is bounded away from zero.
+Since $\delta$ is the minimum of two positive quantities which are both bounded away from zero, $\delta$ is also bounded away from zero.
+Therefore, $t^{i},t^{\prime})=[t^{i},t^{i}+\delta)$ is non-degenerate.
+In both cases — when $\langle m^{i},t^{i}\rangle$ is contained in a move-move or move-wait conflict — a constraint $\langle i,m^{i},[t^{i},t^{\prime})\rangle$ is created for which we have shown above that $t^{i},t^{\prime})$ is non-degenerate.
+This constraint effectively removes $t^{i},t^{\prime})$ from $\mathcal{T}_{m^{i}}$.
+Since $t^{i}=s_{m^{i}}$ for some $s_{m^{i}},e_{m^{i}})\subseteq P_{m^{i}}$, by (12), either the entire safe interval $s_{m^{i}},e_{m^{i}})$ is removed (if $t^{\prime}\geq e_{m^{i}}$) or a non-degenerate prefix $s_{m^{i}},t^{\prime})$ is removed from $P_{m^{i}}$ (if $t^{\prime}<e_{m^{i}}$).
+This is a useful result as it shows that $\delta$-BR applied to any timed move action in a conflict necessarily leads to a non-degenerate time interval being removed from the possible times to execute the move action.
+5.3.4 Non-degenerate Reductions For Wait Actions
+In this section, we consider the constraints created by $\delta$-BR when applied to a wait action $\langle w^{j},t^{j}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ with intersection interval $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$ and collision interval $I^{c}=[t_{1}^{c},t_{2}^{c})$.
+$\delta$-BR creates the vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and motion constraints $\langle j,m,[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ for each $m\in\mathcal{A}:\mathit{from}(m)=v(w^{j})$.
+It will be shown that there exist two cases: when the overlap between the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and the wait interval $t^{j},t^{j}+w_{D}^{j}]$ is either singular or non-degenerate.
+For each case, we will show that either the same conclusions as in Section 5.3.3 can be drawn for the move action $\langle m,t^{j}+w^{j}_{D}\rangle$ following $\langle w^{j},t^{j}\rangle$ in $\pi_{j}$, or a non-degenerate reduction occurs in $\mathcal{T}_{v(w^{j})}$.
 We begin by showing that the constrained interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate by expanding $\bar{t_{1}^{c}}+\delta$ using (7):
@@
 $=\min\left(\bar{t_{1}^{c}}+\gamma(\bar{t_{2}^{c}}-\bar{t_{1}^{c}}),\;t^{j}+w^{j}_{D}\right)$
-$=\min\left((1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}},\;t^{j}+w^{j}_{D}\right)$ (13)
-which shows that $\bar{t_{1}^{c}}+\delta\leq(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}$.
+$=\min\left((1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}},\;t^{j}+w^{j}_{D}\right).$ (13)
+This shows that $\bar{t_{1}^{c}}+\delta\leq(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}$.
 Since $0<\gamma<1$ is fixed to some value less than $1$, it follows that $(1-\gamma)\bar{t_{1}^{c}}+\gamma\bar{t_{2}^{c}}<\bar{t_{2}^{c}}$ and therefore that $\bar{t_{1}^{c}}+\delta<\bar{t_{2}^{c}}$.
 Thus, $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is non-degenerate.
-Next, we show that the constrained $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ necessarily overlap with the time interval $t^{j},t^{j}+w^{j}_{D}]$ when $j$ occupies vertex $v(w^{j})$ while executing $\langle w^{j},t^{j}\rangle$.
+Next, we show that $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ necessarily overlaps with the wait interval $t^{j},t^{j}+w^{j}_{D}]$.
 To do so, we combine two conditions.
-First, under the assumption made in Section III-B that $I^{c}$ is non-degenerate together with $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$ by (5), it follows that $t^{j}<\bar{t_{2}^{c}}$ meaning that $j$’s occupancy of $v(w^{j})$ starts before the constrained interval ends.
-Second, (13) shows that $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$, implying that $j$’s occupancy of $v(w^{j})$ ends at or after the constrained interval begins.
-Together, these conditions are sufficient to guarantee overlap by at least a singular value.
-There exists two cases: the overlap between $t^{j},t^{j}+w^{j}_{D}]$ and $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ is either singular or non-degenerate.
+First, $I^{c}$ is assumed in Section 3.2 to be non-degenerate, and by (5) we have $I^{c}\subseteq[t^{j},t^{j}+w^{j}_{D}]$.
+Thus, it follows that
+$t^{j}<\bar{t_{2}^{c}}.$ (14)
+Second, (13) shows that $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$.
+These two conditions — $t^{j}<\bar{t_{2}^{c}}$ and $\bar{t_{1}^{c}}+\delta\leq t^{j}+w^{j}_{D}$ — together are sufficient to guarantee overlap between $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ and $t^{j},t^{j}+w^{j}_{D}]$ by at least a singular value.
+We now consider each of the cases of singular or non-degenerate overlap:
   • Singular: This implies that $\bar{t_{1}^{c}}+\delta=t^{j}+w^{j}_{D}$.
@@
 Thus, the left-hand-side is $\infty$ implying that the right-hand side containing $w^{j}_{D}$ is also $\infty$.
-Since $w^{j}$ does not have infinite duration and is therefore not the final infinite wait action, there is a subsequent move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ with $\mathit{from}(m)=v(w^{j})$ in $j$’s plan.
-Our branching rule creates a constraint forbidding $m$ from being executed during the non-degenerate interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w^{j}_{D},\bar{t_{2}^{c}})$.
-The same reasoning and conclusions as in Section V-C3 for move actions in move-move and move-wait conflicts applies: a non-degenerate prefix or the entirety of some safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t^{j}+w^{j}_{D}$ identified during the CSIPP search is removed from $P_{m}$.
+Since $w^{j}$ does not have infinite duration and is therefore not the final infinite wait action, there exists a subsequent move action $\langle m,t^{j}+w^{j}_{D}\rangle\in\pi_{j}$ with $\mathit{from}(m)=v(w^{j})$.
+$\delta$-BR creates a constraint forbidding $m$ from being executed during the non-degenerate interval $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w^{j}_{D},\bar{t_{2}^{c}})$.
+The same reasoning and conclusions as in Section 5.3.3 for move actions in move-move and move-wait conflicts applies: a non-degenerate prefix or the entirety of some safe interval $s_{m},e_{m})\subseteq P_{m}$ with $s_{m}=t^{j}+w^{j}_{D}$ identified during the CSIPP search is removed from $P_{m}$.
   • Non-degenerate: We know that $t^{j},t^{j}+w^{j}_{D}]\subseteq\mathcal{T}_{v(w^{j})}$ else the timed wait action $\langle w^{j},t^{j}\rangle$ would not have been returned by CSIPP.
@@
 In other words, this tells us that a non-degenerate part of $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ exists in $\mathcal{T}_{v(w^{j})}$.
-Our branching rule creates a constraint forbidding the occupancy of $v(w^{j})$ during $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$, therefore removing $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ from $\mathcal{T}_{v(w^{j})}$.
+$\delta$-BR creates a vertex constraint $\langle j,v(w^{j}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and thereby removes $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})$ from $\mathcal{T}_{v(w^{j})}$.
 Thus, a non-degenerate interval is removed from $\mathcal{T}_{v(w^{j})}$.
-V-C5 Bounded Plan Impossibility
-Consider a CT node $N\in\varsigma$ mapping to $\tau^{\infty}$.
-At $N$, a plan $\pi\in N_{\Pi}$ with $\pi\sim\tau^{\infty}$ is branched on.
-To satisfy the non-termination condition (I), $\pi_{D}\leq c$ for some fixed $c<\infty$ which requires that every timed action in $\pi$ (except the final infinite wait action, implicitly omitted in the following) ends before $c$.
+5.3.5 Bounded Plan Impossibility
+Consider a CT node $N\in\varsigma$ mapping to $\tau^{\infty}$, meaning that a plan $\pi\in N_{\Pi}$ with $\pi\sim\tau^{\infty}$ is branched on.
+To satisfy the non-termination condition (I), $\pi_{D}\leq c$ for some fixed $c<\infty$ which requires that every timed action in $\pi$ (except the final infinite wait action, which we will momentarily address specifically) ends before $c$.
 For simplicity, we suffice with the less restrictive requirement that all timed move actions must begin before $c$.
@@
 where $\mathbf{v}=\bigcup_{m\in\tau^{\infty}}\left\{\mathit{from}(m),\mathit{to}(m)\right\}$ contains all vertices visited along $\tau^{\infty}$.
-In Section V-C3, we showed that when our branching rule is applied to a move action $\langle m,t\rangle\in\pi$, then either a non-degenerate prefix or the entirety of a maximally connected interval $s_{m},e_{m})\subseteq P_{m}$ is removed from $P_{m}$.
-Similarly, it was shown in Section V-C4 that when our branching rule is applied to a wait action $\langle w,t\rangle\in\pi$, then one of two cases occur.
-In the first case, the same constraining as when our branching rule is applied to a move action occurs for the move action following $\langle w,t\rangle$ in $\pi$.
+Before proceeding, we address the case of $\delta$-BR being applied to an infinite wait action $\langle w^{j}_{n},t^{j}_{n}\rangle\in\pi_{j}$ in a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j}_{n},t^{j}_{n}\rangle\rangle$, with $\pi_{j}\sim\tau^{\infty}$.
+Let this conflict have intersection interval $\bar{I^{c}}$.
+$\delta$-BR creates a vertex constraint $\langle j,v(w^{j}_{n}),[\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})\rangle$ and several motion constraints (which are not necessary to consider for this analysis).
+Thus, to satisfy the vertex constraint and also follow a plan mapping to $\tau^{\infty}$ which ends with waiting at $v(w^{j}_{n})$ indefinitely, agent $j$ can only arrive at $v(w_{n})$ from time $\bar{t_{2}^{c}}$ and later.
+By (14) we know that $t_{n}^{j}<\bar{t_{2}^{c}}$.
+Consequently, the timed move action $\langle m^{j}_{n-1},t^{j}_{n-1}\rangle\in\pi_{j}$ preceding $\langle w^{j}_{n},t^{j}_{n}\rangle$ is no longer valid since it leads to the agent arriving at $v(w_{n})$ at $t^{j}_{n}$, that is, before $\bar{t_{2}^{c}}$.
+Instead, $m^{j}_{n-1}$ can only occur at some time later than $\bar{t_{2}^{c}}-m^{j}_{n,D}$.
+Since $m^{j}_{n-1}$ is the last move action in $\tau^{\infty}$, this branching leads to a non-degenerate increase in the duration of all plans $\pi^{\prime}\sim\tau^{\infty}$.
+In Section 5.3.3, we showed that when $\delta$-BR is applied to a move action $\langle m,t\rangle\in\pi$, then either a non-degenerate prefix or the entirety of a maximally connected interval $s_{m},e_{m})\subseteq P_{m}$ is removed from $P_{m}$.
+Similarly, it was shown in Section 5.3.4 that when $\delta$-BR is applied to a wait action $\langle w,t\rangle\in\pi$, then one of two cases occur.
+In the first case, the same constraining as when $\delta$-BR is applied to a move action occurs for the move action following $\langle w,t\rangle$ in $\pi$.
 In the second case, a non-degenerate reduction of $\mathcal{T}_{v(w)}$ occurs.
-Before proceeding, we analyze the one case which has yet to be addressed: when the action $\langle w_{n},t_{n}\rangle$ in $\pi\sim\tau^{\infty}$ that is branched on at $N$ is an infinite wait action, $w_{n,D}=\infty$.
-When our branching rule is applied to $\langle w_{n},t_{n}\rangle$, a constraint is created forbidding the use of $v(w_{n})$ during some interval $t^{\prime},t^{\prime\prime})$.
-From Section V-C4 we know that $t_{n}<t^{\prime\prime}$ by some non-degenerate amount.
-Thus, the earliest time the agent can arrive at $v(w_{n})$ to remain there indefinitely is at the later time $t^{\prime\prime}$ than $t_{n}$.
-Consequently, the timed move action $\langle m_{n-1},t_{n-1}\rangle\in\pi$ preceding $\langle w_{n},t_{n}\rangle$ is no longer valid since it leads to the agent arriving at $v(w_{n})$ before $t^{\prime\prime}$.
-Instead, $m_{n-1}$ can only occur at some time later than $t^{\prime\prime}-m_{n,D}$.
-Since $m_{n-1}$ is the last move action in $\tau^{\infty}$, these constraint additions lead to a non-degenerate increase in the duration of all plans $\pi^{\prime}\sim\tau^{\infty}$.
-From the above, at every branching there occurs a non-degenerate reduction in one of the finite sets $P_{m}$ or $\mathcal{T}_{v}$ (with $m\in\tau^{\infty}$ and $v\in\mathbf{v}$) or a non-degenerate increase the in duration of all valid plans $\pi^{\prime}\sim\tau^{\infty}$.
-Therefore, after a finite many branchings either one of the sets will be depleted or the duration of valid plans mapping to $\tau^{\infty}$ will grow beyond the fixed $c<\infty$.
-In either case, eventually after a finite many branchings no plan $\pi\sim\tau^{\infty}$ satisfying the accumulated constraints $N_{C}$ and $\pi_{D}<c$ will exist.
-To conclude, the requirement of non-termination which was established in Section V-B cannot be satisfied by CCBS under our branching rule.
+From the above, at every branching there occurs a non-degenerate reduction in one of the finite safe interval sets $P_{m}$ or $\mathcal{T}_{v}$ (with $m\in\tau^{\infty}$ and $v\in\mathbf{v}$) or a non-degenerate increase in the duration of all valid plans $\pi^{\prime}\sim\tau^{\infty}$ (when a final infinite wait action is branched on).
+Therefore, after a finite number of iterations, either one of the safe interval sets $P_{m}$ or $\mathcal{T}_{v}$ will be depleted or the duration of valid plans mapping to $\tau^{\infty}$ will grow beyond the fixed $c<\infty$.
+In either case, eventually — after a finite number of branchings — no plan $\pi\sim\tau^{\infty}$ exists that can satisfy both the accumulated constraints from these branchings and the plan duration bound $\pi_{D}<c$.
+To conclude, the requirement of non-termination which was established in Section 5.2 cannot be satisfied by CCBS using $\delta$-BR.
 Therefore, termination is guaranteed after a finite number of iterations.
-V-D Soundness and Solution Completeness
-Theorem V.2 (Sound and solution complete).
-CCBS under our branching rule is sound and solution complete.
+5.4 Soundness and Solution Completeness
+Based on the results from the previous sections, we finalize our proof that CCBS using $\delta$-BR is sound and solution complete.
+Theorem 5.2 (Sound and solution complete).
+CCBS using $\delta$-BR is sound and solution complete.
 Proof.
-Our branching rule was shown in Section V-A to be sound.
-By Theorem III.2, CCBS is therefore sound when using a sound branching rule.
-Additionally, Section V-C proved that CCBS under our branching rule terminates within finite iterations.
-Since CCBS under our branching rule is both sound and guaranteed to terminate, Theorem III.3 establishes that it is solution complete.
+$\delta$-BR was shown in Section 5.1 to be sound, which means by Theorem 3.2 that CCBS using $\delta$-BR is sound.
+Additionally, Section 5.3 proved that CCBS using $\delta$-BR terminates within finite iterations.
+With these soundness and termination guarantees, Theorem 3.3 establishes that CCBS using $\delta$-BR is solution complete.
 ∎
-VI Experimental Evaluation
-We solve the MAPFR problem in Figure 9 for minimal sum-of-costs using two variants of the publicly available CCBS implementation  [^3] https://github.com/PathPlanning/Continuous-CBS.: the unmodified CCBS release using the implemented branching rule, and an otherwise identical version using our branching rule  [^4] https://github.com/Adcombrink/S-and-SC-CCBS..
-Experiments are run on a Mac Studio, 16-core M4 Max CPU, 64 GB RAM, macOS Sequoia (15.3).
-We use $\gamma=0.5$ when calculating $\delta$.
-Figure 9: A MAPFR problem with four agents, each agent $i$ starts at vertex $S(i)$ and must reach its goal vertex $G(i)$.
+6 Experimental Evaluation
+We perform experiments with the publicly available CCBS using the IBR (CCBS-IBR), and an otherwise identical CCBS using the proposed $\delta$-BR (CCBS-$\delta$-BR).
+Both versions are available in our repository [^3] https://github.com/Adcombrink/Optimal-Continuous-CBS.
+along with details in A regarding necessary modifications to the original code base.
+A counterexample is introduced in Section 6.1 for which CCBS-$\delta$-BR produces an optimal solution and CCBS-IBR returns a sub-optimal solution, along with explanations as to why this occurs.
+Thereafter, we compare the solution quality and runtime for the two methods on benchmark problems in Section 6.2.
+Finally, in Section 6.3 we perform an ablation study on the parameter $\gamma$ used in $\delta$-BR, motivating the use of $\gamma=0.9$ unless stated otherwise.
+We use circular agents with radius $\sqrt{2}/4$ that traverse edges in straight lines at constant speed $1$.
+All experiments are run on a 2025 Mac Studio, 16-core M4 Max CPU, 64 GB RAM, macOS Sequoia (15.3).
+6.1 A Counterexample
+Figure 9: A MAPFR problem with four agents with radius $r=\sqrt{2}/4$.
+Each agent $i=1,..,4$ is shown at its start vertex $S(i)$ and must reach its goal vertex $G(i)$ without colliding with any other agent.
 The example in Figure 9 contains four agents, $i=1,..,4$, each with a start vertex $S(i)$ and goal vertex $G(i)$.
@@
 Agent $1$ starts at vertex $S(1)=E$ and must pass vertex $F$ to reach its goal $G(1)=G$, therefore, agent $2$ with $S(2)=G(2)=F$ must move out of the way.
-The solution found by both methods is for agent $2$ to move to $C$ to allow agent $1$ to pass.
+The solution found by both CCBS-IBR and CCBS-$\delta$-BR is for agent $2$ to move to $C$ to allow agent $1$ to pass.
 However, agent $2$ must avoid a collision with agent $3$ which passes vertex $C$ on its way to its goal vertex $D$.
@@
 Due to the difference in length of edges $EF$ and $AB$, the time that agent $4$ must wait for agent $3$ is larger than the time that agent $1$ must wait for agent $2$.
-Therefore, these two possible solutions have different sum-of-costs, only the one where agent $2$ waits for agent $3$ is optimal.
-We now examine how CCBS under the implemented branching rule handles this problem.
-At the CT root $N^{R}$, with $N^{R}_{C}=\varnothing$, each agent’s initial plan is generated without collision avoidance:
+Therefore, these two possible solutions have different sum-of-costs, and only the one where agent $2$ waits for agent $3$ is optimal.
+We now examine how CCBS-IBR handles this problem. At the CT root $N^{R}$, with $N^{R}_{C}=\varnothing$, each agent’s initial plan is generated without collision avoidance:
 $N^{R}_{\Pi}=\left\{\begin{split}\pi_{1}&=\langle\langle EF,0\rangle\langle FG,1.5\rangle\langle G,2.5,\infty\rangle\rangle\\ \pi_{2}&=\langle\langle F,0,\infty\rangle\rangle\\ \pi_{3}&=\langle\langle BC,0\rangle\langle CD,1\rangle\langle D,2,\infty\rangle\rangle\\ \pi_{4}&=\langle\langle AB,0\rangle\langle B,1,\infty\rangle\rangle\end{split}\right\}$
-where $\langle VU,t\rangle$ denotes a timed move action along edge $(V,U)$ starting at time $t$, and $\langle V,t_{1},t_{2}\rangle$ is a timed wait action at vertex $V$ from $t_{1}$ to $t_{2}$.
-Two conflicts appear in $N^{R}_{\Pi}$: $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ and $\langle\langle FG,1.5\rangle,\langle F,0,\infty\rangle\rangle$ (between agents $1$ and $2$).
-CCBS branches on the first conflict, with collision interval $I^{c}=[0.793,1.5)$.
-This yields the constraints $\langle 1,EF,[0,\infty)\rangle$ (forbidding agent $1$ from ever traversing $EF$) and $\langle 2,F,I^{c}\rangle$ (forbidding agent $2$ from occupying $F$ during $I^{c}$).
+where $\langle VU,t\rangle$ denotes a timed move action from $V$ to $U$ starting at time $t$, and $\langle V,t_{1},t_{2}\rangle$ is a timed wait action at vertex $V$ from $t_{1}$ to $t_{2}$.
+Two conflicts appear in $N^{R}_{\Pi}$: $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ and $\langle\langle FG,1.5\rangle,\langle F,0,\infty\rangle\rangle$, both between agents $1$ and $2$.
+CCBS-IBR branches on the first conflict, with intersection interval $\bar{I^{c}}=[0.793,1.5)$.
+This yields the constraints $\langle 1,EF,[0,\infty)\rangle$ (forbidding agent $1$ from ever traversing $EF$) and $\langle 2,F,\bar{I^{c}}\rangle$ (forbidding agent $2$ from occupying $F$ during $\bar{I^{c}}$).
 The first constraint is infeasible, as agent $1$ must traverse $EF$ to reach its goal $G$.
@@
 $\Pi=\left\{\begin{split}\pi_{1}&=\langle\langle EF,0\rangle\langle FG,1.5\rangle\langle G,2.5,\infty\rangle\rangle\\ \pi_{2}&=\langle\langle FC,0\rangle\langle C,1,1.5\rangle\langle CF,1.5\rangle\langle F,0,\infty\rangle\rangle\\ \pi_{3}&=\langle\langle B,0,1.5\rangle\langle BC,1.5\rangle\langle CD,2.5\rangle\langle D,3.5,\infty\rangle\rangle\\ \pi_{4}&=\langle\langle A,0,1.2\rangle\langle AB,1.2\rangle\langle B,2.2,\infty\rangle\rangle\end{split}\right\}.$
-We now examine how CCBS under our branching rule handles the same conflict.
-The conflict $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ with $I^{c}=[0.793,1.5)$ has an intersection interval, i.e., the time during which agent $1$ executing $\langle EF,0\rangle$ is in collision with agent $2$ waiting at vertex $F$, of $\bar{I^{c}}=I^{c}=[0.793,1.5)$).
-This yields
-$\delta=\min\left(0.5\,(1.5-0.793),\infty\right)=0.354$
+We now examine how CCBS-$\delta$-BR handles the same conflict. The conflict $\langle\langle EF,0\rangle,\langle F,0,\infty\rangle\rangle$ with intersection interval $\bar{I^{c}}=[0.793,1.5)$) yields
+$\delta=\min\left(0.5\cdot\,|[0.793,1.5)|,\infty\right)=0.354$
 and the following set of constraints:
@@
   • Agent $2$: vertex constraint $\langle 2,F,[1.17,1.5)\rangle$ and motion constraints $\langle 2,FE,[1.17,1.5)\rangle$, $\langle 2,FC,[1.17,1.5)\rangle$, $\langle 2,FG,[1.17,1.5)\rangle$.
-Unlike the constraints produced by standard CCBS, these restrictions exclude only infeasible behaviors while preserving feasible solutions — including the case where agent $2$ waits until after agent $3$ passes vertex $C$.
+Unlike the constraints produced by CCBS-IBR, these restrictions exclude only infeasible behaviors while preserving feasible solutions — including the case where agent $2$ waits until after agent $3$ passes vertex $C$.
 The resulting solution is:
@@
 where agent $2$ waits at $F$ until agent $3$ has traversed $C$.
-The performance using each branching rule is reported in Table I.
-There, we see that our branching rule yields a solution with lower sum-of-costs, makespan, and computation time than standard CCBS.
-Consequently, this provides additional evidence to that in [28] that CCBS is not sound.
-Additionally, with the support of the theoretical results in Section V, the solution found by CCBS under our branching rule has optimal sum-of-costs.
-TABLE I: Performance comparison between standard CCBS and CCBS using our branching rule.
-Branching rule Sum-of-Costs Makespan Computation Time
-Standard $10.707$ $3.500$ $8.1$ ms
-Ours $9.000$ $3.000$ $1.4$ ms
-VII Conclusions
+The performance using each branching rule is reported in Table 1.
+There, we see that CCBS-$\delta$-BR yields a solution with lower sum-of-costs, makespan, and computation time than CCBS-IBR.
+Consequently, this provides additional evidence to that in [29] and Section 4 that CCBS-IBR is not sound.
+Moreover, the theoretical results in Section 5 ensure that the solution found by $\delta$-BR has optimal sum-of-costs.
+However, we further strengthen this claim by validating the optimality of CCBS-$\delta$-BR’s solution by constructing a tailor-made Satisfiability Modulo Theory (SMT) model for this specific problem and solving it using the SMT solver Z3 [33], confirming after roughly $26$ hours of computation that the solution returned by $\delta$-BR is indeed optimal.
+Details regarding this validation can be found in B, and animations of CCBS-IBR and CCBS-$\delta$-BR’s solutions can be found in the provided repository.
+Table 1: Performance comparison between the CCBS-IBR and CCBS-$\delta$-BR on the example in Figure 9.
+Method Sum-of-Costs Makespan Computation Time
+CCBS-$\delta$-BR $9.000$ $3.000$ $1.0$ ms
+CCBS-IBR $10.707$ $3.500$ $7.1$ ms
+6.2 Benchmark Comparison
+We compare CCBS-IBR and CCBS-$\delta$-BR on the same $2^{k}$-neighborhood gridmaps and roadmaps benchmark sets from [21], and additionally on our own randomly generated gridlike roadmaps.
+The benchmarking scheme is the same as in [21]: for each map, a scenario defines a list of start and goal vertex pairs.
+We then create a problem with $n$ agents using the first $n$ start and goal vertex pairs.
+For each map and scenario, we begin by solving for $n=2$, and repeatedly incrementing $n$ until a problem cannot be solved within a time limit of $30$ seconds.
+The $2^{k}$-neighborhood gridmaps — originally introduced in [8] — define a grid of traversable and blocked cells.
+Each traversable cell is connected to its $2^{k}$ neighborhood, provided that a traversal can be done without colliding with a blocked cell.
+We test with the same parameters as in [21]: k=1,,5k=1,...,5 with each of the maps Den520d, Warehouse, Rooms, and Empty16x16.
+The roadmaps are based on the overall shape of Den520d with three levels of density: Sparse, Dense, and Mega-dense.
+We refer to [21] for additional details regarding these maps.
+In addition to the gridmaps and roadmaps, we introduce gridlike roadmaps which are generated with a desired average degree over the map vertices.
+Two example maps with average degrees of $2.0$ and $3.5$ are shown in Figure 10.
+By lowering the average degree, the map becomes more constrained in the number of paths between two vertices.
+The likelihood of agent interactions consequently increases.
+These maps are generated from a 2020–by–$10$ vertex grid with spacing 11 by applying zero-mean gaussian noise with standard deviation $0.15$ to the vertex positions, and randomly removing edges until a desired average vertex degree is achieved.
+Edges are only removed if the resulting graph remains connected.
+(a) Average vertex degree $3.5$.
+(b) Average vertex degree $2.0$.
+Figure 10: Two gridlike roadmaps with different average degrees and agent start positions.
+On the gridmap and roadmap benchmark sets, we find no difference in solution quality between CCBS-IBR and CCBS-$\delta$-BR; both methods produce equivalent solutions.
+However, on the gridlike roadmap benchmark set we find on $11$ problems that CCBS-$\delta$-BR produces solutions with a lower sum-of-costs than CCBS-IBR, and equivalent solutions on all remaining problems.
+Although these $11$ problems represent only a fraction of the 15 46215\,462 problems that both methods could solve within the time limit, it shows that CCBS-IBR’s sub-optimality issues do not only present themselves in tailor-made problems.
+These results also hint at the difficulty of producing benchmark problems that fully test the limits of these algorithms.
+Additionally, the $11$ problems that trigger CCBS-IBR to return sub-optimal solutions are on maps with an average node degree distribution toward the lower end of the tested range: $1$ with average degree $2.1$, 33 with average degree $2.2$, 44 with average degree $2.4$, and $3$ with average degree $2.8$.
+This is perhaps unsurprising, as CCBS-IBR’s unsoundness is due to how it handles move-wait conflicts; with fewer path options due to a lower average degree, the likelihood that the available paths include a vertex with a stationary agent increases.
+Over these $11$ problems, CCBS-IBR returns mean $0.243$ and maximum $0.569$ higher sum-of-costs than CCBS-$\delta$-BR.
+Finally, we note that it is possible for CCBS-IBR— being unsound – to remove all solutions to a problem and therefore never terminate.
+Such cases are not reflected in our results as we only compare cases where both CCBS-IBR and CCBS-$\delta$-BR return a solution.
+We compare the runtime of CCBS-IBR and CCBS-$\delta$-BR in Figure 11 by measuring the number of problems each method is able to solve within a given time limit.
+To account for different benchmark set sizes, values are normalized to the number of problems CCBS-IBR solves within $30$ seconds.
+Across all benchmark sets, CCBS-IBR is able to solve more problems within the given time limit.
+This advantage stems from CCBS-IBR’s more aggressive handling of move-wait conflicts.
+While this strategy sometimes removes solutions from the search, it often retains at least one optimal solution while reducing the size of the search space.
+Consequently, CCBS-IBR generally requires fewer node expansions than CCBS-$\delta$-BR, as illustrated in Figure 12.
+The difference is most pronounced on gridlike roadmaps, followed by roadmaps, suggesting that move-wait conflicts occur more frequently in these environments compared to gridmaps.
+However, CCBS-IBR’s aggressive pruning comes with trade-offs.
+In certain cases, as shown above, all optimal solutions are eliminated, leading to sub-optimal results.
+This means that we are effectively comparing an optimal method (CCBS-$\delta$-BR) to a sub-optimal method (CCBS-IBR).
+Furthermore, problems where all solutions are eliminated by CCBS-IBR are not reflected in these results, since CCBS-IBR fails to terminate in such cases and contributes to neither the runtime comparison nor the solution quality evaluation above.
+Figure 11: The number of solved problems within a given runtime, for each method on each benchmark set, normalized by the method with most solved problems.
+Figure 12: Probability density histogram of the increase in the number of high-level node expansions using CCBS-$\delta$-BR compared to CCBS-IBR.
+The y-axes are limited to maximum $0.8$; the height of the $(1.0,1.1)$-bins are shown specifically.
+6.3 Ablation Study on Parameter $\gamma$
+The effect of the parameter $\gamma$ — used to compute $\delta$ in CCBS-$\delta$-BR (see Definition 5.1) — on solver runtime is investigated by solving a set of benchmark problems with $\gamma\in\left\{0.1,0.3,...,0.9\right\}$ and recording the number of problems solved within a given time limit.
+The benchmark set contains gridlike roadmaps: for every average degree value $2.0,2.1,\dots,3.5$, we generate $5$ maps, each with $5$ scenarios.
+To avoid selection bias, this benchmark is separate from the gridlike roadmap set used in Section 6.2.
+The results are reported in Figure 13, showing the number of problems that can be solved within $30$ seconds for each value of $\gamma$ and average degree value.
+We note two things: first, larger $\gamma$ values lead to more problems being solved.
+Therefore, on these gridlike roadmaps, larger values of $\gamma$ reduce the solver runtime.
+Second, we see that more problems can be solved within the given time limit on maps with a larger average vertex degree.
+This is likely due to the larger number of possible paths that an agent can take from its start to goal vertex, making these problems less constrained.
+At these higher average vertex degree values, the different between the performance using different $\gamma$ values becomes more prominent when compared to lower vertex degrees.
+Figure 13: The number of problems solved within a $30$ second time limit for various values of $\gamma$, shown over a range of average vertex degrees on gridlike roadmaps.
+7 Conclusions
 This work set out to resolve the recently revealed long-standing gap between the theory and practice of Continuous-time Conflict Based Search (CCBS) for optimal Multi-Agent Path Finding in continuous time (MAPFR).
-We revisited the MAPFR formulation and analyzed the two existing branching schemes — the theoretical rule and the publicly available implementation — through a new analytical framework.
-This analysis yielded simple, sufficient conditions for soundness and solution completeness and showed that the implemented branching rule violates them, thereby supporting the counter-examples that cast doubt on CCBS.
-We then presented a our novel branching rule and proved that it meets all conditions: it never excludes solutions from the search and guarantees termination when used in CCBS.
-Consequently, the resulting CCBS variant is both sound and solution complete, matching the guarantees of discrete-time CBS for the first time in the continuous domain.
-We validated these results by presenting an example MAPFR problem and applying standard CCBS and CCBS using our branching rule.
-Results shows that our branching rule produces a solution with lower sum-of-costs than standard CCBS, thereby showing definitively that standard CCBS is not sound.
-Practically, our rule can be adopted as a drop-in replacement in existing CCBS code-bases to obtain certificates of optimality without risking non-termination.
-Conceptually, the constraint-set perspective and non-termination criterion offer tools for analyzing other CCBS-like MAPFR solvers and extensions, laying a solid foundation for the next generation of multi-agent path finding algorithms.
+We revisited the MAPFR formulation, introduced a new analytical framework for CCBS-style algorithms, and analyzed the publicly available reference implementation of CCBS.
+The framework provided simple, sufficient conditions for soundness and solution completeness, revealing that the reference implementation violates these conditions — thereby supporting recent counterexamples that challenge the validity of CCBS.
+To address this, we proposed a new branching rule ($\delta$-BR) and proved that CCBS using $\delta$-BR satisfies all conditions: it preserves all solutions and guarantees termination within a finite number of iterations on any solvable MAPFR problem.
+Thus, CCBS using $\delta$-BR is both sound and solution complete for sum-of-costs, makespan, and any other objective function that is strictly monotonically increasing with respect to the maximum agent arrival time.
+To our knowledge, this represents the first method for MAPFR with the same guarantees as Conflict Based Search for the discrete-time MAPF problem.
+We illustrated the difference between standard CCBS and CCBS using $\delta$-BR with a counterexample, showing that standard CCBS returns a worse solution than CCBS with $\delta$-BR, thereby showing that standard CCBS is not sound.
+Furthermore, we compared the two approaches on benchmark problems.
+While standard CCBS often finds solutions faster due to its more aggressive pruning of the search space, this comes at the cost of discarding all optimal solutions in some cases.
+In contrast, our branching rule guarantees solution optimality without sacrificing termination.
+Finally, we show with our provided implementation that $\delta$-BR can be adopted as a drop-in replacement in existing CCBS implementations.
+Moreover, the analytical framework and non-termination criterion introduced in this work provide analytical tools for reasoning about CCBS-like solvers and their extensions, thereby providing a rigorous foundation for the next generation of continuous-time multi-agent path finding algorithms.
+8 CRediT Authorship Contribution Statement
+Alvin Combrink: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Writing - Original Draft.
+Sabino Francesco Roselli: Software, Writing - Review & Editing, Supervision.
+Martin Fabian: Writing - Review & Editing, Supervision, Funding acquisition.
+9 Declaration of Competing Interest
+Alvin Combrink reports financial support was provided by Sweden’s Innovation Agency.
+Sabino Francesco Roselli reports financial support was provided by Sweden’s Innovation Agency.
+Martin Fabian reports financial support was provided by Sweden’s Innovation Agency.
+If there are other authors, they declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+10 Acknowledgments
+We gratefully acknowledge the Vinnova project CLOUDS (Intelligent algorithms to support Circular soLutions fOr sUstainable proDuction Systems), and the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation.
+Appendix A Implementation Details
+Our implementation of CCBS-IBR  [^4] https://github.com/Adcombrink/Optimal-Continuous-CBS/tree/originalCCBS closely follows that of the publicly available CCBS implementation [^5] https://github.com/PathPlanning/Continuous-CBS from [21].
+However, minor modifications have been made to ensure a fair comparison between CCBS-IBR and CCBS-$\delta$-BR  [^6] https://github.com/Adcombrink/Optimal-Continuous-CBS.
+For more details than what we describe here, we refer the reader to the provided repositories.
+First, we do not use any high-level heuristics or disjoint splitting for CCBS-IBR or CCBS-$\delta$-BR, as these are not the focus of this work.
+Therefore, ensuring compatibility of CCBS-$\delta$-BR with these additional improvements are left for future work.
+We do however enable Cardinal Constraints.
+Second, in the implementation from [21], move-wait conflict detection is done using the same method as for move-move conflicts: a binary interval search with a specified precision.
+CCBS-$\delta$-BR, however, is highly sensitive to small numerical differences between the detection of move-wait conflicts and the computation of $\bar{I^{c}}=[\bar{t_{1}^{c}},\bar{t_{2}^{c}})$.
+Specifically, for a move-wait conflict $\langle\langle m^{i},t^{i}\rangle,\langle w^{j},t^{j}\rangle\rangle$ for which $\delta=t^{j}+w_{D}^{j}-\bar{t_{1}^{c}}$ (see Definition 5.1), the resulting constraint interval using in the constraints on $j$ is $\bar{t_{1}^{c}}+\delta,\bar{t_{2}^{c}})=[t^{j}+w_{D}^{j},\bar{t_{2}^{c}})$.
+This interval overlaps with $j$’s wait action time interval $t^{j},t^{j}+w_{D}^{j}]$ by only a singular time instant, $t^{j}+w_{D}^{j}$.
+Thus, if the constraint interval is subject to numerical imprecision such that it is computed to $t^{j}+w_{D}^{j}+\epsilon,\bar{t_{2}^{c}})$ for some small $\epsilon>0$, then the resulting constraints using this interval will not forbid $\langle w^{j},t^{j}\rangle$ and therefore the detected move-wait conflict may occur again.
+Our solution to this is straight-forward: we use $\bar{I^{c}}$ for both constraint creation and move-wait collision detection.
+An alternative could be to use a fixed precision value as additional margin on the constraints, however, this could be regarded as a form of discretization which we avoid.
+Besides, since $\bar{I^{c}}$ is computed analytically (under the assumption that edges are traversed in straight lines, as in all experiments), the collision detection is not subject to a precision value and is therefore more precise.
+Moreover, for the case of general motion functions, CCBS-$\delta$-BR’s numerical sensitivity is easily overcome by using the same computations for collision detection as for constraint creation (such as by pre-computing intersection intervals with, e.g., a binary interval search).
+Therefore, these modifications do not require agents to only traverse edges in straight lines.
+As a consequence of using $\bar{I^{c}}$ in the collision detection for CCBS-$\delta$-BR, the same collision detection is used in CCBS-IBR to avoid any differences in the underlying problem that these two methods solve.
+Finally, the computation of $\bar{I^{c}}$ is re-implemented for more robustness to numerical imprecision, particularly for near-tangent collisions where the collision interval is singular or close to singular.
+Appendix B Validating the Optimality of CCBS-$\delta$-BR’s Solution
+We validated the optimality of CCBS-$\delta$-BR’s solution to the counterexample in Section 6.1 using a tailor-made Satisfiability Modulo Theory (SMT) model with the SMT solver Z3 [33].
+After roughly $26$ hours of computation, the model returned an equivalent solution with sum-of-costs $9.0$, thereby confirming that CCBS-$\delta$-BR’s solution is an optimal solution.
+The exact implementation and results can be found in our repository [^7] https://github.com/Adcombrink/Optimal-Continuous-CBS.
+The model requires specifying an exact number of timed actions that each agent executes.
+Thus, the question arises: how many timed actions must be set to ensure that the returned solution is optimal with respect to the MAPFR problem?
+In the following, we will establish that any solution with a given sum-of-costs requires at most a certain number of timed move actions; if any more timed move actions are required by a single agent, then the sum of costs of such a solution is necessarily higher than the given sum-of-costs.
+We find for this specific problem that any solution requiring more than $5$ timed move actions necessarily has a sum-of-costs >9.0>9.0.
+Considering that agents may also execute timed wait actions, we find that the lower bound on the number of timed actions is $11$.
+Furthermore, we motivate that there does not exists an upper bound to the number of timed actions.
+Thus, with these results, it holds that the retained solution (which is equivalent to the solution from CCBS-$\delta$-BR) is optimal.
+Let $d^{\mathit{LB}}_{i}$ denote an optimistic lower bound on the duration of a plan taking agent $i$ from $S(i)$ to $G(i)$, which is obtained by not considering potential collisions with other agents:
+$d^{\mathit{LB}}_{1}=2.5,\qquad d^{\mathit{LB}}_{2}=0,\qquad d^{\mathit{LB}}_{3}=2.0,\qquad d^{\mathit{LB}}_{4}=1.0.$
+For solution $\Pi$, let $n$ be the maximum number of timed move actions contained in any single plan $\pi\in\Pi$.
+Since the minimum move action duration over all move actions in the counterexample is $1.0$, a plan with $n$ move actions has duration $\geq n$.
+The sum-of-costs lower bound $\mathit{SOC}^{\mathit{LB}}_{\Pi}$ for $\Pi$ occurs when one agent executes $n$ moves while others follow their optimistically shortest-duration plans:
+$\mathit{SOC}^{\mathit{LB}}_{\Pi}$ $=\min_{i=1,..4}\left\{n+\sum_{j\in\{1,..,4\}\setminus\left\{i\right\}}d^{\mathit{LB}}_{j}\right\}=\min\left\{\begin{gathered}n+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{3}+d^{\mathit{LB}}_{4}=n+3.0\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{3}+d^{\mathit{LB}}_{4}=n+5.5\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{4}=n+3.5\\ n+d^{\mathit{LB}}_{1}+d^{\mathit{LB}}_{2}+d^{\mathit{LB}}_{3}=n+4.5\end{gathered}\right\}=n+3.0.$
+From this, if any solution has sum-of-costs <9.0<9.0, then $n+3.0<9.0$, implying $n\leq 5$.
+Since consecutive wait actions can be merged, an agent executing $n$ move actions needs at most $n+1$ wait actions.
+Therefore, the maximum number of timed actions per agent is $n+n+1=2n+1=11$.
+Moreover, since wait actions can be split arbitrarily, enforcing that each agent executes exactly $11$ timed actions still allows for an optimal solution to be found.